### v1

# Unsupervised full-field Bayesian inference of orthotropic hyperelasticity from a single biaxial test: a myocardial case study

Abstract
Cardiac muscle tissue exhibits highly non-linear hyperelastic and orthotropic material behavior during passive deformation. Fully capturing this behavior in traditional homogenized tissue testing requires the excitation of multiple deformation modes, i.e. combined triaxial shear tests and biaxial stretch tests. Inherently, such multimodal experimental protocols necessitate multiple tissue samples and extensive sample manipulations. Intrinsic inter-sample variability and manipulation-induced tissue damage might have an adverse effect on the inversely identified tissue behavior. In this work, we aim to overcome this gap by focusing our attention to the use of heterogeneous deformation profiles in a parameter estimation problem. More specifically, we adapt EUCLID, an unsupervised method for the automated discovery of constitutive models, towards the purpose of parameter identification for highly nonlinear, orthotropic constitutive models using a Bayesian inference approach and three-dimensional continuum elements. We showcase its strength to quantitatively infer – with varying noise levels – the material model parameters of synthetic myocardial tissue slabs from a single heterogeneous biaxial stretch test. This method shows good agreement with the ground-truth simulations and with corresponding credibility intervals. Our work highlights the potential for characterizing highly nonlinear and orthotropic material models from a single biaxial stretch test with uncertainty quantification.
keywords:
material model inference , full-field data , uncertainty quantification , anisotropic hyperelasticity , experimental tissue testing , Bayesian inference
journal: arXiv
[…]
1 Introduction
Soft biological tissues exhibit complex mechanical responses due to their fibrous, anisotropic microstructure. Accurately characterizing this behavior is essential for constructing predictive constitutive models that support simulation-driven design, diagnostics, and treatment planning (Peirlinck et al., 2021). However, unlike engineered materials, soft tissues are subject to considerable biological variability, both within a patient, a patient cohort, and experimental study (Reeps et al., 2012; Roccabianca et al., 2014; Budday et al., 2017; Matouš et al., 2017). This intra- and inter-sample variability is even more pronounced in the myocardium, where strong spatial gradients in myofiber orientation and laminar architecture govern regional mechanical function (Rohmer et al., 2007; Alexander J. Wilson, 2022). In this study we investigate the impact that the underlying microstructural organization variability has on the mechanical behavior of a single sample.
Traditional approaches to constitutive modeling of such tissues typically rely on multiple mechanical tests across different deformation modes including uniaxial tension or compression, biaxial tension, or shear testing (Dokos et al., 2002; Sommer et al., 2015; Kakaletsis et al., 2021). These protocols often require combining data from different specimens and, as such, ignore and smooth out this intrinsic intra- and inter-sample variability and lead to a smoothened out, average response (Staber and Guilleminot, 2017). Moreover, these protocols involve strong assumptions, such as spatially homogeneous stress and strain fields within the region of interest (Holzapfel and Ogden, 2009; Guan et al., 2018; Martonová et al., 2024). Such assumptions can be problematic for living tissues, where spatial heterogeneity is the norm, and the mechanical response may depend on subtle differences in fiber orientation, geometry, or mounting. Moreover, invasive sample manipulations, such as cutting, clamping, and gluing, can alter the intrinsic mechanical behavior that these tests seek to measure (Fehervary et al., 2018; Alloisio et al., 2024).
These challenges have prompted the development of single-sample, single-experiment methodologies that aim to infer full constitutive behavior from a single loading protocol. This shift has been enabled by advances in full-field deformation measurements, such as digital image correlation and imaging-based 3D kinematics (Wang et al., 2021; Jailin et al., 2024; Meng et al., 2025). When applied to samples undergoing spatially heterogeneous deformation, such data enable the inference of multiple material parameters simultaneously, reducing the need for multiple loading modes. Indeed, unsupervised identification techniques, such as the virtual fields method (Grédiac et al., 2006; Pierron and Grédiac, 2012), finite element updating (Kavanagh and Clough, 1971; Peirlinck et al., 2018; Elouneg et al., 2021), variational systems identification (Wang et al., 2021), the equilibrium gap method (Claire et al., 2004; Peyraut and Genet, 2025), and EUCLID (Flaschel et al., 2021; Thakolkaran et al., 2022; Joshi et al., 2022; Thakolkaran et al., 2025), demonstrate that heterogeneous strain states, once seen as a nuisance, can be leveraged as rich sources of information, obviating the need for homogenizing assumptions. In this study, we explore the validity of a simple biaxial tensile testing protocol on a single sample towards the identification of the underlying orthotropic behavior of a myocardial tissue slab.
Yet, a key limitation persists in most current workflows: the lack of uncertainty quantification. Material parameters are frequently reported as point estimates, with little accounting for the inherent measurement noise or biological variability. In the context of myocardial modeling, where even adjacent samples can differ markedly in stiffness and fiber orientation (Dokos et al., 2002; Nielles-Vallespin et al., 2017), such deterministic inference may be misleading (Hauseux et al., 2018). Recent work has begun to address this gap through Bayesian approaches to inverse modeling, which offer not only best-fit estimates but also posterior distributions reflecting uncertainty in data and model form (Ozturk et al., 2021; Joshi et al., 2022; Aggarwal et al., 2023; Peshave et al., 2024; Anton et al., 2024; Taç et al., 2024; Linka et al., 2025). Within the unsupervised setting, Bayesian-EUCLID (Joshi et al., 2022) forms an interesting contribution towards the inference of myocardial tissue behavior from full-field displacement fields and force measurements. Joshi et al. demonstrated the feasibility of using Gibbs sampling for full-field inference in two-dimensional problems, where the material response could be inferred from a linear system of equations for the unknown material parameters in a rich constitutive feature library. In contrast, the setting of this study involves two distinct and compounding challenges: first, the underlying constitutive models are highly nonlinear (Costa et al., 2001; Holzapfel and Ogden, 2009; Martonová et al., 2024), with strongly cross-correlated parameters (Guan et al., 2018; Laita et al., 2025) that give rise to complex posterior landscapes; and second, the transition to three-dimensional full-field data. While the latter provides richer information, it also substantially increases the computational cost of evaluating the likelihood function. These factors independently and jointly limit the practical utility of Markov Chain Monte Carlo (MCMC) methods, including Gibbs sampling, due to the high cost of repeated likelihood evaluations and the poor mixing behavior in high-dimensional, correlated spaces. These limitations motivate us to extend the Bayesian EUCLID framework with more scalable inference strategies, such as Stochastic Variational Inference (SVI), which approximates posterior distributions through optimization rather than sampling, enabling tractable inference in settings with both complex models and large data volumes.
Towards these goals, this study presents an unsupervised, full-field, Bayesian framework for the inference of orthotropic hyperelastic material parameters from a single heterogeneous biaxial tensile test. Our approach builds on the Bayesian-EUCLID framework (Joshi et al., 2022), extended via SVI (Thorat et al., 2025) to capture parametric uncertainty in a highly nonlinear, and a-priori chosen, constitutive model form. The myocardium is modeled using the well-established Holzapfel–Ogden constitutive law (Holzapfel and Ogden, 2009; Avazmohammadi et al., 2019), and heterogeneity in the deformation field is introduced either through microstructural variation (i.e. intrinsic fiber-sheet-interlaminar architectural heterogeneity), geometric tissue sample modification, or both. We demonstrate that the spatial richness of the heterogeneous displacement field combined with reaction force data enables robust parameter inference with uncertainty quantification from a single test. We additionally validate the inferred material models using unseen invariant values and assess sensitivity to displacement noise.
This work highlights the potential of full-field, unsupervised, and uncertainty-aware approaches to redefine how we efficiently characterize complex living tissues, taking into account minimal tissue availability and minimizing tissue manipulation.
2 Unsupervised Stochastic Variational constitutive parameter inference
Conventional approaches for characterizing myocardial material behavior typically rely on spatially averaged measurements and simplifying assumptions of spatially homogeneous microstructural organization. However, this intrinsic homogenization obscures the true spatial heterogeneity present within a tissue sample. To address this, we infer the constitutive parameter vector in an unsupervised setting from experimentally accessible data: (i) readily available global reaction forces, (ii) full-field deformation, and (iii) the microstructural organization across the slab. In this work, we assume digital volume correlation (DVC) data derived from concomitant through-thickness imaging, such as ultrasound (Navy et al., 2025), Magnetic Resonance Imaging (Estrada et al., 2020), or X-Ray Imaging (Davis et al., 2024) to obtain three-dimensional deformation fields. Microstructural maps in (iii) can be obtained via SHG or micro-CT (Alberini et al., 2024; Kakaletsis et al., 2021; Maes et al., 2022). Among established inverse strategies such as Finite Element Model Updating (FEMU) (Meijer et al., 1999), the Virtual Fields Method (VFM) (Grédiac et al., 2006), the Equilibrium Gap Method (EGM) (Claire et al., 2004; Roux and Hild, 2020), and Efficient Unsupervised Unsupervised Constitutive Law Identification (EUCLID) (Flaschel et al., 2021), we build on the latter. EUCLID formulates the inverse problem via the weak form of momentum balance (see Section 2.1). Here, we extend it to a fully three-dimensional setting with nonlinear orthotropic materials, and recast the classic deterministic parameter optimization into an unsupervised Bayesian inference problem. Leveraging Stochastic Variational Inference (SVI), we identify orthotropic hyperelastic parameters from a single biaxial test; the overall workflow is summarized in Figure 1 and detailed in Section 2.2.
Figure 1: Schematic overview of our unsupervised full-field stochastic variational inference framework for orthotropic hyperelastic tissue behavior. Quantification of the tissue’s microstructural organization, such as spatially varying fiber orientations, provides essential architectural information across the sample domain. In parallel, point-wise measurements of displacements $\hat{\bm{u}}$ and reaction forces $\hat{\bm{R}}^{\beta}$ are acquired under quasi-static biaxial loading using digital volume correlation. The sample geometry is discretized into a finite element mesh, enabling reconstruction of continuous deformation fields $\bm{F}$ from the measured data. A nonlinear orthotropic constitutive model $\psi$, parameterized by a set of cross correlated material parameters $\bm{\theta}$, serves as the mechanistic basis for inference. Given the reconstructed deformation and microstructural organization fields, the model predicts stress responses at the element level, which are used to compute internal and external nodal forces. Residuals are formulated through the weak form of the conservation of linear momentum, minimized pointwise for unconstrained degrees of freedom and in aggregate at boundaries with known reaction forces. The inverse problem is cast as a stochastic variational inference task, where a variational guide distribution $q(\bm{\theta})$ approximates the true posterior $\pi\left(\bm{\theta}\,|\,\hat{\bm{u}},\hat{\bm{R}}^{\beta}\right)$ by minimizing the Kullback-Leibler divergence $\mathrm{KL}\left(\pi\left(\bm{\theta}\,|\,\hat{\bm{u}},\hat{\bm{R}}^{\beta}\right)\,||\,q\left(\bm{\theta}\right)\right)$. Our physics-informed constitutive inference framework enables efficient and scalable inference from complex high dimensional full-field deformation datasets, accommodating nonlinear material behaviors and cross-correlated parameter effects.
2.1 EUCLID preliminaries
Consider a myocardial tissue specimen undergoing quasi-static deformation in a three-dimensional reference domain $\Omega\subset\mathbb{R}^{3}$. Dependent on the tissue slab’s extraction orientation protocol (Fig. 3), the underlying myocardial microstruture variation induces diverse and heterogeneous strain states, while the material itself is assumed to be homogeneous. Optionally, additional tissue manipulations, e.g. cutting a hole in the tissue slab or local cardiac ablation, can further promote a highly heterogeneous deformation state during biaxial tensile testing.
Boundary conditions are applied such that Dirichlet conditions are enforced on $\partial\Omega_{u}\subset\partial\Omega$ and Neumann conditions on the remainder, $\partial\Omega_{t}=\partial\Omega\setminus\partial\Omega_{u}$. For simplicity, our analysis focuses on displacement-controlled loading (i.e., Dirichlet boundary conditions), while noting that applied forces in load-controlled scenarios are equivalent to reaction forces under displacement control. The dataset comprises of $n_{t}$ snapshots of displacement measurements,
$\mathcal{U}=\left\{\bm{u}^{a,t}\in\mathbb{R}^{3}:a=1,\dots,n_{n};\ t=1,\dots,n_{t}\right\},$ (1)
recorded at $n_{n}$ reference points
$\mathcal{X}=\left\{\bm{X}^{a}\in\Omega:a=1,\dots,n_{n}\right\}.$ (2)
Additionally, for each snapshot, $n_{\beta}$ reaction forces
$\left\{R^{\beta,t}:\beta=1,\dots,n_{\beta};\ t=1,\dots,n_{t}\right\}$ (3)
are measured at selected Dirichlet boundaries (e.g., using load cells). Since $n_{\beta}\ll n_{n}$, the force data is sparse compared to the dense displacement field. For brevity, the superscript $(\cdot)^{t}$ is omitted in the subsequent discussion, although the numerical procedure is applied independently to every snapshot. Given these limited measurements, the primary goal is to infer the a priori assumed constitutive model $\psi(\bm{F},\bm{\theta})$’s parameter set $\bm{\theta}$ governing the stress–strain response.
To reconstruct the displacement field, the reference domain $\mathcal{X}$ is discretized using linear tetrahedral elements (each with a single quadrature point at its barycenter),
$\bm{u}(\bm{X})=\sum_{a=1}^{n_{n}}N^{a}(\bm{X})\ \bm{u}^{a}.$ (4)
Here, $N^{a}:\Omega\to\mathbb{R}$ represents the shape function associated with node $\bm{X}^{a}$. We subsequently approximate the corresponding deformation gradient field as
$\bm{F}(\bm{X})=\bm{I}+\sum_{a=1}^{n_{n}}\bm{u}^{a}\otimes\nabla N^{a}(\bm{X}),$ (5)
where $\bm{I}$ is the identity matrix, and $\nabla$ represents the gradient operator with respect to the reference coordinates $\bm{X}$.
Following Flaschel et al. (2021) and Thakolkaran et al. (2022), we leverage the conservation of linear momentum to guide the learning of the constitutive model parameter set, which eliminates the need of stress labels. Assuming quasi-static loading conditions and negligible body forces, the weak form of the linear momentum balance in our reference domain $\Omega$ is given by
$\int_{\Omega}\bm{P}(\bm{F}):\nabla\bm{v}\,\mathrm{d}V-\int_{\partial\Omega_{t}}\hat{\bm{t}}\cdot\bm{v}\,\mathrm{d}S=0\quad\forall\ \text{admissible }\bm{v}\,$, (6)
where $\hat{\bm{t}}$ denotes the prescribed traction acting on $\partial\Omega_{t}$ and $\bm{v}$ is a test function that is sufficiently regular and vanishes on the Dirichlet boundary $\partial\Omega_{u}$. We prefer the weak formulation over the strong form, as it avoids the need for second order derivatives with respect to the displacement-field, which are sensitive to measurement noise.
Assuming no dissipative energy losses within the material and rewriting the Clausius–Duhem entropy inequality Planck (1897) following the Coleman and Noll principle Coleman and Noll (1959); Peirlinck et al. (2024), we get
$\bm{P}(\bm{F})=\frac{\partial\psi(\bm{F},\bm{\theta})}{\partial\bm{F}}$ (7)
as the constitutive relation between the first Piola-Kirchoff stress $\bm{P}$ and the deformation gradient $\bm{F}$ defined through the free energy function $\psi$ and parametrized using the constitutive parameter set $\bm{\theta}$.
Let $\mathcal{D}=\{(a,i):a=1,\dots,n_{n};\ i=1,2,3\}$ denote all the displacement degrees of freedom in our domain $\Omega$, which we partition into $\mathcal{D}^{\text{free}}$, the unconstrained degrees of freedom, and $\mathcal{D}^{\text{fix}}_{\beta}$ (with $\beta=1,\dots,n_{\beta}$) denoting the degrees of freedom under Dirichlet constraints that contribute to the observed reaction force $R^{\beta}$.
Approximating the test function $\bm{v}$ by
$\bm{v}(\bm{X})=\sum_{a=1}^{n_{n}}N^{a}(\bm{X})\,\bm{v}^{a},\qquad\text{with}\quad v_{i}^{a}=0\quad\forall\ (a,i)\in\bigcup_{\beta=1}^{n_{\beta}}\mathcal{D}^{\text{fix}}_{\beta}\,$, (8)
and introducing our constitutive relation (7), our weak form (6) reduces to
$\sum_{a=1}^{n_{n}}v_{i}^{a}f_{i}^{a}=0\,,\quad\text{with}\quad f_{i}^{a}=\underbrace{\int_{\Omega}P_{ij}\,\nabla_{j}N^{a}\,\mathrm{d}V}_{\text{internal force}}-\underbrace{\int_{\partial\Omega_{t}}\hat{t}_{i}N^{a}\,\mathrm{d}S}_{\text{external force}}\,.$ (9)
Here, the integrals are computed using numerical quadrature over the same discretization and mesh as (4) resulting in a Bubnov-Galerkin form. Since the test functions are arbitrary, the force residual must vanish at every unconstrained degree of freedom,
$f_{i}^{a}=0\quad\forall\ (a,i)\in\mathcal{D}^{\text{free}}\,.$ (10)
At the fixed degrees of freedom, the internal and external forces are counteracted by the reaction force imposed by the Dirichlet constraints. Since point-wise reaction forces cannot be measured experimentally, they are assumed to be unavailable. Instead, only the total reaction forces integrated over the boundary segments are known. As a result, the global force balance for each measured reaction force is expressed as
$\sum_{(a,i)\in\mathcal{D}^{\text{fix}}_{\beta}}f^{a}_{i}=R^{\beta}\qquad\forall\qquad\beta=1,\dots,n_{\beta},$ (11)
where the summation is carried out over all point-wise forces associated with the degrees of freedom in the $\beta^{\text{th}}$ Dirichlet constraint, denoted as $\mathcal{D}^{\text{fix}}_{\beta}$. As noted earlier, the superscript $(\cdot)^{t}$ has been omitted for brevity, but the above force balance conditions apply to all data snapshots at $t=1,\dots,n_{t}$.
Together, (10) and (11) constitute a system of equations, which defines the inverse problem of constitutive model inference (Flaschel et al., 2021; Thakolkaran et al., 2022, 2025) as the minimization of the force balance residuals with respect to the trainable parameter set $\bm{\theta}$ of the free energy function $\psi_{\bm{}}$ (7):
$\bm{\hat{\theta}}\leftarrow\arg\min_{\bm{\theta}}\sum_{t=1}^{n_{t}}\Bigg[\underbrace{\sum_{(a,i)\ \in\ \mathcal{D}^{\text{free}}}\left(f_{i}^{a,t}\right)^{2}}_{\text{free degrees of freedom}}+\underbrace{\sum_{\beta=1}^{n_{\beta}}\Bigl(R^{\beta,t}-\sum_{(a,i)\ \in\ \mathcal{D}^{\text{fix}}_{\beta}}f_{i}^{a,t}\Bigr)^{2}}_{\text{fixed degrees of freedom}}\Bigg]\,$, (12)
such that the displacement and reaction force data satisfy the physics-based constraints (10) and (11).
2.2 Stochastic variational inference
In this study, we assume an a priori known constitutive law (see Eq. 27) and aim to stochastically infer the underlying parameter set $\bm{\theta}$. Leveraging Bayes’ theorem, we express both our prior beliefs about the parameter set and the influence of noisy measurements within a unified probabilistic framework.
We recast the deterministic optimization problem in (12) into Bayesian inference by treating the force residuals as the result of additive measurement noise (both for displacements and reaction forces). Let $\bm{\rho}(\bm{\theta})$ denote the global residual obtained by concatenating contributions from all load steps $t=1,\ldots,n_{t}$; for brevity, we omit explicit superscripts $(\cdot)^{t}$ when the meaning is clear from context. The residual stacks two components: (a) the internal-force equilibrium at unconstrained degrees of freedom and (b) the reaction-force consistency at fixed, i.e. instrumented, boundary segments:
$\bm{\rho}(\bm{\hat{\theta}})=\begin{bmatrix}\{f_{i}^{a}\}_{(a,i)\in\mathcal{D}^{\text{free}}}\\ \{R^{\beta}-\sum_{(a,i)\in\mathcal{D}^{\text{fix}}_{\beta}}f_{i}^{a}\}_{\beta=1}^{n_{\beta}}\end{bmatrix}$ (13)
Note that $\bm{\rho}(\bm{\theta})$ implicitly depends on observed displacements through the predicted forces in $f_{i}^{a}$, and thus the material model parameters $\bm{\theta}$. We assume that these residuals are distributed according to a multivariate normal distribution with zero mean and isotropic variance, reflecting the assumption of additive Gaussian noise:
$\bm{\rho}(\bm{\hat{\theta}})\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})$ (14)
In our case, both force and displacement measurements are affected by sensor noise. However, because displacements propagate in nonlinear fashion to the force residuals, the uncertainty introduced by the displacement noise is assumed to dominate. Under this assumption, the likelihood of the observed displacement $\hat{\bm{u}}$ and reaction force $\hat{R}^{\beta}$ data, given the parameters $\bm{\theta}$ and noise variance $\sigma^{2}$, becomes:
$\pi\left(\hat{\bm{u}},\hat{R}^{\beta}\mid\bm{\theta},\sigma^{2}\right)\,\propto\,\exp\left(-\frac{1}{2\sigma^{2}}\|\,\bm{\rho}(\bm{\theta})\,\|^{2}\right)$ (15)
Again, for brevity, the superscript $(\cdot)^{t}$ is omitted in the subsequent discussion, although the likelihood distribution is computed as the product of the likelihood distribution for the residual computed at each independent quasi-static loading time snapshot.
Next, we place priors on the the noise variance $\sigma^{2}$ and the parameter vector $\bm{\theta}$ respectively to complete our Bayesian treatment of the inference problem. Because the variance $\sigma^{2}$ of the measurement noise is unknown and must be inferred along with the parameters, we place an inverse-gamma prior on it:
$\sigma^{2}\sim\mathcal{IG}\,\left(\alpha_{\sigma},\beta_{\sigma}\right)$ (16)
Concomitantly, we enforce physical constraints on the parameters by requiring them to be positive and therefore enforce a truncated normal distribution for the prior. Here, we introduce a scaling hyperprior $V_{s}$ which allows us to decouple the overall measurement residual noise variance $\sigma$ from the underlying parameter variance.
$\bm{\theta}\sim\mathcal{N}_{+}\,\left(\bm{0},v_{s}\,\sigma^{2}\bm{I}\right)$ (17)
For this scaling hyperprior, we again enforce parameter positivity by setting an inverse gamma prior distribution:
$v_{s}\sim\mathcal{IG}\,\left(\alpha_{v_{s}},\beta_{v_{s}}\right)$ (18)
Combining the prior distributions (16) – (18) with our likelihood function (15) yields the posterior:
$\pi\left(\bm{\theta},\sigma^{2},v_{s}\mid\hat{\bm{u}},\hat{R}^{\beta}\right)\,\propto\,\pi\left(\hat{\bm{u}},\hat{R}^{\beta}\mid\bm{\theta},\sigma^{2}\right)\,\pi\left(\bm{\theta}\mid\sigma^{2},v_{s}\right)\,\pi\left(\sigma^{2}\right)\,\pi\left(v_{s}\right),$ (19)
This posterior serves as the foundation for either sampling-based inference or variational inference approaches, chosen according to model complexity and data volume. In simpler inference problems, sampling-based methods such as MCMC can often be relied on to approximate the posterior distribution. For example, Gibbs sampling has been successfully applied to two-dimensional full-field problems, where the material response could be inferred by solving a linear system of equations on a library of constitutive features (Joshi et al., 2022). In those cases, the constitutive models were potentially nonlinear in input but remained linear in their parameters, which keeps the posterior geometry relatively tractable for MCMC. In our setting, however, the constitutive models are nonlinear in both inputs and parameters, which induces strong parameter interactions and complex (often multimodal) posteriors. Together with the cost of evaluating data-heavy 3D full-field likelihoods, this makes traditional MCMC prohibitively expensive. In turn, we use variational inference to approximate the posterior distribution $\pi$ with a more tractable family of parametric guide distributions $q$. We therefore adopt stochastic variational inference (SVI) to approximate the posterior efficiently, enabling gradient-based optimization such as stochastic gradient descent and ADAM (Kingma and Welling, 2013; Kingma and Ba, 2017). SVI has been shown to substantially accelerate inference relative to MCMC (Rodrigues, 2022; Thorat et al., 2025). Additionally, we employ a mean-field variational family $q$ (Parisi, 1988; Consonni and Marin, 2007), which assumes statistical independence between the quantities of interest.
$q\left(\bm{\theta},\,v_{s},\,\sigma^{2}\right)=q\left(\bm{\theta}\right)\,q\left(\sigma^{2}\right)\,q\left(v_{s}\right)$ (20)
where each factor is defined as:
$q\left(\bm{\theta}\right)\quad$ $=$ $\quad\mathcal{N}_{+}(\bm{\mu}^{q},\bm{\Sigma}^{q}\circ\bm{I})$ (21)
$q\left(\sigma^{2}\right)\quad$ $=$ $\quad\mathcal{IG}(\alpha_{\sigma}^{q},\beta_{\sigma}^{q})$ (22)
$q\left(v_{s}\right)\quad$ $=$ $\quad\mathcal{IG}(\alpha_{vs}^{q},\beta_{vs}^{q})$ (23)
and $\circ$ denotes the element-wise (Hadamard) product. The full set of variational parameters is denoted by $\bm{\Phi}$ and our hyperparameters are reported in Table 1. The hierarchical structure (Joshi et al., 2022; Thorat et al., 2025; Peirlinck et al., 2019) of our prior distributions is showcased in Figure 2.
Figure 2: Schematic of the hierarchical inference model. This hierarchical structure of the Bayesian inference model allows for simulteneous placement of prior distributions on the noise in the likelihood and the constitutive model parameters. The red and circular distributions represent hyperpriors for $V_{s}$ and $\sigma^{2}$, the lightblue circular distribution represent the prior on the material model parameters and the black and rectangular distribution represents the likelihood of the linear momentum balance. The black arrows denote the hierarchical structure dependencies. The hyperpriors are distributed according to inverse-gamma distributions. The model parameters are distributed according to a truncated multivariate normal distribution. Truncation occurs at numerical zero to enforce polyconvexity in the material model. The likelihood is assumed to be normally distributed.
To find the best approximation within this family, we minimize the Kullback-Leibler (KL) divergence (Kullback, 1968; Cover T., 1968) between the variational distribution $q\left(\bm{\Phi}\right)$ and the true posterior $\pi\left(\bm{\Phi}\mid\hat{\bm{u}},\hat{R}^{\beta}\right)$, where $\bm{\Phi}=\{\bm{\theta},v_{s},\sigma^{2}\}$:
$\mathrm{KL}\left[\pi(\bm{\Phi}\mid\hat{\bm{u}},\hat{R}^{\beta})\mid\mid q(\bm{\Phi})\right]=\mathbb{E}_{q(\bm{\Phi})}\left[-\log q\left(\bm{\Phi}\right)-\log\pi\left(\bm{\Phi}\right)-\log\pi\left(\hat{\bm{u}},\hat{R}^{\beta}\mid\bm{\Phi}\right)\right]+\log\pi\left(\hat{\bm{u}},\hat{R}^{\beta}\right)$ (24)
The entire inference pipeline is implemented using NumPyro (Phan et al., 2019), which supports automatic differentiation and stochastic optimization with ADAM using the reparameterization trick (Kingma and Ba, 2017; Kingma and Welling, 2013), which alleviates the need to explicitly define the minimization procedure as was done in (Thorat et al., 2025). We initialize the optimization method with a learning rate of $r_{0}=0.01$ and use exponential learning rate decay $\gamma$. The learning rate is defined for each epoch $i$ as $r_{i}=r_{0}\gamma^{i/n_{T}}$.
Table 1: Hyperparameters and simulation settings used in the stochastic variational inference framework and synthetic tissue slab data generation. The upper rows list optimization settings and hierarchical Bayesian prior parameters used for the SVI guide distributions and model priors. The bottom rows summarize the geometry and loading conditions for the training and validation slab specimens, respectively.
Parameter Notation Value
SVI hyperparameters:
  weight for reaction force balance $\lambda_{r}$ $10$
  Optimizer - Adam
  Number of epochs $n_{T}$ 100,000100{,}000
  Learning rate scheduler - exponential
  Base learning rate $r_{0}$ $0.01$
  Learning decay rate $\gamma$ $0.1$
SVI guide hyperparameters:
  Initial mean of $\mathcal{N}_{+}$ distribution $\mu_{\theta}$ $10+\epsilon^{2}\times\{1,\dots,1\},\quad\epsilon\sim\mathcal{N}(0,1)$
  Initial variance of $\mathcal{N}_{+}$ distribution $\sigma^{2}_{\theta}$ $10^{-18}\times\{1,\dots,1\}$
  Initial scale parameter for $v_{s}$ $\alpha_{vs}$ $10^{4}$
  Initial rate parameter for $v_{s}$ $\beta_{vs}$ $10^{4}$
  Initial scale parameter for $\sigma^{2}$ $a_{\sigma^{2}}$ $10^{4}$
  Initial rate parameter for $\sigma^{2}$ $b_{\sigma^{2}}$ $10^{4}$
SVI model prior hyperparameters:
  prior mean of $\mathcal{N}_{+}$ distribution $\mu_{\theta}$ $10^{-10}\times\{1,\dots,1\}$
  prior variance of $\mathcal{N}_{+}$ distribution $\sigma^{2}_{\theta}$ $\{1,\dots,1\}$
  prior scale parameter for $v_{s}$ $\alpha_{vs}$ 11
  prior rate parameter for $v_{s}$ $\beta_{vs}$ 11
  prior scale parameter for $\sigma^{2}$ $\alpha_{\sigma^{2}}$ 11
  prior rate parameter for $\sigma^{2}$ $\beta_{\sigma^{2}}$ 11
Training slab specimen:
  Number of nodes in mesh for FEM-based data generation - $16,856$
  Number of reaction force constraints $n_{\beta}$ 55
  Number of data snapshots $n_{t}$ 2
  Loading parameter $\delta$ $\{0.5\times t:t=2,3\}$
  Loading ratios - $\{(1:2),(1:1),(2:1)\}$
Ground-truth material model parameters $\bm{\theta}_{gt}$:
  Linear scaling parameters $a_{\star}$ $[kPa]$ $a,a_{f},a_{n},a_{fs}$ $\{0.809,1.911,0.227,0.547\}$
  Exponential scaling parameters $_{\star} [][-] b,b$$b_{f},b_{n},b_{fs}$ $\{7.474,22.063,34.802,5.691\}$
  Bulk modulus $[MPa]$ $K$ $0.1$
3 Synthetic data generation
To assess our Bayesian inference framework under controlled conditions, we generate synthetic biaxial tensile testing data from high-resolution simulations of varying microstructurally and geometrically heterogeneous myocardial tissue slabs.
3.1 Myofiber architecture and tissue slab slicing
Informed by prior histological characterizations of myocardial tissue, we model our studied specimens as a layered orthotropic material in which cardiomyocytes align along main myofiber directions, packed by endomysial collagen into sheets, and connected via perimysial collagen (Rohmer et al., 2007; Holzapfel and Ogden, 2009; Alexander J. Wilson, 2022). This organization defines three locally orthogonal directions: the fiber direction $\bm{f}$, the sheet direction $\bm{s}$, and the sheet-normal direction $\bm{n}=\bm{f}\times\bm{s}$. To investigate the role of heterogeneity in unsupervised constitutive parameter identifiability, we simulate myocardial tissue slabs that exhibit both microstructure and geometry induced heterogeneity.
Figure 3: Specimen slicing and microstructural organization. Top: Schematic of the left ventricle and three considered slicing orientations to extract myocardial tissue slabs: circumferential–longitudinal (circ-long), circumferential–radial (circ-rad), and a rotated variant of the latter (rad-heli). Bottom: Each slicing orientation leads to a distinct microstructural configuration within the specimen. In circ-long slabs, the fiber direction lies in-plane and the sheet direction spans the through-thickness. Circ-rad slabs display a transmural fiber rotation from epicardium to endocardium, with the sheet direction aligned in the radial plane. The rad-heli orientation rotates the slab around the radial axis, aligning all three microstructural directions approximately within the plane of the specimen. Fiber directions ($\bm{f}$) are shown in red, sheet directions ($\bm{s}$) in pink, and normal directions ($\bm{n}$) in brown.
Leveraging the intrinsic spatially varying myofiber architecture in the human heart (Lombaert et al., 2012), we extract three microstructurally distinct synthetic tissue slabs from the left ventricular free wall. All samples are assumed to be $10\times 10\times 1$ mm3 in size, corresponding to representative slabs from a healthy left ventricular wall which is at least 10 mm thick. The sample geometry and applied boundary conditions are also visualized in Figure 1 - left. Figure 3 showcases the microstructural variability across the slab domain, with directional fiber $\bm{f}$, sheet $\bm{s}$ and normal $\bm{n}$ vectors discretized element-wise, depend on how the tissue slicing and extraction procedure is performed:
  • 1.
    circ-long: sections aligned with the ventricular circumferential–longitudinal plane. The in-plane $\bm{f}$ orientation and through-thickness $\bm{s}$ direction enable near-planar fiber alignment. We consider both aligned and off-axis variants depending on their angular deviation from the square edge. This slab specimen type aligns with the biaxial tension testing slabs studied by Sommer et al. (2015).
  • 2.
    circ-rad: sections cut in the circumferential–radial plane. These samples exhibit a transmural fiber rotation from approximately $60^{\circ}$ at the epicardium to $60^{\circ}$ at the endocardium, while the sheet direction $\bm{s}$ remains in-plane and radially aligned.
  • 3.
    rad-heli: samples based on a radial-helical cut. Given the natural transmural variation of the myofiber angle $\bm{f}$ from $60^{\circ}$ in the sub-epicardium to $60^{\circ}$ in the sub-endocardium (Lombaert et al., 2012), this configuration places all three microstructural directions ($\bm{f},\bm{s},\bm{n}$) approximately within the sample plane by rotating the sample an additional $45^{\circ}$ along the radial axis. The resultant fiber orientation ranges from $15^{\circ}$ up to $105^{\circ}$, aligning the $\bm{f}$ fiber in the sample plane near the bottom edge of the sample and aligning the $\bm{n}$ fiber in the sample plane near the top edge.
To investigate the influence of geometric heterogeneity, we introduce a second set of specimens with a central circular occlusion with a radius of $1$ mm, see Figure 1 - right. These heterogeneous samples introduce additional strain concentration zones that we hypothesize may improve the identifiability of model parameters under limited data conditions, e.g. a biaxial tensile test to characterize orthotropic material behavior.
3.2 Biaxial loading emulation
We simulate a digital volume correlation experiment by modeling a biaxial loading protocol on the microstructurally and geometrically homogeneous and heterogeneous tissue slabs described above using the finite element method (FEM) implemented in ABAQUS/Standard (Dassault Systèmes Simulia Corp., 2025), see for example (Peirlinck et al., 2018). Our virtual tissue slabs are subjected to displacement-controlled symmetric biaxial loading, with loading ratios $\lambda_{1}:\lambda_{2}=(1:2),(1:1),(2:1)$ based on stretch. The maximum stretch applied to the samples is $15\%$ their initial length, which relates to maximal values used in experiments (Sommer et al., 2015). We use linear tetrahedral elements to discretize the domain and record nodal displacements along with reaction forces for a sequence of $n_{t}$ load steps (see Figure 1)
Leveraging the local volume change of deformation $J=\det\left(\bm{F}\right)$, we multiplicatively decompose the deformation gradient $\bm{F}$ (5) into its volumetric $\bm{F}^{\rm{vol}}$ and isochoric $\bar{\bm{F}}$ parts (Flory, 1961),
$\bm{F}=\bm{F}^{\rm{vol}}\bar{\bm{F}}\quad\mbox{with}\quad\bm{F}^{\rm{vol}}=J^{\frac{1}{3}}\bm{I}\quad\mbox{and}\quad\bar{\bm{F}}=J^{-\frac{1}{3}}\bm{F}.$ (25)
and introduce the following deformation (pseudo-)invariants (Spencer, 1984; Menzel, 2004):
$\bar{I}_{1}$ $=$ $[\bar{\bm{F}}^{{\rm{T}}{}}\bar{\bm{F}}]:\bm{I}$ (26)
$\bar{I}_{\rm{4ff}}$ $=$ $[\bar{\bm{F}}^{{\rm{T}}{}}\bar{\bm{F}}]:[\bm{f}^{0}\otimes\bm{f}^{0}]$
$\bar{I}_{\rm{4nn}}$ $=$ $[\bar{\bm{F}}^{{\rm{T}}{}}\bar{\bm{F}}]:[\bm{n}^{0}\otimes\bm{n}^{0}]$
$\bar{I}_{\rm{4fs}}$ $=$ $[\bar{\bm{F}}^{{\rm{T}}{}}\bar{\bm{F}}]:[\bm{f}^{0}\otimes\bm{s}^{0}]$
where $\bm{f}^{0}$, $\bm{s}^{0}$, $\bm{n}^{0}$ denote the local microstructure fiber, sheet, and normal unit vector orientations in the undeformed reference configuration. Sometimes these invariants are centered around zero, in which case the following notation is used: $K_{1}=I_{1}-3,K_{4ff}=I_{4ff}-1,K_{4nn}=I_{4nn}-1$. We adopt the orthotropic, compressible Holzapfel-Ogden (HO) material model (Holzapfel and Ogden, 2009; Guan et al., 2018) using the following strain energy density function:
$\psi(\bm{F},\bm{\theta})$ $=\frac{a}{2b}\left[\exp\left(b(\bar{I}_{1}-3)\right)-1\right]+\sum_{i\,=\,f,n}\frac{a_{i}}{2b_{i}}\left[\exp\left(b_{i}\left<\bar{I}_{4ii}-1\right>^{2}\right)-1\right]$ (27)
$+\frac{a_{fs}}{2b_{fs}}\left[\exp\left(b_{fs}\bar{I}_{4fs}^{2}\right)-1\right]+\frac{K}{2}\left(\frac{J^{2}-1}{2}-\ln J\right)$
to compute the First Piola Kirchoff stress tensor $\bm{P}$ following (7). This variation of the HO-model was implemented using the universal material subroutine (Peirlinck et al., 2024). An analytical form of the stress tensor $\bm{P}$ is provided in B.
In (27), $\bm{\theta}=\{a,b,a_{f},b_{f},a_{n},b_{n},a_{fs},b_{fs},K\}$ (see Table 1) denotes the to-be-inferred ground-truth material parameters. The Macaulay brackets $\langle x\rangle$ ensure that fiber and sheet term strain energy contributions activate only under tension (Wittrick, 1965; Wriggers et al., 2016; Peirlinck et al., 2024). Lastly, a bulk modulus is chosen based on the distribution of $J$ in the circ-long-hom samples and common compressibility considerations of the myocardium (McEvoy et al., 2018; Yin et al., 1996).
Figure 4: Deformation profiles of geometrically homogeneous tissue slabs with varying microstructural heterogeneity. Maximum principal stretches are shown for circ-long-hom (top row), circ-rad-hom (middle row), and rad-heli-hom (bottom row) specimens subjected to biaxial loading protocols ($\lambda_{1}:\lambda_{2}=(1:2),(1:1),(2:1)$) up to $\lambda^{\text{max}}=1.15$. Rightmost columns display distributions of invariants $I_{1}-3$, $I_{4ff}-1$, $I_{4nn}-1$, and $I_{4fs}$ for each specimen type respectively using approximated kernel density estimations denoted as $\hat{\pi}(I_{*})$. The kernel density estimation is truncated at an arbitrary level of 20 to show the spread of the invariants more clearly.
Figure 4 presents the deformation response of geometrically homogeneous slabs with varying microstructural architectures. The qualitative results of the maximum principal stretch are shown for all loading conditions and homogeneous tissue samples. On the right side of the figure we show kernel density estimations of the biaxially activated invariants - truncated at $20$ to visualize the spread of these invariants. The circ-long-hom configuration represents a classic homogenized myocardial sample, as excised by (Sommer et al., 2015), with assumed uniform fiber-sheet-normal orientations. Under biaxial loading, this results in a uniform deformation pattern with isotropic invariants $I_{1}$ ranging from $3.158$ to $3.257$ across the three loading protocols. Anisotropic contributions are moderate, with $I_{4ff}$ and $I_{4nn}$ spanning ranges from $1.025$ to $1.139$ and $1.024$ to $1.144$ respectively. Fiber-sheet shear contributions remain negligible, with $I_{4fs}$ limited to 0 up to machine precision. Introducing transmural fiber variation in the circ-rad-hom slab increases local deformation heterogeneity. The isotropic invariant $I_{1}$ expands to a broader range of $3.160$ to $3.367$, while $I_{4ff}$ and $I_{4nn}$ reach $1.194$ and $1.252$, respectively. Notably, fiber-sheet shear contributions increase, with $I_{4fs}$ reaching $0.099$. In the rad-heli-hom specimen, which introduces a more asymmetric helical fiber distribution, deformation heterogeneity is further amplified. The isotropic invariant $I_{1}$ ranges from $3.147$ to $3.359$, while $I_{4ff}$ increases significantly, spanning up to $1.248$. The $I_{4nn}$ component reaches $1.251$, and fiber-sheet shear contributions grow to $0.114$. The distributions of invariants shown in the right column of Figure 4 emphasize these trends. The circ-long-hom specimen exhibits single homogeneous singular values for $I_{4ff}$ and $I_{4nn}$, while $I_{4fs}$ remains zero. In contrast, circ-rad-hom and rad-heli-hom configurations display broader distributions across all invariants, particularly for fiber-sheet shear deformations. Consistent with these observations, we note tissue bulging effects in the thickness direction. Circ-rad configurations exhibit bulging centrally, whereas rad-heli configurations deform near the specimen edges. In particular, the rad-heli-hom configuration shows noticeable thickness bulging coinciding with bands of reduced strain, where fibers align in-plane. These bands are visually distinct due to their contrasting strain magnitudes relative to surrounding regions. These results underline the capacity of microstructural heterogeneity alone, even in geometrically simple specimens, to enrich deformation diversity. Such diversity is critical for reliably inferring nonlinear orthotropic constitutive behavior, motivating their role as training configurations in this study.
Figure 5: Deformation profiles of geometrically heterogeneous tissue slabs with varying microstructural heterogeneity. Maximum principal stretches circ-long-het (top row), circ-rad-het (middle row), and rad-heli-het (bottom row) specimens under biaxial loading ($\lambda_{1}:\lambda_{2}=(1:2),(1:1),(2:1)$). Invariant distributions of $I_{1}-3$, $\bar{I}_{4ff}-1$, $\bar{I}_{4nn}-1$, and $I_{4fs}$ are shown in the rightmost panels using approximated kernel density estimations denoted as $\hat{\pi}(I_{*})$. The kernel density estimation is truncated at an arbitrary level of 20 to show the spread of the invariants more clearly.
Figure 5 illustrates the deformation response of geometrically heterogeneous slabs, where a central occlusion is combined with varying microstructural organization. In the circ-long-het specimen, geometric heterogeneity alone induces significant deformation diversity. The isotropic invariant $I_{1}$ spans from $3.070$ to $3.390$, with $I_{4ff}$ and $I_{4nn}$ reaching up to $1.232$ and $1.209$, respectively. Fiber-sheet shear contributions, represented by $I_{4fs}$, remain limited, peaking at $0.034$. Adding microstructural heterogeneity in the circ-rad-het configuration further amplifies deformation heterogeneity. Here, $I_{1}$ varies between $3.141$ and $3.5477$, while $I_{4ff}$ and $I_{4nn}$ attain maximum principal stretches of $1.218$ and $1.243$ respectively. Notably, fiber-sheet shear contributions increase substantially, with $I_{4fs}$ rising to $0.4342$. In the rad-heli-het specimen, characterized by asymmetric helical fiber variation and geometric heterogeneity, isotropic $I_{1}$ deformation spans $3.136$ to $3.518$. Anisotropic contributions reach $1.242$ for $I_{4ff}$, $1.241$ for $I_{4nn}$, and $0.289$ for the fiber-sheet shear term $I_{4fs}$, which confirms this configuration as the most mechanically diverse. The distributions of invariants reflect these differences. The circ-long-het slab shows moderate spreading compared to its homogeneous counterpart, but fiber-sheet shearing remains limited. In contrast, the circ-rad-het and rad-heli-het configurations exhibit pronounced broadening in all invariants, particularly in $I_{4fs}$, underscoring the synergistic effect of combined geometrical and microstructural heterogeneities. Similar to the homogeneous cases, bulging effects are evident. Circ-rad configurations exhibit central bulging, while rad-heli configurations deform near specimen edges. The rad-heli-het specimen shows the most pronounced bulging, exceeding initial thickness. Localized deformation bands with reduced strain are again visible, while regions surrounding the central occlusion display elevated strains, highlighting the amplification of local deformation due to geometric discontinuities. These observations reinforce the rationale for using such heterogeneous configurations as training specimens, enhancing parameter identifiability through enriched mechanical diversity.
4 Orthotropic constitutive parameter inference
[…]
$e_{\theta}$ $=\frac{|\hat{\theta}-\theta_{\text{gt}}|}{\theta_{\text{gt}}}\qquad$ $\tilde{\theta}$ $=\frac{\hat{\theta}-\mu_{\hat{\theta}}}{\mu_{\hat{\theta}}}\qquad$ $s_{\tilde{\theta}}$ $=\frac{\sqrt{N(N-1)}}{N-2}\cdot\frac{\frac{1}{N}\sum_{n=1}^{N}(\tilde{\theta}_{n}-\mu_{\tilde{\theta}})^{3}}{\left(\frac{1}{N}\sum_{n=1}^{N}(\tilde{\theta}_{n}-\mu_{\tilde{\theta}})^{2}\right)^{3/2}}$ (28)
with $N$ equaling the amount of samples drawn from the posterior distribution. To gain insight into the shape, mean, and spread of the approximated posterior distribution, we draw 1,0001{,}000 samples from the inferred guide distribution shown in the bottom rows of Figures 6 and 7 respectively.
Figure 6: Orthotropic constitutive parameter inference from geometrically homogeneous tissue slabs with varying microstructure. From left to right: a circ-long-hom specimen, a circ-rad-hom specimen, and a rad-heli-hom specimen. From top to bottom: (i) evolution of inferred material parameters $\hat{\theta}$ over the epochs, (ii) evolution of relative parameter errors $e_{\theta}$, (iii) final relative errors at final epoch, and (iv) rescaled posterior distributions. Distributions are shown in grayscale when unbiased skewness $|s_{\theta}|<0.5$. Despite microstructural heterogeneity, geometric homogeneity limits identifiability of shear-related parameters ($a_{fs}$, $b_{fs}$), which remain poorly inferred. In contrast, isotropic and primary anisotropic parameters are consistently identified with high accuracy, exhibiting narrow and near-Gaussian posterior distributions. Increasing microstructural heterogeneity improves convergence rates and reduces inference error for shear parameters, though challenges persist for $b_{fs}$, particularly in the rad-heli-hom specimen.
Focusing on geometrically homogeneous samples (Figure 6), we evaluated three configurations with increasing microstructural heterogeneity: fully homogeneous (circ-long-hom), symmetric transmural fiber rotation (circ-rad-hom), and asymmetric helical fiber variation (rad-heli-hom). For the circ-long-hom configuration, all parameters converged rapidly except for those associated with the fiber-sheet shear coupling parameters ($a_{fs}$, $b_{fs}$). Final relative errors were minimal for isotropic parameters ($e_{a}=4.77\times 10^{-5}$, $e_{b}=-2.30\times 10^{-6}$) and remained low for anisotropic terms in $f$ and $n$ directions (i.e., $e_{a_{f}}=-5.50\times 10^{-5}$, $e_{b_{f}}=-1.47\times 10^{-6}$, $e_{a_{n}}=9.66\times 10^{-5}$ , $e_{b_{n}}=1.29\times 10^{-5}$) and the bulk modulus ($e_{k}=2.423\times 10^{-6}$). In contrast, shear-related parameters exhibited substantial errors ($e_{a_{fs}}=-141.7$, $e_{b_{fs}}=0.92$). These trends stabilized within 8,0008{,}000 epochs, except for $b_{fs}$, which remained poorly identified. Introducing symmetric transmural fiber heterogeneity, i.e. in the circ-rad-hom slab, markedly improved fiber-sheet shear parameter identification, reducing relative errors to $e_{a_{fs}}=8.05\times 10^{-3}$ and $e_{b_{fs}}=0.17$, while maintaining accuracy for isotropic and anisotropic terms (see Table 4). Convergence for $b_{fs}$ was slower, requiring up to 30,00030{,}000 epochs. Further increasing fiber and sheet deformation heterogeneity (Figure 4), that is in the rad-heli-hom slab, yielded similar improvements , though inferring $b_{fs}$ remained challenging, with a relative error of $0.374$ and no clear stabilization after 100,000100{,}000 epochs. Other parameters exhibited relative errors below $10^{-4}$, comparable to the circ-rad-hom case (Table 4). Posterior distributions (Figure 6 - bottom row) confirmed these trends. Well-identified parameters displayed narrow, symmetric distributions, with standard deviations several orders of magnitude smaller than their means. In contrast, $a_{fs}$ and $b_{fs}$ exhibited skewed or truncated distributions, particularly in the circ-long-hom configuration, indicating limited identifiability. Notably, skewness magnitudes exceeded $|s_{\theta}|>0.5$ only for $a_{fs}$ and $b_{fs}$, while other inferred parameter distributions remained near-Gaussian. Overall, these results highlight that geometric homogeneity severely limits identifiability of shear-related parameters, while even moderate microstructural heterogeneity substantially improves inference quality.
Figure 7: Orthotropic constitutive parameter inference from geometrically heterogeneous tissue slabs with varying microstructure. From left to right: a circ-long-het specimen, a circ-rad-het specimen, and a rad-heli-het specimen. From top to bottom: (i) evolution of inferred material parameters $\hat{\theta}$ over the epochs, (ii) evolution of relative parameter errors $e_{\theta}$, (iii) final relative errors at final epoch, and (iv) rescaled posterior distributions. Distributions are shown in grayscale when unbiased skewness $|s_{\theta}|<0.5$. Geometric heterogeneity significantly enhances identifiability of shear-related parameters, especially $b_{fs}$, reducing relative errors to sub-percent levels in circ-rad-het and rad-heli-het specimens. Posterior distributions for these parameters become narrower and less skewed compared to homogeneous counterparts. All other parameters maintain high accuracy and exhibit symmetric, low-variance distributions. These results underline the critical role of geometric heterogeneity in resolving parameters sensitive to shear-deformation modes.
Focusing on geometrically heterogeneous samples (Figure 7), we evaluated three configurations with increasing microstructural complexity: fully homogeneous fibers (circ-long-het), symmetric transmural fiber rotation (circ-rad-het), and asymmetric helical fiber variation (rad-heli-het). For the circ-long-het configuration, all parameters were accurately identified except for $b_{fs}$ (relative error of roughly $1\%$). Relative errors remained low for isotropic parameters ($e_{a}=1.83\times 10^{-5}$, $e_{b}=1.22\times 10^{-5}$), anisotropic parameters in $f$ and $n$ directions ($e_{a_{f}}=-4.18\times 10^{-5}$, $e_{b_{f}}=9.77\times 10^{-6}$, $e_{a_{n}}=7.87\times 10^{-5}$, $e_{b_{n}}=-1.40\times 10^{-5}$), and the bulk modulus ($e_{k}=4.84\times 10^{-6}$). However, fiber-sheet shear coupling remained challenging, with errors of $e_{a_{fs}}=1.02\times 10^{-2}$ and $e_{b_{fs}}=0.879$. All parameters stabilized within 8,0008{,}000 epochs. Introducing symmetric fiber heterogeneity in the circ-rad-het slab markedly improved shear parameter inference, reducing $e_{a_{fs}}$ and $e_{b_{fs}}$ to $2.52\times 10^{-3}$ and $1.54\times 10^{-3}$, respectively. Isotropic and anisotropic parameters remained highly accurate (e.g., $e_{a}=-1.64\times 10^{-5}$, $e_{b}=-2.80\times 10^{-5}$), with convergence achieved within 5,0005{,}000 epochs. Further increasing microstructural complexity in the rad-heli-het configuration yielded comparable results. Relative errors for $a_{fs}$ and $b_{fs}$ were $1.14\times 10^{-3}$ and $6.90\times 10^{-3}$, respectively, while other parameters maintained relative errors below $10^{-4}$, as shown in the third row of Figure 7 and Table 4. Convergence required up to 10,00010{,}000 epochs, slightly slower than the circ-rad-het case. Posterior distributions (Figure 7 - bottom row) confirmed these trends. Parameters associated with isotropic and anisotropic invariants displayed narrow, near-Gaussian distributions with rescaled standard deviations between $10^{-3}$ and $10^{-5}$. In contrast, shear-related parameters ($a_{fs}$, $b_{fs}$) exhibited wider, skewed distributions, particularly in circ-long-het. However, skewness and standard deviations were notably reduced compared to homogeneous slab configurations, indicating improved identifiability. In general, these results demonstrate that geometric heterogeneity substantially enhances inference of shear-related parameters, effectively complementing microstructural variations to improve parameter identifiability.
In general, we notice in the top rows of figures 6 and 7 that the parameters $b$ , $b_{f}$, $a_{n}$ and $a_{fs}$ drop in value after roughly 1,0001{,}000 epochs, only to reach a minimal value around 5,0005{,}000 epochs and then shoot up to converge to their optimal values. Similar behavior can be seen for the remaining parameters, though they stabilize at lower values than their initial values. In the second row of these figures we can also see interesting jumps between $10$ and 5,0005{,}000 epochs. Comparing these jumps to the parameter values in the top row reveals that these are the epochs where the estimated parameter value crosses the ground-truth value without achieving a stabilizing solution. This underlines the changing posterior geometry and thus also how the loss landscape changes due to parameter optimization. The summary of all parameter fitting results under noise free conditions is provided in C.
5 Impact of measurement noise
We examined inference robustness by introducing displacement measurement noise at two levels: low ($1\times 10^{-4}\,\mathrm{mm}$) and high ($1\times 10^{-3}\,\mathrm{mm}$), applied to geometrically heterogeneous training samples, because they showed structurally better performance in the previous sections. The choice of displacement noise levels is based on similar works (Flaschel et al., 2021; Joshi et al., 2022; Thakolkaran et al., 2022; Rahmani et al., 2013). We note here that Gaussian white noise can be considered harsher than what is present in realistic data because the spatial correlation between measurements is not taken into account. Some relevant works have implemented spatially correlated noise, or adapted fictitious noise to DIC measurements directly (Jafari et al., 2025; Peyraut and Genet, 2025). By considering Gaussian white noise, we aim to validate how robust our method is.
[…]
In Figure 8, we assess parameter inference performance with low levels of displacement noise ($\sigma=10^{-4}$) for geometrically heterogeneous tissue slabs with varying levels of microstructural heterogeneity. From left to right, results are shown for (i) a fully homogeneous circ-long-het specimen, (ii) a circ-rad-het specimen with moderate transmural fiber rotation, and (iii) a rad-heli-het specimen exhibiting pronounced helical fiber variation. In the top row of the figure, we observe that the circ-long-het configuration fails to converge for the parameter value $a_{fs}$. Consequently, more stochastic behavior as well as a slight downward drift can be observed in $b_{fs}$. Although most parameter values reach good relative error values ranging from $1\times 10^{-4}$ to $5\times 10^{-3}$, the parameters associated with $I_{4fs}$ show relative errors of the same order of magnitude as the true value. This error is further highlighted by the skewness and clear truncation of the posterior distribution shown in the bottom row. The spread in values also highlights the uncertainty around these parameter values. The helically distributed fiber configurations show no such skewness in any of their parameter distributions. They do obtain higher relative errors in their parameter values. Ranging between $5\times 10^{-3}$ to $1\times 10^{-1}$ for the circ-rad configuration and $3\times 10^{-4}$ to $2\times 10^{-2}$ (excluding $a_{fs}$ and $b_{fs}$). Strikingly, the shear parameters stabilize in the circ-rad-het and rad-heli-het configurations after roughly 5,0005{,}000 and 40,00040{,}000 epochs respectively, even though their relative errors are remarkably high. Combined with the lack of skewness and truncation in their distributions, this indicates that the parameter values have moved sufficiently far away from the prior and have settled in a (sub-optimal) local minimum.
[…]
We assess the performance of the parameter fitting approach for the high-noise case in a similar way. In Figure 9, we use the same lay-out as before and observe striking differences with the low displacement noise setting. All configurations now fail to characterize the shearing parameters $a_{fs}$ and $b_{fs}$. Moreover, we now also observe that the circ-rad-het configuration fails to show stabilizing behavior for the $a_{n}$ and $b_{n}$ parameters. We also observe an increase in relative error for all parameter values in all configurations. Remarkably, the circ-long-het specimen obtains the lowest relative errors for all parameters. The circ-rad-het configuration clearly performs the worst of the three showing relative errors ranging from $2\times 10^{-1}$ to $1$. Compared to the low-noise case this indicates a sudden drop in fitting performance. All configurations show overly skew distributions in the $a_{fs}$ and $b_{fs}$ parameters, although the circ-rad configuration also shows skewness in the distributions for $a_{n}$ and $b_{n}$. The remaining distributions show standard deviations orders of magnitude lower than the mean, indicating the confidence the model has in these values. Similarly to the low-noise case we see that parameter that do stabilize, do so around 8,0008{,}000. The rad-heli-het configuration seems to show stabilizing behavior for all its parameters roughly 2,0002{,}000 epochs later.
Summarizing Figures 7, 8, and 9, we observe that the introduction of noise steadily degrades the relative error in the inferred parameter values. Furthermore, some inferred constitutive parameters start converging towards numerical zero as the amount of displacement measurement noise increases. D provides a summary of the obtained constitutive parameter values, the relative error achieved, and their standard deviation.
6 Validation
We reproduce multimodal, homogeneous loading protocols, i.e. classical incompressible biaxial tension and triaxial shear tests, following the experimental protocols of (Dokos et al., 2002; Sommer et al., 2015). Constitutive parameters inferred from our unsupervised single-shot heterogeneous biaxial protocol are then used to assess, in a controlled fashion, the predictive capability of the proposed framework relative to established multimodal calibration pipelines for myocardial tissue (Guan et al., 2018; Peirlinck et al., 2018; Kakaletsis et al., 2021; Martonová et al., 2024). Specifically, we prescribe homogeneous deformations with deformation gradient:
$\bm{F}=\left[\begin{array}[]{lll}\lambda_{\rm{f}}&\gamma_{\rm{fs}}&\gamma_{\rm{fn}}\\ \gamma_{\rm{sf}}&\lambda_{\rm{s}}&\gamma_{\rm{sn}}\\ \gamma_{\rm{nf}}&\gamma_{\rm{ns}}&\lambda_{\rm{n}}\\ \end{array}\right].$ (29)
For the biaxial tension cases, the fiber and normal stretches $\lambda_{\rm f}$ and $\lambda_{\rm n}$ are controlled, the sheet stretch follows from incompressibility, $\lambda_{\rm s}=1/(\lambda_{\rm f}\lambda_{\rm n})$, and all shear strains are set to zero, $\gamma_{\rm{fs}}=\gamma_{\rm{sf}}=\gamma_{\rm{nf}}=\gamma_{\rm{fn}}=\gamma_{\rm{sn}}=\gamma_{\rm{ns}}\equiv 0$. The hydrostatic pressure $p$ is determined by enforcing zero normal traction in the sheet direction, $\sigma_{\rm ss}=0$, which yields the following non-zero normal stresses:
[…]
Here $\psi_{\star}$ indicates the partial derivative of the strain energy density function with respect to the corresponding invariant $I_{\star}$, i.e. $\partial\psi/\partial I_{\star}$. In the triaxial shear tests, the principal stretches are fixed to unity, $\lambda_{\rm f}=\lambda_{\rm s}=\lambda_{\rm n}\equiv 1$, and only one shear component is activated per test; all others are set to zero, $\gamma_{\rm{fs}}=\gamma_{\rm{sf}}=\gamma_{\rm{nf}}=\gamma_{\rm{fn}}=\gamma_{\rm{sn}}=\gamma_{\rm{ns}}\equiv 0$, with a single component allowed to vary. Each test therefore excites a single shear mode. Given that $\gamma_{\star}\geq 0$, the resulting shear stresses are:
$\sigma_{\rm sf}=2\,\gamma_{\rm fs}\,\psi_{1}+\psi_{\rm 4fs},$ $\sigma_{\rm fs}=2\,\gamma_{\rm sf}\!\left[\psi_{1}+\psi_{\rm 4ff}\right]+\psi_{\rm 4fs},$ (31)
$\sigma_{\rm nf}=2\,\gamma_{\rm fn}\!\left[\psi_{1}+\psi_{\rm 4nn}\right],$ $\sigma_{\rm fn}=2\,\gamma_{\rm nf}\!\left[\psi_{1}+\psi_{\rm 4ff}\right],$
$\sigma_{\rm ns}=2\,\gamma_{\rm sn}\!\left[\psi_{1}+\psi_{\rm 4nn}\right],$ $\sigma_{\rm sn}=2\,\gamma_{\rm ns}\psi_{1}.$
Figure 10: Noise-free validation single-shot versus multi-modal biaxial and triaxial shear tests. Comparing Cauchy stresses based on the ground-truth parameter values, shown in tri markers, and the inferred noise-free full-field single-shot biaxial test parameter values sets, shown in shaded colors representing the $95\%$ confidence intervals resulting from the inferred posterior parameter set distributions. $R^{2}$ scores for inferred versus ground-truth stress values, for each individual biaxial tensile and triaxial shear test, are shown in the top left corners.
Figure 10 reports noise-free validation across three heterogeneous configurations: circ–long–het (left), circ–rad–het (middle), and rad–heli–het (right). Each column compares Cauchy stresses from ground-truth parameters (tri markers) with predictions from the single-shot, noise-free full-field inference constitutive parameter sets (shaded 95% credible intervals) under classical biaxial tension and triaxial shear protocols. In the circ–long–het configuration, the inferred biaxial stress–stretch curves match the ground-truth with $R^{2}=1.00$ for both biaxial tension cases. The nf-, sn-, and ns-shear responses are also indistinguishable from the ground-truth, whereas the cross-fiber shear components show lower agreement, with $R^{2}=0.89$ for $\sigma_{\mathrm{fs}}$ and $R^{2}=0.77$ for $\sigma_{\mathrm{sf}}$ respectively. Aggregated over all tests in this configuration, the overall coefficient of determination is $R^{2}=0.96$. These deviations are consistent with limited excitation of the $I_{4fs}$ shear-coupling invariant in the circ–long-het configuration as shown in Figure 5. In the circ–rad–het configuration, the inferred constitutive parameter sets reproduce the ground-truth stresses across biaxial tension and all shear tests with $R^{2}=1.00$ throughout. Posterior confidence intervals remain tight, indicating precise recovery of the active parameters for this orientation. In the rad–heli–het configuration, inferred parameter set curves overlay the ground-truth for all biaxial tensile and shear tests with $R^{2}=1.00$. Overall, the noise-free validation shows that single-shot full-field inference generalizes across all three heterogeneous configurations, with slight deviations confined to cross-fiber shear in the circ–long–het case.
Figure 11: High-noise validation single-shot versus multi-modal biaxial and triaxial shear tests. Comparing Cauchy stresses based on the ground-truth parameter values, shown in tri markers, and the inferred high-noise $\sigma=10^{-3}$ full-field single-shot biaxial test parameter values sets, shown in shaded colors representing the $95\%$ confidence intervals resulting from the inferred posterior parameter set distributions. $R^{2}$ scores for inferred versus ground-truth stress values, for each individual biaxial tensile and triaxial shear test, are shown in the top left corners.
Figure 11 reports high-noise ($\sigma=10^{-3}$) validation across three heterogeneous configurations: circ–long–het (left), circ–rad–het (middle), and rad–heli–het (right). Each column compares Cauchy stresses from ground-truth parameters with predictions from the single-shot full-field inference under displacement noise $\sigma=10^{-3}$ (shaded 95% confidence intervals). In the circ–long–het configuration, inferred biaxial tension behavior remains perfectly aligned with the ground-truth stress behavior. Similarly, the nf-, sn-, and ns-shear responses retain correct magnitude and slope. Noise primarily affects the inferred cross-fiber shear stress $\sigma_{\mathrm{fs}}$ and $\sigma_{\mathrm{sf}}$ response, where credible intervals widen and small deviations between ground-truth and inferred behaviors appear. Aggregated over all tests in this configuration, the overall coefficient of determination is $R^{2}=0.96$. In the circ–rad–het configuration, the stresses computed from the inferred parameter sets under noisy full-field biaxial tensile tests, no longer fit the ground-truth stress curves perfectly. Mean trends are preserved across biaxial tensile and shear protocols, but panel-wise $R^{2}$ values are lower than those in the high-noise circ-long-het and rad-heli-het configurations. Nevertheless, the aggregated coefficient of determination for this configuration still remains $R^{2}=0.77$. In the rad–heli–het configuration, we see the best overall inferred stress responses across all biaxial tensile and shear stress responses, with $R^{2}$-values greater than 0.88 across the board. Here, the total aggregated coefficient of determination equals $R^{2}=0.93$. Overall, despite elevated displacement and thus deformation noise, the single-shot SVI inference approach maintains very good quantitative agreement with ground-truth stresses across all three heterogeneous configurations.
7 Discussion
The ability to infer complex orthotropic hyperelastic models from a single full-field biaxial tensile testing experiment represents an interesting avenue in biomechanical tissue characterization. Traditional workflows typically rely on multiple specialized loading configurations and assume simplified model forms to ensure parameter identifiability (Dokos et al., 2002; Sommer et al., 2015; Avazmohammadi et al., 2018). In contrast, our framework leverages specimen-intrinsic microstructural (and geometrical) heterogeneity, combined with stochastic variational inference, to robustly recover full sets of nonlinear, cross-correlated material parameters under realistic noise conditions. This approach not only reduces experimental tissue manipulation but also expands opportunities to characterize spatially heterogeneous biological tissues, where sample availability or repeat testing is limited. Moreover, it lays the groundwork for practical integration of physics-informed Bayesian inference into future experimental biomechanics and computational modeling pipelines. Additionally, we offer the possibility of taking inter-sample and intra-sample variability into account for materials where this was previously not an option.
Power of tissue testing heterogeneity in constitutive parameter identifiability. Our results show that introducing both geometric and microstructural heterogeneity in the training specimen substantially improves the identifiability of constitutive model parameters. The addition of a simple circular occlusion generates spatial strain gradients and local stress concentrations that increase the sensitivity of internal forces to variations in model parameters (Attar et al., 2022; Conde et al., 2023; Pottier et al., 2011; Pierron and Grédiac, 2021). Similarly, strategically aligning fiber orientations relative to the applied boundary conditions enhances the excitation of multiple deformation invariants (Zhang et al., 2022; Kakaletsis et al., 2021; Linka et al., 2021). These experimental design choices enabled stable recovery of all parameters associated with dominant energy modes ($I_{1}$, $I_{4ff}$), and even parameters associated with more subtle contributions ($I_{4fs}$) when measurement noise was low. Importantly, both forms of heterogeneity complemented each other: geometric features enriched the macroscopic stress states, while fiber orientation tuning ensured multiaxial activation of the microstructure. These findings suggest that future experimental protocols could deliberately incorporate both microstructural and geometrical tailoring to maximize parameter observability.
Homogeneous loading conditions heavily impact orthotropic inference power. Samples with homogeneously distributed fiber orientations and without geometric features exhibited markedly worse parameter convergence (Pottier et al., 2011). In these homogeneous specimens, the dominant loading paths aligned primarily with the $I_{1}$ and $I_{4ff}$ invariants, while parameters associated with fiber-sheet coupling ($I_{4fs}$) received little excitation, as previously postulated by (Holzapfel and Ogden, 2009; Guan et al., 2018). As a result, posterior distributions for associated parameters such as $a_{fs}$ and $b_{fs}$ remained close to their prior, showing truncation effects and large uncertainty. Interestingly, under high measurement noise, homogeneous configurations experienced smaller decrease in performance in validation compared to heterogeneous specimens. This reflects a robustness trade-off: simpler configurations concentrate strain energy into a few modes, reducing sensitivity to noise at the cost of losing full identifiability. These findings highlight that while homogeneity can be beneficial under extreme noise, achieving accurate full-model inference requires careful experimental enrichment.
Measurement noise, as expected, degrades inference performance. At a displacement noise level of $1\times 10^{-4}$ mm, the method still achieved low relative errors and tightly concentrated posterior distributions for most constitutive parameters. The shear related parameter $a_{fs}$ and $b_{fs}$, however, show larger deviation from the ground-truth. At $1\times 10^{-3}$ mm noise, significant broadening and bias in the posterior appeared, particularly for parameters linked to minor strain energy contributions. Namely, the fourth coupling invariant $I_{4fs}$ showed strong vulnerability to noise, with associated parameters $a_{fs}$ and $b_{fs}$ tending toward zero or staying close to the prior resulting in a truncated posterior distribution, which is in agreement with the loss in accuracy for anisotropic features in (Rahmani et al., 2013). This effectively removed their contribution to the strain energy function in the inferred models. Despite these deviations, the validation strain energy and stresses remained close to the ground-truth. This suggests that some parameter errors can be tolerated without substantial loss of predictive accuracy for macroscopic mechanical behavior. Improving noise robustness through enhanced experimental design, probabilistic displacement modeling, or variational reweighting strategies offer important future directions. (Marek et al., 2019; Jafari et al., 2025; Peyraut and Genet, 2025)
Single-shot inferred material models validate well to multimodal testing protocols - even under high noise. We show that we can reproduce the behavior from classical homogeneous multimodal biaxial tensile and triaxial shear testing. Even at high noise levels, we see better overall agreement with the ground-truth stresses from a single heterogeneous full-field biaxial tensile test than we generally see for separate biaxial tensile-only, or triaxial shear-only training. Interestingly, the circ-rad-het configuration seems to be most sensitive to displacement noise when considering pure biaxial tensile and triaxial shear test loading, despite this configuration’s richer heterogeneous deformation profile shown in Figure 5. Both the circ-long-het and rad-heli-het samples show good extrapolative behavior, even under high noise levels when the shearing component in the strain energy vanishes, forming interesting slicing configurations for future experimental studies of myocardial tissue behavior.
8 Conclusion
We presented a physics-informed stochastic variational framework that recovers the full constitutive parameter set of a nonlinear orthotropic hyperelastic model from a single biaxial stretch experiment. By exploiting intrinsic microstructural heterogeneity, or introducing simple geometric heterogeneity features when needed, the method identifies not only the dominant tensile parameters but also the shear-coupling terms that are typically weakly excited in a biaxial tensile test. Our results demonstrate that an effective balance between underlying microstructure and biaxial loading improves identifiability, and this balance can be set during slicing and tissue harvesting. Circumferential–radial slabs and rotated cuts that place both fiber and sheet directions in the loading plane enable full recovery under noise-free conditions. Under realistic displacement measurement noise, the approach remains robust and yields stresses and strain energies in close agreement with ground-truth simulations. Adding geometric heterogeneity, for example a central occlusion, systematically improves convergence and tightens posterior distributions, particularly for shear parameters. These results indicate that slicing protocols and experimental design can be used deliberately to enhance observability in full-field inverse characterization. More broadly, the framework enables localized characterization of orthotropic nonlinear biological tissues using a single-mode setup. This reduces experimental tissue manipulation burden, improves reproducibility, and supports studies where available tissue is limited.
Appendix A Tissue slab geometry and loading conditions
We summarize the tissue slab geometry and loading conditions used for synthetic data generation in Figure 1.
Figure 1: Synthetic data generation: training sample dimensions and loading conditions Tissue slab geometry and applied boundary conditions for the training samples. We subject all training samples undergo to a biaxial loading protocol $\lambda_{1}:\lambda_{2}-(1:2),(1:1),(2:1)$ up to $15\%$ of the initial length of the sample. All samples are of size 10 $\times$ 10 $\times$ 1 mm3. The geometrically heterogeneous tissue samples have a circular occlusion in the center with a radius of $R=1$ mm.
Appendix B Analytical derivations
We summarize the analytical derivation of the Holzapfel-Ogden model stress used in our code below:
$\bm{P}\left(\bm{F},\bm{\theta}\right)=$ $2\bm{F}\frac{\partial\psi(\bm{F})}{\partial\bm{C}}$ (32)
$=$ $2\bm{F}\left(\frac{\partial\bar{\psi}(\bar{I}_{1})}{\partial\bar{I}_{1}}\frac{\partial\bar{I}_{1}}{\partial\bm{C}}+\frac{\partial\bar{\psi}(\bar{I}_{4f})}{\partial\bar{I}_{4f}}\frac{\partial\bar{I}_{4f}}{\partial\bm{C}}\right$.
$+\left.\frac{\partial\bar{\psi}(\bar{I}_{4s})}{\partial\bar{I}_{4s}}\frac{\partial\bar{I}_{4s}}{\partial\bm{C}}+\frac{\partial\bar{\psi}(\bar{I}_{4fs})}{\partial\bar{I}_{4fs}}\frac{\partial\bar{I}_{4fs}}{\partial\bm{C}}+K\frac{\bm{J}^{2}-1}{2}\bm{C}^{-1}\right)$
$\frac{\partial\bar{\psi}(\bar{I}_{1})}{\partial\bar{I}_{1}}$ $=\frac{a}{2}exp(b(\bar{I}_{1}-3))$ (33)
$\frac{\partial\bar{\psi}(\bar{I}_{4i})}{\partial\bar{I}_{4i}}$ $=a_{i}\langle\bar{I}_{4i}-1\rangle\exp(b_{i}\langle\bar{I}_{4i}-1\rangle^{2})$
$\frac{\partial\bar{\psi}(\bar{I}_{4fs})}{\partial\bar{I}_{4fs}}$ $=a_{fs}\bar{I}_{4fs}exp(b_{fs}\bar{I}_{4fs}^{2})$
$\frac{\partial\bar{I}_{1}}{\partial\bm{C}}$ $=J^{-2/3}\frac{\partial I_{1}}{\partial\bm{C}}+I_{1}\frac{\partial J^{-2/3}}{\partial\bm{C}}=\bm{J}^{-2/3}\bm{I}-\frac{1}{3}\bar{I}_{1}\bm{C}^{-1}$ (34)
$\frac{\partial\bar{I}_{4i}}{\partial\bm{C}}$ $=J^{-2/3}\frac{\partial I_{4i}}{\partial\bm{C}}+I_{4i}\frac{\partial J^{-2/3}}{\partial\bm{C}}=\bm{J}^{-2/3}(\bm{i}\otimes\bm{i})-\frac{1}{3}\bar{I}_{4i}\bm{C}^{-1}$
$\frac{\partial\bar{I}_{4fs}}{\partial\bm{C}}$ $=J^{-2/3}\frac{\partial I_{4fs}}{\partial\bm{C}}+I_{4fs}\frac{\partial J^{-2/3}}{\partial\bm{C}}=\frac{1}{2}\bm{J}^{-2/3}(\bm{f}\otimes\bm{s}+\bm{s}\otimes\bm{f})-\frac{1}{3}\bar{I}_{4fs}\bm{C}^{-1}$
And the Kullback-Leibler derivation:
$KL\left[\pi(\bm{\theta},V_{s},\sigma^{2}|\bm{b})||q(\bm{\theta},V_{s},\sigma^{2})\right]$ $=KL\left[\pi(\bm{\Phi}|\bm{b})||q(\bm{\Phi})\right]=\int q(\bm{\Phi})log\left[\frac{q(\bm{\Phi})}{\pi(\bm{\Phi}|\bm{b})}\right]d\bm{\Phi}$ (35)
$=\int q(\bm{\Phi})log\left[\frac{q(\bm{\Phi})\pi(\bm{b})}{\pi(\bm{\Phi},\bm{b})}\right]d\bm{\Phi}$
$=\int q(\bm{\Phi})log\left[\frac{q(\bm{\Phi})}{\pi(\bm{\Phi},\bm{b})}\right]d\bm{\Phi}+\int q(\bm{\Phi})log[\pi(\bm{b})]d\bm{\Phi}$
$=\int q(\bm{\Phi})log\left[\frac{q(\bm{\Phi})}{\pi(\bm{\Phi},\bm{b})}\right]d\bm{\Phi}+log[\pi(\bm{b})]$
$=-\int q(\bm{\Phi})log\left[\frac{\pi(\bm{b}|\bm{\Phi})\pi(\bm{\Phi})}{q(\bm{\Phi})}\right]d\bm{\Phi}+log[\pi(\bm{b})]$
$=-\int q(\bm{\Phi})\left(log\left[\frac{\pi(\bm{\Phi})}{q(\bm{\Phi})}\right]+log[\pi(\bm{b}|\bm{\Phi})]\right)d\bm{\Phi}+log[\pi(\bm{b})]$
$=-\int q(\bm{\Phi})log\left[\frac{\pi(\bm{\Phi})}{q(\bm{\Phi})}\right]d\bm{\Phi}-\int q(\bm{\Phi})log[\pi(\bm{b}|\bm{\Phi})]d\bm{\Phi}+log[\pi(\bm{b})]$
Appendix C Noise-free orthotropic constitutive parameter inference results
This appendix contains the numerical results for the parameter fitting procedure when no displacement noise is applied to the experiments. In Table 2, we show the optimized hyperparameters at the final epoch. The parameters themselves and their accompanying standard deviation are reported in Table 3. Finally, we also report the relative error between the estimated material model parameters and the ground-truth in Table 4.
Table 2: Hyperparameter estimates. The hyperparameters at the final epoch of the optimization procedure for the samples with and without circular occlusion. The hyperparameters are both for inverse-gamma distributions where $\alpha_{\star}$, $\beta_{\star}$ are the scale and rate parameters.
Hyperparameters circ-long-hom circ-rad-hom rad-heli-hom circ-long-het circ-rad-het rad-heli-het
$\alpha_{vs}$ $1.914\times 10^{1}$ $1.928\times 10^{1}$ $1.928\times 10^{1}$ $1.916\times 10^{1}$ $1.930\times 10^{1}$ $1.930\times 10^{1}$
$\alpha_{\sigma}$ $5.215\times 10^{6}$ $5.050\times 10^{6}$ $5.101\times 10^{6}$ $5.400\times 10^{6}$ $5.127\times 10^{6}$ $5.292\times 10^{6}$
$\beta_{vs}$ $7.401\times 10^{7}$ $7.201\times 10^{7}$ $7.169\times 10^{7}$ $924\times 10^{7} 7.7$$20\times 10^{7} 7.7$$714\times 10^{7}$
$\beta_{\sigma}$ $1.901\times 10^{1}$ $1.836\times 10^{1}$ $1.854\times 10^{1}$ $1.834\times 10^{1}$ $1.733\times 10^{1}$ $1.789\times 10^{1}$
Table 3: Parameter estimates. The parameters at the final epoch of the optimization procedure for the samples with and without circular occlusion. Parameters in $\bm{\theta}$ are written out in full to correspond to the HO-model parameters. The corresponding standard deviation has also been reported.
Parameters circ-long-hom circ-rad-hom rad-heli-hom circ-long-het circ-rad-het rad-heli-het
mean:
$a$ $8.090\times 10^{-4}$ $8.090\times 10^{-4}$ $8.090\times 10^{-4}$ $8.090\times 10^{-4}$ $8.090\times 10^{-4}$ $8.090\times 10^{-4}$
$b$ $74\times 10^{0} 7.474$$\times 10^{0} 7.474$$\times 10^{0} 7.474$$\times 10^{0} 7.474$$\times 10^{0} 7.474$$74\times 10^{0}$
$a_{f}$ $1.911\times 10^{-3}$ $1.911\times 10^{-3}$ $1.911\times 10^{-3}$ $1.911\times 10^{-3}$ $1.911\times 10^{-3}$ $1.911\times 10^{-3}$
$b_{f}$ $206\times 10^{1} 2.2$$06\times 10^{1} 2.2$$06\times 10^{1} 2.2$$06\times 10^{1} 2.2$$06\times 10^{1} 2.2$$206\times 10^{1}$
$a_{n}$ $270\times 10^{-4} 2.2$$70\times 10^{-4} 2.2$$70\times 10^{-4} 2.2$$70\times 10^{-4} 2.2$$70\times 10^{-4} 2.2$$270\times 10^{-4}$
$b_{n}$ $3.480\times 10^{1}$ $3.479\times 10^{1}$ $3.480\times 10^{1}$ $3.480\times 10^{1}$ $3.479\times 10^{1}$ $3.480\times 10^{1}$
$a_{fs}$ $7.806\times 10^{-2}$ $5.426\times 10^{-4}$ $5.480\times 10^{-4}$ $5.414\times 10^{-4}$ $5.484\times 10^{-4}$ $5.476\times 10^{-4}$
$b_{fs}$ $4.548\times 10^{-1}$ $4.715\times 10^{0}$ $3.561\times 10^{0}$ $6.862\times 10^{-1}$ $5.700\times 10^{0}$ $5.652\times 10^{0}$
$K$ $0\times 10^{-1} 1.000×10$$0\times 10^{-1}$ $0\times 10^{-1} 1.000×10$$0\times 10^{-1}$ $0\times 10^{-1} 1.000×10$$0\times 10^{-1}$
variance:
$\sigma_{a}$ $1.792\times 10^{-8}$ $1.513\times 10^{-8}$ $1.518\times 10^{-8}$ $1.756\times 10^{-8}$ $1.684\times 10^{-8}$ $1.600\times 10^{-8}$
$\sigma_{b}$ $1.223\times 10^{-4}$ $9.733\times 10^{-5}$ $9.650\times 10^{-5}$ $1.268\times 10^{-4}$ $064\times 10^{-4} 1.1$$125\times 10^{-4}$
$\sigma_{af}$ $4.608\times 10^{-9}$ $2.851\times 10^{-8}$ $2.972\times 10^{-8}$ $5.507\times 10^{-9}$ $4.784\times 10^{-8}$ $2.798\times 10^{-8}$
$\sigma_{bf}$ $2.990\times 10^{-5}$ $2.072\times 10^{-4}$ $2.109\times 10^{-4}$ $3.661\times 10^{-5}$ $6.346\times 10^{-4}$ $2.022\times 10^{-4}$
$\sigma_{an}$ $514\times 10^{-9} 1.1$$100\times 10^{-7}$ $8.551\times 10^{-9}$ $1.380\times 10^{-9}$ $1.084\times 10^{-7}$ $8.144\times 10^{-9}$
$\sigma_{bn}$ $6.611\times 10^{-5}$ $3.699\times 10^{-2}$ $4.123\times 10^{-4}$ $7.526\times 10^{-5}$ $3.117\times 10^{-2}$ $3.880\times 10^{-4}$
$\sigma_{afs}$ $1.906\times 10^{0}$ $4.088\times 10^{-5}$ $4.207\times 10^{-5}$ $6.564\times 10^{-5}$ $4.102\times 10^{-6}$ $1.034\times 10^{-5}$
$\sigma_{bfs}$ $1.370\times 10^{1}$ $4.410\times 10^{0}$ $3.996\times 10^{0}$ $1.283\times 10^{1}$ $4.534\times 10^{-2}$ $2.728\times 10^{-1}$
$\sigma_{k}$ $2.363\times 10^{-7}$ $6.853\times 10^{-7}$ $5.688\times 10^{-7}$ $2.296\times 10^{-7}$ $8.327\times 10^{-7}$ $5.612\times 10^{-7}$
Table 4: Relative error of the parameter estimates. The relative error at the final epoch between the parameter values and the ground-truth for samples with and without circular occlusion and various microstructures.
Relative errors circ-long-hom circ-rad-hom rad-heli-hom circ-long-het circ-rad-het rad-heli-het
$e_{rel,a}$ $4.765\times 10^{-5}$ $1.864\times 10^{-5}$ $2.925\times 10^{-5}$ $1.832\times 10^{-5}$ $1.644\times 10^{-5}$ $2.593\times 10^{-5}$
$e_{rel,b}$ $2.298\times 10^{-6}$ $3.659\times 10^{-5}$ $3.854\times 10^{-5}$ $1.223\times 10^{-5}$ $2.795\times 10^{-5}$ $3.428\times 10^{-5}$
$e_{rel,af}$ $5.495\times 10^{-5}$ $8.243\times 10^{-5}$ $7.541\times 10^{-5}$ $4.184\times 10^{-5}$ $1.356\times 10^{-5}$ $8.322\times 10^{-5}$
$e_{rel,bf}$ $1.471\times 10^{-6}$ $4.349\times 10^{-5}$ $4.047\times 10^{-5}$ $9.769\times 10^{-6}$ $8.735\times 10^{-6}$ $3.724\times 10^{-5}$
$e_{rel,an}$ $9.655\times 10^{-5}$ $1.844\times 10^{-4}$ $3.895\times 10^{-5}$ $7.867\times 10^{-5}$ $2.017\times 10^{-4}$ $4.702\times 10^{-5}$
$e_{rel,bn}$ $1.290\times 10^{-5}$ $2.250\times 10^{-4}$ $5.078\times 10^{-5}$ $1.396\times 10^{-5}$ $4.682\times 10^{-4}$ $4.864\times 10^{-5}$
$e_{rel,afs}$ $1.417\times 10^{2}$ $8.050\times 10^{-3}$ $1.901\times 10^{-3}$ $1.023\times 10^{-2}$ $2.520\times 10^{-3}$ $1.137\times 10^{-3}$
$e_{rel,bfs}$ $9.201\times 10^{-1}$ $1.716\times 10^{-1}$ $3.743\times 10^{-1}$ $8.794\times 10^{-1}$ $1.535\times 10^{-3}$ $6.900\times 10^{-3}$
$e_{rel,K}$ $2.423\times 10^{-6}$ $2.932\times 10^{-5}$ $4.502\times 10^{-5}$ $4.837\times 10^{-6}$ $1.661\times 10^{-5}$ $4.184\times 10^{-5}$
Appendix D Noise-affected orthotropic constitutive parameter inference results
This appendix contains the numerical results for the parameter fitting procedure when no displacement noise is applied to the experiments. In Table 5, we show the optimized hyperparameters at the final epoch. The parameters themselves and their accompanying standard deviation are reported in Table 6. Finally, we also report the relative error between the estimated material model parameters and the ground-truth in Table 7.
Table 5: Hyperparameter estimates for noisy experiments. The hyperparameters at the final epoch of the optimization procedure for the samples with circular occlusion and additive displacement noise at high ($\sigma=1e-3$) and low ($\sigma=1e-4$) levels. The hyperparameters are both for inverse-gamma distributions where $\alpha_{\star}$, $\beta_{\star}$ are the scale and rate parameters.
Hyperparameters circ-long-hom circ-rad-hom rad-heli-hom circ-long-het circ-rad-het rad-heli-het
$\sigma=1e^{-4}$ $\sigma=1e^{-4}$ $\sigma=1e^{-4}$ $\sigma=1e^{-3}$ $\sigma=1e^{-3}$ $\sigma=1e^{-3}$
$\alpha_{vs}$ 1.922e+01 1.930e+01 1.928e+01 1.922e+01 1.918e+01 1.923e+01
$\alpha_{\sigma}$ 2.615e+06 2.678e+06 2.683e+06 8.849e+05 1.047e+06 9.640e+05
$\beta_{vs}$ 1.977e+07 2.076e+07 2.479e+07 2.168e+06 1.052e+06 2.585e+06
$\beta_{\sigma}$ 3.598e+01 3.166e+01 3.250e+01 1.079e+02 8.942e+01 9.801e+01
Table 6: Noisy parameter estimates. The parameters at the final epoch of the optimization procedure for the samples with circular occlusion and additive displacement noise at high ($\sigma=1e-3$) and low ($\sigma=1e-4$) levels. Parameters in $\bm{\theta}$ are written out in full to correspond to the HO-model parameters. The corresponding standard deviation has also been reported.
Parameters circ-long-hom circ-rad-hom rad-heli-hom circ-long-het circ-rad-het rad-heli-het
$\sigma=1e^{-4}$ $\sigma=1e^{-4}$ $\sigma=1e^{-4}$ $\sigma=1e^{-3}$ $\sigma=1e^{-3}$ $\sigma=1e^{-3}$
mean:
$a$ 8.086e-04 8.051e-04 8.054e-04 8.101e-04 9.909e-04 9.237e-04
$b$ 7.475e+$0$ 7.535e+$0$ 7.489e+$0$ 7.823e+$0$ 8.773e+$0$ 7.919e+$0$
$a_{f}$ 1.919e-03 1.988e-03 1.946e-03 2.004e-03 3.670e-03 2.041e-03
$b_{f}$ 2.202e+01 2.134e+01 2.183e+01 2.129e+01 9.790e+$0$ 2.086e+01
$a_{n}$ 2.259e-04 2.366e-04 2.288e-04 2.325e-04 3.209e-10 2.496e-04
$b_{n}$ 3.484e+01 3.133e+01 3.479e+01 3.377e+01 1.865e-01 3.404e+01
$a_{fs}$ 1.644e-08 2.590e-04 4.239e-05 1.224e-08 8.290e-09 1.039e-08
$b_{fs}$ 5.301e-01 9.085e+$0$ 2.487e+01 5.017e-01 1.953e-01 1.499e-02
$K$ 1.000e-01 9.952e-02 1.001e-01 9.922e-02 7.003e-02 9.155e-02
variance:
$\sigma_{a}$ 6.561e-08 5.847e-08 5.717e-08 5.420e-07 3.137e-07 4.306e-07
$\sigma_{b}$ 4.958e-04 3.947e-04 4.232e-04 4.235e-03 1.809e-03 2.792e-03
$\sigma_{af}$ 1.933e-08 1.765e-07 8.679e-08 1.813e-07 1.922e-06 7.681e-07
$\sigma_{bf}$ 1.254e-04 2.288e-03 6.918e-04 1.026e-03 1.473e-02 5.697e-03
$\sigma_{an}$ 5.369e-09 3.939e-07 2.461e-08 5.067e-08 5.002e-08 2.214e-07
$\sigma_{bn}$ 2.759e-04 1.138e-01 1.364e-03 2.398e-03 4.350e+$0$ 1.103e-02
$\sigma_{afs}$ 3.926e-06 8.179e-06 6.189e-06 3.030e-06 1.820e-06 2.152e-06
$\sigma_{bfs}$ 1.365e+01 1.662e-01 1.076e+$0$ 1.323e+01 4.074e+$0$ 1.253e+01
$\sigma_{K}$ 9.586e-07 2.898e-06 2.220e-06 8.185e-06 1.974e-05 1.785e-05
Table 7: Relative errors of the noisy parameter estimates. The relative error at the final epoch between the parameter values and the ground-truth for samples with circular occlusion, additive displacement noise at high ($\sigma=1e-3$) and low ($\sigma=1e-4$) levels and various microstructures.
Parameters circ-long-hom circ-rad-hom rad-heli-hom circ-long-het circ-rad-het rad-heli-het
$\sigma=1e^{-4}$ $\sigma=1e^{-4}$ $\sigma=1e^{-4}$ $\sigma=1e^{-3}$ $\sigma=1e^{-3}$ $\sigma=1e^{-3}$
$e_{rel,a}$ 4.632e-04 4.835e-03 4.466e-03 -1.355e-03 -2.248e-01 -1.418e-01
$e_{rel,b}$ -1.793e-04 -8.132e-03 -2.026e-03 -4.671e-02 -1.738e-01 -5.960e-02
$e_{rel,af}$ -4.334e-03 -4.031e-02 -1.806e-02 -4.855e-02 -9.203e-01 -6.801e-02
$e_{rel,bf}$ 1.847e-03 3.258e-02 1.073e-02 3.515e-02 5.563e-01 5.470e-02
$e_{rel,an}$ 5.049e-03 -4.218e-02 -7.769e-03 -2.432e-02 1.000e+$0$ -9.956e-02
$e_{rel,bn}$ -1.119e-03 9.979e-02 2.589e-04 2.969e-02 9.946e-01 2.204e-02
$e_{rel,afs}$ 1.000e+$0$ 5.264e-01 9.225e-01 1.000e+$0$ 1.000e+$0$ 1.000e+$0$
$e_{rel,bfs}$ 9.069e-01 -5.964e-01 -3.369e+$0$ 9.118e-01 9.657e-01 9.974e-01
$e_{rel,K}$ -1.137e-04 4.845e-03 -1.068e-03 7.770e-03 2.997e-01 8.453e-02

### v2

@@
 Cardiac muscle tissue exhibits highly non-linear hyperelastic and orthotropic material behavior during passive deformation.
-Fully capturing this behavior in traditional homogenized tissue testing requires the excitation of multiple deformation modes, i.e.
-combined triaxial shear tests and biaxial stretch tests.
-Inherently, such multimodal experimental protocols necessitate multiple tissue samples and extensive sample manipulations.
-Intrinsic inter-sample variability and manipulation-induced tissue damage might have an adverse effect on the inversely identified tissue behavior.
-In this work, we aim to overcome this gap by focusing our attention to the use of heterogeneous deformation profiles in a parameter estimation problem.
-More specifically, we adapt EUCLID, an unsupervised method for the automated discovery of constitutive models, towards the purpose of parameter identification for highly nonlinear, orthotropic constitutive models using a Bayesian inference approach and three-dimensional continuum elements.
-We showcase its strength to quantitatively infer – with varying noise levels – the material model parameters of synthetic myocardial tissue slabs from a single heterogeneous biaxial stretch test.
-This method shows good agreement with the ground-truth simulations and with corresponding credibility intervals.
-Our work highlights the potential for characterizing highly nonlinear and orthotropic material models from a single biaxial stretch test with uncertainty quantification.
+Traditional constitutive identification protocols therefore combine multiple loading modes and typically require multiple specimens and substantial handling.
+In soft living tissues, such protocols are challenged by inter- and intra-sample variability and by manipulation-induced alterations of mechanical response, which can bias inverse calibration.
+In this work we exploit spatially heterogeneous full-field kinematics as an information-rich alternative to multimodal testing.
+We adapt EUCLID, an unsupervised method for the automated discovery of constitutive models, towards Bayesian parameter inference for highly nonlinear, orthotropic constitutive models.
+Using synthetic myocardial tissue slabs, we demonstrate that a single heterogeneous biaxial experiment, combined with sparse reaction-force measurements, enables robust recovery of Holzapfel–Ogden parameters with quantified uncertainty, across multiple noise levels.
+The inferred responses agree closely with ground-truth simulations and yield credible intervals that reflect the impact of measurement noise on orthotropic material model inference.
+Our work supports single-shot, uncertainty-aware characterization of nonlinear orthotropic material models from a single biaxial test, reducing sample demand and experimental manipulation.
 keywords:
-material model inference , full-field data , uncertainty quantification , anisotropic hyperelasticity , experimental tissue testing , Bayesian inference
+material model inference , full-field kinematics , uncertainty quantification , anisotropic hyperelasticity , experimental tissue testing , Bayesian inference
 journal: arXiv
@@
 Soft biological tissues exhibit complex mechanical responses due to their fibrous, anisotropic microstructure.
-Accurately characterizing this behavior is essential for constructing predictive constitutive models that support simulation-driven design, diagnostics, and treatment planning (Peirlinck et al., 2021).
-However, unlike engineered materials, soft tissues are subject to considerable biological variability, both within a patient, a patient cohort, and experimental study (Reeps et al., 2012; Roccabianca et al., 2014; Budday et al., 2017; Matouš et al., 2017).
-This intra- and inter-sample variability is even more pronounced in the myocardium, where strong spatial gradients in myofiber orientation and laminar architecture govern regional mechanical function (Rohmer et al., 2007; Alexander J.
-Wilson, 2022).
-In this study we investigate the impact that the underlying microstructural organization variability has on the mechanical behavior of a single sample.
-Traditional approaches to constitutive modeling of such tissues typically rely on multiple mechanical tests across different deformation modes including uniaxial tension or compression, biaxial tension, or shear testing (Dokos et al., 2002; Sommer et al., 2015; Kakaletsis et al., 2021).
-These protocols often require combining data from different specimens and, as such, ignore and smooth out this intrinsic intra- and inter-sample variability and lead to a smoothened out, average response (Staber and Guilleminot, 2017).
-Moreover, these protocols involve strong assumptions, such as spatially homogeneous stress and strain fields within the region of interest (Holzapfel and Ogden, 2009; Guan et al., 2018; Martonová et al., 2024).
-Such assumptions can be problematic for living tissues, where spatial heterogeneity is the norm, and the mechanical response may depend on subtle differences in fiber orientation, geometry, or mounting.
-Moreover, invasive sample manipulations, such as cutting, clamping, and gluing, can alter the intrinsic mechanical behavior that these tests seek to measure (Fehervary et al., 2018; Alloisio et al., 2024).
+Accurately characterizing this behavior is essential for constructing predictive constitutive models that support simulation-driven design, diagnostics, and treatment planning [1, 2].
+However, unlike engineered materials, soft tissues are subject to considerable biological variability, both within a patient, across a patient cohort, and across experimental studies [3, 4, 5, 6, 7].
+This intra- and inter-sample variability is even more pronounced in the myocardium, where strong spatial gradients in myofiber orientation and laminar architecture govern regional mechanical function [8, 9].
+Here, we investigate the impact that the underlying microstructural organization variability has on the mechanical behavior of a single sample.
+Traditional approaches to constitutive modeling of such tissues typically rely on multiple mechanical tests across different deformation modes including uniaxial tension or compression, biaxial tension, or shear testing [10, 11, 12].
+These protocols often require combining data from different specimens and, as such, ignore and smooth out intrinsic intra- and inter-sample variability into an average response [13].
+Moreover, these protocols involve strong assumptions, such as spatially homogeneous stress and strain fields within the region of interest [14, 15, 16].
+Such assumptions are problematic for living tissues, where spatial heterogeneity is the norm, and the mechanical response may depend on subtle differences in fiber orientation, geometry, or mounting.
+In addition, invasive sample manipulations, such as cutting, clamping, and gluing, can alter the intrinsic mechanical behavior that these tests seek to measure [17, 18, 19].
 These challenges have prompted the development of single-sample, single-experiment methodologies that aim to infer full constitutive behavior from a single loading protocol.
-This shift has been enabled by advances in full-field deformation measurements, such as digital image correlation and imaging-based 3D kinematics (Wang et al., 2021; Jailin et al., 2024; Meng et al., 2025).
+This shift has been enabled by advances in full-field deformation measurements, such as digital image correlation and imaging-based 3D kinematics [20, 21, 22].
 When applied to samples undergoing spatially heterogeneous deformation, such data enable the inference of multiple material parameters simultaneously, reducing the need for multiple loading modes.
-Indeed, unsupervised identification techniques, such as the virtual fields method (Grédiac et al., 2006; Pierron and Grédiac, 2012), finite element updating (Kavanagh and Clough, 1971; Peirlinck et al., 2018; Elouneg et al., 2021), variational systems identification (Wang et al., 2021), the equilibrium gap method (Claire et al., 2004; Peyraut and Genet, 2025), and EUCLID (Flaschel et al., 2021; Thakolkaran et al., 2022; Joshi et al., 2022; Thakolkaran et al., 2025), demonstrate that heterogeneous strain states, once seen as a nuisance, can be leveraged as rich sources of information, obviating the need for homogenizing assumptions.
-In this study, we explore the validity of a simple biaxial tensile testing protocol on a single sample towards the identification of the underlying orthotropic behavior of a myocardial tissue slab.
+Indeed, unsupervised identification techniques, such as the virtual fields method (VFM) [23, 24], finite element model updating (FEMU) [25, 26, 27], variational systems identification [20], the equilibrium gap method (EGM) [28, 29], and EUCLID [30, 31, 32, 33], demonstrate that heterogeneous strain states, once seen as a nuisance, can be leveraged as rich sources of information, obviating the need for homogenizing assumptions.
+In this study, we assess the extent to which a simple biaxial tensile testing protocol on a single myocardial tissue slab can support identification of the underlying orthotropic behavior.
 Yet, a key limitation persists in most current workflows: the lack of uncertainty quantification.
 Material parameters are frequently reported as point estimates, with little accounting for the inherent measurement noise or biological variability.
-In the context of myocardial modeling, where even adjacent samples can differ markedly in stiffness and fiber orientation (Dokos et al., 2002; Nielles-Vallespin et al., 2017), such deterministic inference may be misleading (Hauseux et al., 2018).
-Recent work has begun to address this gap through Bayesian approaches to inverse modeling, which offer not only best-fit estimates but also posterior distributions reflecting uncertainty in data and model form (Ozturk et al., 2021; Joshi et al., 2022; Aggarwal et al., 2023; Peshave et al., 2024; Anton et al., 2024; Taç et al., 2024; Linka et al., 2025).
-Within the unsupervised setting, Bayesian-EUCLID (Joshi et al., 2022) forms an interesting contribution towards the inference of myocardial tissue behavior from full-field displacement fields and force measurements.
-Joshi et al.
-demonstrated the feasibility of using Gibbs sampling for full-field inference in two-dimensional problems, where the material response could be inferred from a linear system of equations for the unknown material parameters in a rich constitutive feature library.
-In contrast, the setting of this study involves two distinct and compounding challenges: first, the underlying constitutive models are highly nonlinear (Costa et al., 2001; Holzapfel and Ogden, 2009; Martonová et al., 2024), with strongly cross-correlated parameters (Guan et al., 2018; Laita et al., 2025) that give rise to complex posterior landscapes; and second, the transition to three-dimensional full-field data.
+In the context of myocardial modeling, where even adjacent samples can differ markedly in stiffness and fiber orientation [10, 34], such deterministic inference may be misleading [35].
+Recent work has begun to address this gap through Bayesian approaches to inverse modeling, which offer not only best-fit estimates but also posterior distributions reflecting uncertainty in data and model form [36, 32, 37, 38, 39, 40, 41].
+Within the unsupervised setting, Bayesian-EUCLID forms an interesting contribution towards the inference of myocardial tissue behavior from full-field displacement fields and force measurements.
+In prior work, Joshi et al.
+demonstrated the feasibility of using Gibbs sampling for full-field inference in two-dimensional problems [32], where the material response could be inferred from a linear system of equations for the unknown material parameters in a rich constitutive feature library.
+In contrast, the setting of this study involves two distinct and compounding challenges.
+First, the underlying constitutive models are highly nonlinear [42, 14, 16], with strongly cross-correlated parameters [15, 43] that give rise to complex posterior landscapes.
+Second, we transition to three-dimensional full-field data.
 While the latter provides richer information, it also substantially increases the computational cost of evaluating the likelihood function.
 These factors independently and jointly limit the practical utility of Markov Chain Monte Carlo (MCMC) methods, including Gibbs sampling, due to the high cost of repeated likelihood evaluations and the poor mixing behavior in high-dimensional, correlated spaces.
-These limitations motivate us to extend the Bayesian EUCLID framework with more scalable inference strategies, such as Stochastic Variational Inference (SVI), which approximates posterior distributions through optimization rather than sampling, enabling tractable inference in settings with both complex models and large data volumes.
+These limitations motivates adapting the Bayesian EUCLID framework with more scalable inference strategies, such as Stochastic Variational Inference (SVI), which approximates posterior distributions through optimization rather than sampling, enabling tractable inference in settings with both complex models and large data volumes.
 Towards these goals, this study presents an unsupervised, full-field, Bayesian framework for the inference of orthotropic hyperelastic material parameters from a single heterogeneous biaxial tensile test.
-Our approach builds on the Bayesian-EUCLID framework (Joshi et al., 2022), extended via SVI (Thorat et al., 2025) to capture parametric uncertainty in a highly nonlinear, and a-priori chosen, constitutive model form.
-The myocardium is modeled using the well-established Holzapfel–Ogden constitutive law (Holzapfel and Ogden, 2009; Avazmohammadi et al., 2019), and heterogeneity in the deformation field is introduced either through microstructural variation (i.e.
-intrinsic fiber-sheet-interlaminar architectural heterogeneity), geometric tissue sample modification, or both.
+Our approach builds on the Bayesian-EUCLID framework [32], extended via SVI to capture parametric uncertainty in a highly nonlinear, a priori chosen constitutive model form.
+The myocardium is modeled using the well-established Holzapfel–Ogden constitutive law [14, 44], and heterogeneity in the deformation field is introduced either through microstructural variation (i.e., intrinsic fiber-sheet-interlaminar architectural heterogeneity), geometric tissue sample modification, or both.
 We demonstrate that the spatial richness of the heterogeneous displacement field combined with reaction force data enables robust parameter inference with uncertainty quantification from a single test.
 We additionally validate the inferred material models using unseen invariant values and assess sensitivity to displacement noise.
-This work highlights the potential of full-field, unsupervised, and uncertainty-aware approaches to redefine how we efficiently characterize complex living tissues, taking into account minimal tissue availability and minimizing tissue manipulation.
+Overall, this work highlights the potential of full-field, unsupervised, and uncertainty-aware approaches to redefine how we efficiently characterize complex biological tissues, taking into account minimal tissue availability and minimizing tissue manipulation.
 2 Unsupervised Stochastic Variational constitutive parameter inference
 Conventional approaches for characterizing myocardial material behavior typically rely on spatially averaged measurements and simplifying assumptions of spatially homogeneous microstructural organization.
-However, this intrinsic homogenization obscures the true spatial heterogeneity present within a tissue sample.
-To address this, we infer the constitutive parameter vector in an unsupervised setting from experimentally accessible data: (i) readily available global reaction forces, (ii) full-field deformation, and (iii) the microstructural organization across the slab.
-In this work, we assume digital volume correlation (DVC) data derived from concomitant through-thickness imaging, such as ultrasound (Navy et al., 2025), Magnetic Resonance Imaging (Estrada et al., 2020), or X-Ray Imaging (Davis et al., 2024) to obtain three-dimensional deformation fields.
-Microstructural maps in (iii) can be obtained via SHG or micro-CT (Alberini et al., 2024; Kakaletsis et al., 2021; Maes et al., 2022).
-Among established inverse strategies such as Finite Element Model Updating (FEMU) (Meijer et al., 1999), the Virtual Fields Method (VFM) (Grédiac et al., 2006), the Equilibrium Gap Method (EGM) (Claire et al., 2004; Roux and Hild, 2020), and Efficient Unsupervised Unsupervised Constitutive Law Identification (EUCLID) (Flaschel et al., 2021), we build on the latter.
+However, this intrinsic homogenization obscures the spatial heterogeneity present within a tissue sample.
+To address this, we infer the constitutive parameter vector in an unsupervised setting from experimentally accessible data: (i) global reaction forces, (ii) full-field deformation, and (iii) microstructural organization across the slab.
+In this work, we assume digital volume correlation (DVC) data derived from concomitant through-thickness imaging, such as ultrasound [45], Magnetic Resonance Imaging [46], or X-Ray Imaging [47] to obtain three-dimensional deformation fields.
+In parallel, microstructural maps can be obtained via SHG or micro-CT [48, 12, 49].
+Among established inverse constitutive identification strategies such as FEMU, VFM, and EGM, we build on EUCLID [30].
 EUCLID formulates the inverse problem via the weak form of momentum balance (see Section 2.1).
-Here, we extend it to a fully three-dimensional setting with nonlinear orthotropic materials, and recast the classic deterministic parameter optimization into an unsupervised Bayesian inference problem.
+Here, we adapt it to a fully three-dimensional setting with nonlinear orthotropic materials and recast the classic deterministic parameter optimization into an unsupervised Bayesian inference problem.
 Leveraging Stochastic Variational Inference (SVI), we identify orthotropic hyperelastic parameters from a single biaxial test; the overall workflow is summarized in Figure 1 and detailed in Section 2.2.
@@
 Quantification of the tissue’s microstructural organization, such as spatially varying fiber orientations, provides essential architectural information across the sample domain.
-In parallel, point-wise measurements of displacements $\hat{\bm{u}}$ and reaction forces $\hat{\bm{R}}^{\beta}$ are acquired under quasi-static biaxial loading using digital volume correlation.
+In parallel, point-wise measurements of displacements $\bm{u}$ and reaction forces $\bm{R}^{\beta}$ are acquired under quasi-static biaxial loading using digital volume correlation.
 The sample geometry is discretized into a finite element mesh, enabling reconstruction of continuous deformation fields $\bm{F}$ from the measured data.
@@
 Residuals are formulated through the weak form of the conservation of linear momentum, minimized pointwise for unconstrained degrees of freedom and in aggregate at boundaries with known reaction forces.
-The inverse problem is cast as a stochastic variational inference task, where a variational guide distribution $q(\bm{\theta})$ approximates the true posterior $\pi\left(\bm{\theta}\,|\,\hat{\bm{u}},\hat{\bm{R}}^{\beta}\right)$ by minimizing the Kullback-Leibler divergence $\mathrm{KL}\left(\pi\left(\bm{\theta}\,|\,\hat{\bm{u}},\hat{\bm{R}}^{\beta}\right)\,||\,q\left(\bm{\theta}\right)\right)$.
+The inverse problem is cast as a stochastic variational inference task, where a variational guide distribution $q(\bm{\theta})$ approximates the true posterior $\pi\left(\bm{\theta}\,|\,\bm{u},\bm{R}^{\beta}\right)$ by minimizing the Kullback-Leibler divergence $\mathrm{KL}\left(q\left(\bm{\theta}\right)\,||\,\pi\left(\bm{\theta}\,|\,\bm{u},\bm{R}^{\beta}\right)\right)$.
 Our physics-informed constitutive inference framework enables efficient and scalable inference from complex high dimensional full-field deformation datasets, accommodating nonlinear material behaviors and cross-correlated parameter effects.
 2.1 EUCLID preliminaries
+We first define the kinematic and force data available from a single experiment, and then derive the equilibrium residuals that underpin deterministic parameter identification.
 Consider a myocardial tissue specimen undergoing quasi-static deformation in a three-dimensional reference domain $\Omega\subset\mathbb{R}^{3}$.
-Dependent on the tissue slab’s extraction orientation protocol (Fig.
-3), the underlying myocardial microstruture variation induces diverse and heterogeneous strain states, while the material itself is assumed to be homogeneous.
-Optionally, additional tissue manipulations, e.g.
-cutting a hole in the tissue slab or local cardiac ablation, can further promote a highly heterogeneous deformation state during biaxial tensile testing.
+Depending on the tissue slab’s extraction orientation protocol (see Section  3), the underlying myocardial microstructure variation induces diverse and heterogeneous strain states.
+In this work, this heterogeneity enters through spatially varying structural fields, while the constitutive parameter vector $\bm{\theta}$ is assumed to be spatially homogeneous and thus constant over $\Omega$.
+This separation attributes spatial variability to the structural inputs while keeping the inverse problem focused on a single parameter vector.
+Optionally, additional tissue manipulations, e.g., cutting a hole in the tissue slab or local cardiac ablation, can further promote a highly heterogeneous deformation state during biaxial tensile testing.
 Boundary conditions are applied such that Dirichlet conditions are enforced on $\partial\Omega_{u}\subset\partial\Omega$ and Neumann conditions on the remainder, $\partial\Omega_{t}=\partial\Omega\setminus\partial\Omega_{u}$.
 For simplicity, our analysis focuses on displacement-controlled loading (i.e., Dirichlet boundary conditions), while noting that applied forces in load-controlled scenarios are equivalent to reaction forces under displacement control.
-The dataset comprises of $n_{t}$ snapshots of displacement measurements,
-$\mathcal{U}=\left\{\bm{u}^{a,t}\in\mathbb{R}^{3}:a=1,\dots,n_{n};\ t=1,\dots,n_{t}\right\},$ (1)
-recorded at $n_{n}$ reference points
-$\mathcal{X}=\left\{\bm{X}^{a}\in\Omega:a=1,\dots,n_{n}\right\}.$ (2)
+The dataset comprises $n_{t}$ snapshots of displacement measurements,
+$\left\{\bm{u}^{a,t}\in\mathbb{R}^{3}:a=1,\dots,n_{n};\ t=1,\dots,n_{t}\right\},$ (1)
+recorded at $n_{n}$ reference points comprising the measurement node set:
+$\left\{\bm{X}^{a}\in\Omega:a=1,\dots,n_{n}\right\}.$ (2)
 Additionally, for each snapshot, $n_{\beta}$ reaction forces
-$\left\{R^{\beta,t}:\beta=1,\dots,n_{\beta};\ t=1,\dots,n_{t}\right\}$ (3)
+$\left\{R^{\beta,t}\in\mathbb{R}:\beta=1,\dots,n_{\beta};\ t=1,\dots,n_{t}\right\}$ (3)
 are measured at selected Dirichlet boundaries (e.g., using load cells).
-Since $n_{\beta}\ll n_{n}$, the force data is sparse compared to the dense displacement field.
+Since $n_{\beta}\ll n_{n}$, the reaction force data are sparse compared to the dense displacement field.
 For brevity, the superscript $(\cdot)^{t}$ is omitted in the subsequent discussion, although the numerical procedure is applied independently to every snapshot.
-Given these limited measurements, the primary goal is to infer the a priori assumed constitutive model $\psi(\bm{F},\bm{\theta})$’s parameter set $\bm{\theta}$ governing the stress–strain response.
-To reconstruct the displacement field, the reference domain $\mathcal{X}$ is discretized using linear tetrahedral elements (each with a single quadrature point at its barycenter),
+Given $\left\{\bm{u}^{a}\right\}$ and $\left\{R^{\beta}\right\}$, our goal is to infer the parameter set $\bm{\theta}$ of an a priori assumed constitutive model $\psi(\bm{F},\bm{\theta})$.
+To connect the measurements to equilibrium, we first reconstruct a continuous kinematic field and its deformation gradient.
+The reference domain $\Omega$ is discretized using linear tetrahedral elements, evaluated with a single quadrature point at its barycenter, yielding
 $\bm{u}(\bm{X})=\sum_{a=1}^{n_{n}}N^{a}(\bm{X})\ \bm{u}^{a}.$ (4)
-Here, $N^{a}:\Omega\to\mathbb{R}$ represents the shape function associated with node $\bm{X}^{a}$. We subsequently approximate the corresponding deformation gradient field as
+Here, $N^{a}:\Omega\to\mathbb{R}$ represents the shape function associated with the measurement node $\bm{X}^{a}$. We subsequently approximate the corresponding deformation gradient field as
 $\bm{F}(\bm{X})=\bm{I}+\sum_{a=1}^{n_{n}}\bm{u}^{a}\otimes\nabla N^{a}(\bm{X}),$ (5)
 where $\bm{I}$ is the identity matrix, and $\nabla$ represents the gradient operator with respect to the reference coordinates $\bm{X}$.
-Following Flaschel et al.
-(2021) and Thakolkaran et al.
-(2022), we leverage the conservation of linear momentum to guide the learning of the constitutive model parameter set, which eliminates the need of stress labels.
+Assuming hyperelasticity, i.e., no dissipative energy losses within the material, the first Piola Kirchhoff stress follows as
+$\bm{P}(\bm{F})=\frac{\partial\psi(\bm{F},\bm{\theta})}{\partial\bm{F}}$ (6)
+as the constitutive relation between the first Piola Kirchhoff stress $\bm{P}$ and the deformation gradient $\bm{F}$ defined through the free energy function $\psi$ and parameterized using the constitutive parameter set $\bm{\theta}$ [50, 51, 52].
+In particular, $\bm{P}$ depends on the unknown parameters through $\psi(\bm{F},\bm{\theta})$.
+We leverage the conservation of linear momentum to guide the learning of the constitutive model parameter set, which eliminates the need for stress labels [30, 31].
 Assuming quasi-static loading conditions and negligible body forces, the weak form of the linear momentum balance in our reference domain $\Omega$ is given by
-$\int_{\Omega}\bm{P}(\bm{F}):\nabla\bm{v}\,\mathrm{d}V-\int_{\partial\Omega_{t}}\hat{\bm{t}}\cdot\bm{v}\,\mathrm{d}S=0\quad\forall\ \text{admissible }\bm{v}\,$, (6)
-where $\hat{\bm{t}}$ denotes the prescribed traction acting on $\partial\Omega_{t}$ and $\bm{v}$ is a test function that is sufficiently regular and vanishes on the Dirichlet boundary $\partial\Omega_{u}$.
-We prefer the weak formulation over the strong form, as it avoids the need for second order derivatives with respect to the displacement-field, which are sensitive to measurement noise.
-Assuming no dissipative energy losses within the material and rewriting the Clausius–Duhem entropy inequality Planck (1897) following the Coleman and Noll principle Coleman and Noll (1959); Peirlinck et al.
-(2024), we get
-$\bm{P}(\bm{F})=\frac{\partial\psi(\bm{F},\bm{\theta})}{\partial\bm{F}}$ (7)
-as the constitutive relation between the first Piola-Kirchoff stress $\bm{P}$ and the deformation gradient $\bm{F}$ defined through the free energy function $\psi$ and parametrized using the constitutive parameter set $\bm{\theta}$.
+$\int_{\Omega}\bm{P}(\bm{F},\bm{\theta}):\nabla\bm{v}\,\mathrm{d}V-\int_{\partial\Omega_{t}}\bar{\bm{t}}\cdot\bm{v}\,\mathrm{d}S=0\quad\forall\ \text{admissible }\bm{v}\,$, (7)
+where $\bar{\bm{t}}$ denotes the prescribed traction acting on $\partial\Omega_{t}$ per unit reference area and $\bm{v}$ is a test function that is sufficiently regular and vanishes on the Dirichlet boundary $\partial\Omega_{u}$.
+We prefer the weak formulation over the strong form, as it avoids the need for second order derivatives with respect to the displacement field, which are sensitive to measurement noise.
+In our displacement-controlled setting, non-instrumented free surfaces are typically traction free (so $\bar{\bm{t}}=0$ there).
+Measured boundary forces enter through reactions on $\partial\Omega_{u}$, which we incorporate below.
 Let $\mathcal{D}=\{(a,i):a=1,\dots,n_{n};\ i=1,2,3\}$ denote all the displacement degrees of freedom in our domain $\Omega$, which we partition into $\mathcal{D}^{\text{free}}$, the unconstrained degrees of freedom, and $\mathcal{D}^{\text{fix}}_{\beta}$ (with $\beta=1,\dots,n_{\beta}$) denoting the degrees of freedom under Dirichlet constraints that contribute to the observed reaction force $R^{\beta}$.
-Approximating the test function $\bm{v}$ by
+Using a Bubnov–Galerkin discretization, we approximate the admissible test function $\bm{v}$ as
 $\bm{v}(\bm{X})=\sum_{a=1}^{n_{n}}N^{a}(\bm{X})\,\bm{v}^{a},\qquad\text{with}\quad v_{i}^{a}=0\quad\forall\ (a,i)\in\bigcup_{\beta=1}^{n_{\beta}}\mathcal{D}^{\text{fix}}_{\beta}\,$, (8)
 and introducing our constitutive relation (6), our weak form (7) reduces to
-$\sum_{a=1}^{n_{n}}v_{i}^{a}f_{i}^{a}=0\,,\quad\text{with}\quad f_{i}^{a}=\underbrace{\int_{\Omega}P_{ij}\,\nabla_{j}N^{a}\,\mathrm{d}V}_{\text{internal force}}-\underbrace{\int_{\partial\Omega_{t}}\hat{t}_{i}N^{a}\,\mathrm{d}S}_{\text{external force}}\,.$ (9)
-Here, the integrals are computed using numerical quadrature over the same discretization and mesh as (4) resulting in a Bubnov-Galerkin form.
-Since the test functions are arbitrary, the force residual must vanish at every unconstrained degree of freedom,
+$\sum_{a=1}^{n_{n}}v_{i}^{a}f_{i}^{a}=0\,,\quad\text{with}\quad f_{i}^{a}=\underbrace{\int_{\Omega}P_{ij}\,\nabla_{j}N^{a}\,\mathrm{d}V}_{\text{internal force}}-\underbrace{\int_{\partial\Omega_{t}}\bar{t}_{i}N^{a}\,\mathrm{d}S}_{\text{external force}}\,.$ (9)
+The integrals are evaluated by numerical quadrature on the same discretization as (4).
+Since the test functions are arbitrary on $\mathcal{D}^{\text{free}}$, the force residual must vanish at every unconstrained degree of freedom,
 $f_{i}^{a}=0\quad\forall\ (a,i)\in\mathcal{D}^{\text{free}}\,.$ (10)
 At the fixed degrees of freedom, the internal and external forces are counteracted by the reaction force imposed by the Dirichlet constraints.
-Since point-wise reaction forces cannot be measured experimentally, they are assumed to be unavailable.
-Instead, only the total reaction forces integrated over the boundary segments are known.
-As a result, the global force balance for each measured reaction force is expressed as
+Because point-wise reactions are not accessible experimentally, we treat them as unavailable.
+Instead, we use the measured reaction forces integrated over instrumented boundary segments, which yields the global balance constraint
 $\sum_{(a,i)\in\mathcal{D}^{\text{fix}}_{\beta}}f^{a}_{i}=R^{\beta}\qquad\forall\qquad\beta=1,\dots,n_{\beta},$ (11)
 where the summation is carried out over all point-wise forces associated with the degrees of freedom in the $\beta^{\text{th}}$ Dirichlet constraint, denoted as $\mathcal{D}^{\text{fix}}_{\beta}$.
-As noted earlier, the superscript $(\cdot)^{t}$ has been omitted for brevity, but the above force balance conditions apply to all data snapshots at $t=1,\dots,n_{t}$.
-Together, (10) and (11) constitute a system of equations, which defines the inverse problem of constitutive model inference (Flaschel et al., 2021; Thakolkaran et al., 2022, 2025) as the minimization of the force balance residuals with respect to the trainable parameter set $\bm{\theta}$ of the free energy function $\psi_{\bm{}}$ (7):
-$\bm{\hat{\theta}}\leftarrow\arg\min_{\bm{\theta}}\sum_{t=1}^{n_{t}}\Bigg[\underbrace{\sum_{(a,i)\ \in\ \mathcal{D}^{\text{free}}}\left(f_{i}^{a,t}\right)^{2}}_{\text{free degrees of freedom}}+\underbrace{\sum_{\beta=1}^{n_{\beta}}\Bigl(R^{\beta,t}-\sum_{(a,i)\ \in\ \mathcal{D}^{\text{fix}}_{\beta}}f_{i}^{a,t}\Bigr)^{2}}_{\text{fixed degrees of freedom}}\Bigg]\,$, (12)
+Together, (10) and (11) constitute a system of equations that defines the inverse problem of constitutive model inference [30, 31, 33].
+We deterministically solve for $\bm{\theta}^{\star}$ by minimizing the equilibrium residuals over all snapshots:
+$\bm{\theta}^{\star}\leftarrow\arg\min_{\bm{\theta}}\sum_{t=1}^{n_{t}}\Bigg[\underbrace{\sum_{(a,i)\ \in\ \mathcal{D}^{\text{free}}}\left(f_{i}^{a,t}\right)^{2}}_{\text{free degrees of freedom}}+\underbrace{\sum_{\beta=1}^{n_{\beta}}\Bigl(R^{\beta,t}-\sum_{(a,i)\ \in\ \mathcal{D}^{\text{fix}}_{\beta}}f_{i}^{a,t}\Bigr)^{2}}_{\text{fixed degrees of freedom}}\Bigg]\,$, (12)
 such that the displacement and reaction force data satisfy the physics-based constraints (10) and (11).
 2.2 Stochastic variational inference
-In this study, we assume an a priori known constitutive law (see Eq.
-27) and aim to stochastically infer the underlying parameter set $\bm{\theta}$.
-Leveraging Bayes’ theorem, we express both our prior beliefs about the parameter set and the influence of noisy measurements within a unified probabilistic framework.
-We recast the deterministic optimization problem in (12) into Bayesian inference by treating the force residuals as the result of additive measurement noise (both for displacements and reaction forces).
-Let $\bm{\rho}(\bm{\theta})$ denote the global residual obtained by concatenating contributions from all load steps $t=1,\ldots,n_{t}$; for brevity, we omit explicit superscripts $(\cdot)^{t}$ when the meaning is clear from context.
-The residual stacks two components: (a) the internal-force equilibrium at unconstrained degrees of freedom and (b) the reaction-force consistency at fixed, i.e.
-instrumented, boundary segments:
-$\bm{\rho}(\bm{\hat{\theta}})=\begin{bmatrix}\{f_{i}^{a}\}_{(a,i)\in\mathcal{D}^{\text{free}}}\\ \{R^{\beta}-\sum_{(a,i)\in\mathcal{D}^{\text{fix}}_{\beta}}f_{i}^{a}\}_{\beta=1}^{n_{\beta}}\end{bmatrix}$ (13)
-Note that $\bm{\rho}(\bm{\theta})$ implicitly depends on observed displacements through the predicted forces in $f_{i}^{a}$, and thus the material model parameters $\bm{\theta}$.
-We assume that these residuals are distributed according to a multivariate normal distribution with zero mean and isotropic variance, reflecting the assumption of additive Gaussian noise:
-$\bm{\rho}(\bm{\hat{\theta}})\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})$ (14)
-In our case, both force and displacement measurements are affected by sensor noise.
-However, because displacements propagate in nonlinear fashion to the force residuals, the uncertainty introduced by the displacement noise is assumed to dominate.
-Under this assumption, the likelihood of the observed displacement $\hat{\bm{u}}$ and reaction force $\hat{R}^{\beta}$ data, given the parameters $\bm{\theta}$ and noise variance $\sigma^{2}$, becomes:
-$\pi\left(\hat{\bm{u}},\hat{R}^{\beta}\mid\bm{\theta},\sigma^{2}\right)\,\propto\,\exp\left(-\frac{1}{2\sigma^{2}}\|\,\bm{\rho}(\bm{\theta})\,\|^{2}\right)$ (15)
-Again, for brevity, the superscript $(\cdot)^{t}$ is omitted in the subsequent discussion, although the likelihood distribution is computed as the product of the likelihood distribution for the residual computed at each independent quasi-static loading time snapshot.
-Next, we place priors on the the noise variance $\sigma^{2}$ and the parameter vector $\bm{\theta}$ respectively to complete our Bayesian treatment of the inference problem.
-Because the variance $\sigma^{2}$ of the measurement noise is unknown and must be inferred along with the parameters, we place an inverse-gamma prior on it:
+In this study, we assume an a priori known constitutive law (28) and aim to infer the underlying constitutive parameter set $\bm{\theta}$ together with its uncertainty.
+Leveraging Bayes’ theorem, we combine prior beliefs about $\bm{\theta}$ with a likelihood model that accounts for the impact of measurement noise on the physics-based residual constraints conditioned on the displacement and reaction force data.
+We recast the deterministic optimization problem in (12) into Bayesian inference by treating the equilibrium residuals as the result of additive measurement noise.
+For a given snapshot $t=1,\ldots,n_{t}$, we define the residual vector
+$\bm{\rho}^{t}(\bm{\theta})=\begin{bmatrix}\left\{f_{i}^{a,t}\right\}_{(a,i)\in\mathcal{D}^{\text{free}}}\\ \left\{R^{\beta,t}-\sum_{(a,i)\in\mathcal{D}^{\text{fix}}_{\beta}}f_{i}^{a,t}\right\}_{\beta=1}^{n_{\beta}}\end{bmatrix}$, (13)
+and stack all snapshots into a single global residual
+$\bm{\rho}(\bm{\theta})=\left[\begin{array}[]{c}\bm{\rho}^{1}(\bm{\theta})\\ \vdots\\ \bm{\rho}^{n_{t}}(\bm{\theta})\end{array}\right]$ (14)
+This construction removes the need to carry superscript $(\cdot)^{t}$ in the subsequent probabilistic expressions while still retaining the multi-snapshot information.
+Note that $\bm{\rho}(\bm{\theta})$ implicitly depends on observed displacements and reaction forces through the constitutive response leading to the predicted forces in $f_{i}^{a}$.
+We assume that the stacked residuals follow a multivariate normal distribution with zero mean and isotropic variance, reflecting the assumption of additive Gaussian noise:
+$\bm{\rho}(\bm{\theta})\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})$ (15)
+which yields the likelihood
+$\pi\left(\bm{u},R^{\beta}\mid\bm{\theta},\sigma^{2}\right)\,\propto\,\exp\left(-\frac{1}{2\sigma^{2}}\|\,\bm{\rho}(\bm{\theta})\,\|^{2}\right)$ (16)
+In our case, both displacement field and reaction force measurements are affected by sensor noise, and we treat $\sigma^{2}$ as an effective residual variance that captures this combined uncertainty.
+Given that displacement propagates in a nonlinear fashion to the force residuals, we expect the residual variance to be dominated by displacement noise.
+Next, we place priors on the residual noise variance $\sigma^{2}$ and the parameter vector $\bm{\theta}$.
+Since the variance of the measurement noise $\sigma^{2}$ is unknown and must be inferred, we assign an inverse-gamma prior:
 $\sigma^{2}\sim\mathcal{IG}\,\left(\alpha_{\sigma},\beta_{\sigma}\right)$ (17)
-Concomitantly, we enforce physical constraints on the parameters by requiring them to be positive and therefore enforce a truncated normal distribution for the prior.
-Here, we introduce a scaling hyperprior $V_{s}$ which allows us to decouple the overall measurement residual noise variance $\sigma$ from the underlying parameter variance.
+Concomitantly, we enforce physical constraints on the parameters by requiring them to be positive and therefore enforce a truncated normal distribution for the prior (truncated on the left at zero).
+Here, we introduce a scaling hyperprior $v_{s}$ to decouple the overall measurement residual noise variance $\sigma^{2}$ from the underlying parameter prior variance:
 $\bm{\theta}\sim\mathcal{N}_{+}\,\left(\bm{0},v_{s}\,\sigma^{2}\bm{I}\right)$ (18)
-For this scaling hyperprior, we again enforce parameter positivity by setting an inverse gamma prior distribution:
+We place an inverse-gamma hyperprior on $v_{s}$:
 $v_{s}\sim\mathcal{IG}\,\left(\alpha_{v_{s}},\beta_{v_{s}}\right)$ (19)
-Combining the prior distributions (16) – (18) with our likelihood function (15) yields the posterior:
-$\pi\left(\bm{\theta},\sigma^{2},v_{s}\mid\hat{\bm{u}},\hat{R}^{\beta}\right)\,\propto\,\pi\left(\hat{\bm{u}},\hat{R}^{\beta}\mid\bm{\theta},\sigma^{2}\right)\,\pi\left(\bm{\theta}\mid\sigma^{2},v_{s}\right)\,\pi\left(\sigma^{2}\right)\,\pi\left(v_{s}\right),$ (19)
-This posterior serves as the foundation for either sampling-based inference or variational inference approaches, chosen according to model complexity and data volume.
+The hierarchical structure [32, 53, 54] of our prior distributions is showcased in Figure 2.
+Figure 2: Schematic of the hierarchical inference model.
+This hierarchical Bayesian model places priors on both the residual variance in the likelihood and the constitutive model parameters.
+The red circular distributions represent hyperpriors for $v_{s}$ and $\sigma^{2}$, the light-blue circular distribution represents the prior on the material model parameters, and the black rectangular distribution represents the likelihood induced by the linear momentum balance residuals.
+The black arrows denote hierarchical dependencies.
+The hyperpriors are inverse-gamma distributed, the consitutive parameters are distributed according to a truncated multivariate normal distribution, and the likelihood is assumed to be normally distributed.
+Combining the prior distributions (17)–(19) with the likelihood (16) yields the posterior:
+$\pi\left(\bm{\theta},\sigma^{2},v_{s}\mid\bm{u},R^{\beta}\right)\,\propto\,\pi\left(\bm{u},R^{\beta}\mid\bm{\theta},\sigma^{2}\right)\,\pi\left(\bm{\theta}\mid\sigma^{2},v_{s}\right)\,\pi\left(\sigma^{2}\right)\,\pi\left(v_{s}\right),$ (20)
+This posterior can, in principle, be approximated using sampling-based inference or variational inference.
 In simpler inference problems, sampling-based methods such as MCMC can often be relied on to approximate the posterior distribution.
-For example, Gibbs sampling has been successfully applied to two-dimensional full-field problems, where the material response could be inferred by solving a linear system of equations on a library of constitutive features (Joshi et al., 2022).
+For example, Gibbs sampling has been successfully applied to two-dimensional full-field problems, where the material response could be inferred by solving a linear system of equations on a library of constitutive features [32].
 In those cases, the constitutive models were potentially nonlinear in input but remained linear in their parameters, which keeps the posterior geometry relatively tractable for MCMC.
 In our setting, however, the constitutive models are nonlinear in both inputs and parameters, which induces strong parameter interactions and complex (often multimodal) posteriors.
-Together with the cost of evaluating data-heavy 3D full-field likelihoods, this makes traditional MCMC prohibitively expensive.
-In turn, we use variational inference to approximate the posterior distribution $\pi$ with a more tractable family of parametric guide distributions $q$.
-We therefore adopt stochastic variational inference (SVI) to approximate the posterior efficiently, enabling gradient-based optimization such as stochastic gradient descent and ADAM (Kingma and Welling, 2013; Kingma and Ba, 2017).
-SVI has been shown to substantially accelerate inference relative to MCMC (Rodrigues, 2022; Thorat et al., 2025).
-Additionally, we employ a mean-field variational family $q$ (Parisi, 1988; Consonni and Marin, 2007), which assumes statistical independence between the quantities of interest.
+Together with the cost of evaluating data-heavy 3D full-field likelihoods, this makes conventional MCMC-based posterior exploration computationally unattractive in practice.
+We therefore adopt stochastic variational inference (SVI) to approximate the posterior efficiently using gradient-based optimization.
+SVI has been shown to substantially accelerate inference relative to MCMC in related settings [55, 53].
+In SVI, we approximate the posterior distribution $\pi$ with a tractable family of parametric guide distributions $q$.
+We employ a mean-field variational family $q$ [56, 57], which assumes statistical independence between the quantities of interest,
 $q\left(\bm{\theta},\,v_{s},\,\sigma^{2}\right)=q\left(\bm{\theta}\right)\,q\left(\sigma^{2}\right)\,q\left(v_{s}\right)$ (21)
-where each factor is defined as:
-$q\left(\bm{\theta}\right)\quad$ $=$ $\quad\mathcal{N}_{+}(\bm{\mu}^{q},\bm{\Sigma}^{q}\circ\bm{I})$ (21)
-$q\left(\sigma^{2}\right)\quad$ $=$ $\quad\mathcal{IG}(\alpha_{\sigma}^{q},\beta_{\sigma}^{q})$ (22)
-$q\left(v_{s}\right)\quad$ $=$ $\quad\mathcal{IG}(\alpha_{vs}^{q},\beta_{vs}^{q})$ (23)
-and $\circ$ denotes the element-wise (Hadamard) product.
-The full set of variational parameters is denoted by $\bm{\Phi}$ and our hyperparameters are reported in Table 1.
-The hierarchical structure (Joshi et al., 2022; Thorat et al., 2025; Peirlinck et al., 2019) of our prior distributions is showcased in Figure 2.
-Figure 2: Schematic of the hierarchical inference model.
-This hierarchical structure of the Bayesian inference model allows for simulteneous placement of prior distributions on the noise in the likelihood and the constitutive model parameters.
-The red and circular distributions represent hyperpriors for $V_{s}$ and $\sigma^{2}$, the lightblue circular distribution represent the prior on the material model parameters and the black and rectangular distribution represents the likelihood of the linear momentum balance.
-The black arrows denote the hierarchical structure dependencies.
-The hyperpriors are distributed according to inverse-gamma distributions.
-The model parameters are distributed according to a truncated multivariate normal distribution.
-Truncation occurs at numerical zero to enforce polyconvexity in the material model.
-The likelihood is assumed to be normally distributed.
-To find the best approximation within this family, we minimize the Kullback-Leibler (KL) divergence (Kullback, 1968; Cover T., 1968) between the variational distribution $q\left(\bm{\Phi}\right)$ and the true posterior $\pi\left(\bm{\Phi}\mid\hat{\bm{u}},\hat{R}^{\beta}\right)$, where $\bm{\Phi}=\{\bm{\theta},v_{s},\sigma^{2}\}$:
-$\mathrm{KL}\left[\pi(\bm{\Phi}\mid\hat{\bm{u}},\hat{R}^{\beta})\mid\mid q(\bm{\Phi})\right]=\mathbb{E}_{q(\bm{\Phi})}\left[-\log q\left(\bm{\Phi}\right)-\log\pi\left(\bm{\Phi}\right)-\log\pi\left(\hat{\bm{u}},\hat{R}^{\beta}\mid\bm{\Phi}\right)\right]+\log\pi\left(\hat{\bm{u}},\hat{R}^{\beta}\right)$ (24)
-The entire inference pipeline is implemented using NumPyro (Phan et al., 2019), which supports automatic differentiation and stochastic optimization with ADAM using the reparameterization trick (Kingma and Ba, 2017; Kingma and Welling, 2013), which alleviates the need to explicitly define the minimization procedure as was done in (Thorat et al., 2025).
-We initialize the optimization method with a learning rate of $r_{0}=0.01$ and use exponential learning rate decay $\gamma$.
-The learning rate is defined for each epoch $i$ as $r_{i}=r_{0}\gamma^{i/n_{T}}$.
+with factors:
+$q\left(\bm{\theta}\right)\quad$ $=$ $\quad\mathcal{N}_{+}\left(\bm{\mu}^{q},\bm{\Sigma}^{q}\circ\bm{I}\right)$ (22)
+$q\left(\sigma^{2}\right)\quad$ $=$ $\quad\mathcal{IG}\left(\alpha_{\sigma}^{q},\beta_{\sigma}^{q}\right)$ (23)
+$q\left(v_{s}\right)\quad$ $=$ $\quad\mathcal{IG}\left(\alpha_{vs}^{q},\beta_{vs}^{q}\right),$ (24)
+where $\circ$ denotes the element-wise Hadamard product.
+The full set of variational parameters is denoted by $\bm{\Phi}=\left\{\bm{\mu}^{q},\bm{\Sigma}^{q},\alpha_{\sigma}^{q},\beta_{\sigma}^{q},\alpha_{vs}^{q},\beta_{vs}^{q}\right\}$, and hyperparameters are reported in Table 1.
+We specifically highlight $\bm{\mu}^{q}$, and $\bm{\Sigma}^{q}$ to be the location and scale parameters of the truncated normal distribution.
+These correspond to the mean and standard deviation of the underlying and untruncated normal distribution used in the definition of $\mathcal{N}_{+}$.
+To find the best approximation within this family, we minimize the Kullback-Leibler divergence [58, 59] between the variational distribution $q$ and the true posterior $\pi$, i.e.
+$\bm{\Phi}^{\star}\leftarrow\arg\min_{{\bm{\Phi}}}\operatorname{KL}\left(q_{\bm{\Phi}}\left(\bm{\theta},v_{s},\sigma^{2}\right)\|\,\pi\left(\bm{\theta},v_{s},\sigma^{2}\mid\bm{u},R^{\beta}\right)\right),$ (25)
+The entire inference pipeline is implemented in NumPyro [60], enabling automatic differentiation and stochastic optimization with ADAM using the reparameterization trick [61, 62].
+Doing so, we alleviate the need to explicitly define the minimization procedure as was done in Thorat et al.
+[53].
+We initialize the optimizer with a learning rate of $r_{0}=0.01$, and apply an exponential learning rate decay $\gamma$, i.e.
+$r_{i}=r_{0}\gamma^{i/n_{T}}$ at epoch i with $n_{T}$ denoting the total number of epochs.
 Table 1: Hyperparameters and simulation settings used in the stochastic variational inference framework and synthetic tissue slab data generation.
@@
   Initial variance of $\mathcal{N}_{+}$ distribution $\sigma^{2}_{\theta}$ $10^{-18}\times\{1,\dots,1\}$
-  Initial scale parameter for $v_{s}$ $\alpha_{vs}$ $10^{4}$
-  Initial rate parameter for $v_{s}$ $\beta_{vs}$ $10^{4}$
+  Initial scale parameter for $v_{s}$ $\alpha_{v_{s}}$ $10^{4}$
+  Initial rate parameter for $v_{s}$ $\beta_{v_{s}}$ $10^{4}$
   Initial scale parameter for $\sigma^{2}$ $a_{\sigma^{2}}$ $10^{4}$
@@
   Linear scaling parameters $a_{\star}$ [kPa] $a,a_{f},a_{n},a_{fs}$ $\{0.809,1.911,0.227,0.547\}$
-  Exponential scaling parameters $_{\star} [][-] b,b$$b_{f},b_{n},b_{fs}$ $\{7.474,22.063,34.802,5.691\}$
+  Exponential scaling parameters $_{\star} [-] b,b$$b_{f},b_{n},b_{fs}$ $\{7.474,22.063,34.802,5.691\}$
   Bulk modulus [MPa] $K$ $0.1$
@@
 3.1 Myofiber architecture and tissue slab slicing
-Informed by prior histological characterizations of myocardial tissue, we model our studied specimens as a layered orthotropic material in which cardiomyocytes align along main myofiber directions, packed by endomysial collagen into sheets, and connected via perimysial collagen (Rohmer et al., 2007; Holzapfel and Ogden, 2009; Alexander J.
-Wilson, 2022).
-This organization defines three locally orthogonal directions: the fiber direction $\bm{f}$, the sheet direction $\bm{s}$, and the sheet-normal direction $\bm{n}=\bm{f}\times\bm{s}$.
-To investigate the role of heterogeneity in unsupervised constitutive parameter identifiability, we simulate myocardial tissue slabs that exhibit both microstructure and geometry induced heterogeneity.
+Informed by prior histological characterizations of myocardial tissue [8, 9, 14], we model our studied specimens as a layered orthotropic material in which cardiomyocytes align along main myofiber directions, packed by endomysial collagen into sheets, and connected via perimysial collagen.
+This organization defines three locally orthogonal directions: the fiber direction $\bm{f}$, the sheet direction $\bm{s}$, and the sheet-normal direction $\bm{n}$.
+To investigate the role of heterogeneity in unsupervised constitutive parameter identifiability, we simulate myocardial tissue slabs that exhibit both microstructure-induced and geometry-induced heterogeneity.
 Figure 3: Specimen slicing and microstructural organization.
@@
 Fiber directions ($\bm{f}$) are shown in red, sheet directions ($\bm{s}$) in pink, and normal directions ($\bm{n}$) in brown.
-Leveraging the intrinsic spatially varying myofiber architecture in the human heart (Lombaert et al., 2012), we extract three microstructurally distinct synthetic tissue slabs from the left ventricular free wall.
-All samples are assumed to be $10\times 10\times 1$ mm3 in size, corresponding to representative slabs from a healthy left ventricular wall which is at least 10 mm thick.
-The sample geometry and applied boundary conditions are also visualized in Figure 1 - left.
-Figure 3 showcases the microstructural variability across the slab domain, with directional fiber $\bm{f}$, sheet $\bm{s}$ and normal $\bm{n}$ vectors discretized element-wise, depend on how the tissue slicing and extraction procedure is performed:
+Leveraging the intrinsic spatially varying myofiber architecture in the human heart [63], we extract three microstructurally distinct synthetic tissue slabs from the left ventricular free wall.
+All samples are assumed to be $10\times 10\times 1$ mm3 in size, representative of slabs from a healthy human left ventricular wall.
+The sample geometry and applied boundary conditions are summarized in A.
+Figure 3 showcases the microstructural variability across the slab domain, with directional fiber $\bm{f}$, sheet $\bm{s}$ and normal $\bm{n}$ vectors discretized element-wise, and their spatial distribution depends on the chosen slicing orientation:
   • 1.
@@
 We consider both aligned and off-axis variants depending on their angular deviation from the square edge.
-This slab specimen type aligns with the biaxial tension testing slabs studied by Sommer et al.
-(2015).
+This slab specimen type aligns with the biaxial tension testing slabs studied by [11].
   • 2.
@@
     rad-heli: samples based on a radial-helical cut.
-Given the natural transmural variation of the myofiber angle $\bm{f}$ from $60^{\circ}$ in the sub-epicardium to $60^{\circ}$ in the sub-endocardium (Lombaert et al., 2012), this configuration places all three microstructural directions ($\bm{f},\bm{s},\bm{n}$) approximately within the sample plane by rotating the sample an additional $45^{\circ}$ along the radial axis.
+Given the natural transmural variation of the myofiber angle $\bm{f}$ from $60^{\circ}$ in the sub-epicardium to $60^{\circ}$ in the sub-endocardium [63], this configuration places all three microstructural directions ($\bm{f},\bm{s},\bm{n}$) approximately within the sample plane by rotating the sample an additional $45^{\circ}$ along the radial axis.
 The resultant fiber orientation ranges from $15^{\circ}$ up to $105^{\circ}$, aligning the $\bm{f}$ fiber in the sample plane near the bottom edge of the sample and aligning the $\bm{n}$ fiber in the sample plane near the top edge.
-To investigate the influence of geometric heterogeneity, we introduce a second set of specimens with a central circular occlusion with a radius of $1$ mm, see Figure 1 - right.
-These heterogeneous samples introduce additional strain concentration zones that we hypothesize may improve the identifiability of model parameters under limited data conditions, e.g.
-a biaxial tensile test to characterize orthotropic material behavior.
+To investigate the influence of geometric heterogeneity, we introduce a second set of specimens with a central circular occlusion with a radius of $1$ mm, see A.
+These samples are expected to introduce additional deformation heterogeneity that we hypothesize may improve constitutive parameter identifiability under limited data conditions, e.g., when inferring orthotropic behavior from a single biaxial tensile test.
 3.2 Biaxial loading emulation
-We simulate a digital volume correlation experiment by modeling a biaxial loading protocol on the microstructurally and geometrically homogeneous and heterogeneous tissue slabs described above using the finite element method (FEM) implemented in ABAQUS/Standard (Dassault Systèmes Simulia Corp., 2025), see for example (Peirlinck et al., 2018).
-Our virtual tissue slabs are subjected to displacement-controlled symmetric biaxial loading, with loading ratios $\lambda_{1}:\lambda_{2}=(1:2),(1:1),(2:1)$ based on stretch.
-The maximum stretch applied to the samples is $15\%$ their initial length, which relates to maximal values used in experiments (Sommer et al., 2015).
+We simulate a digital volume correlation experiment by modeling a biaxial loading protocol on the microstructurally and geometrically homogeneous and heterogeneous tissue slabs described above using the finite element method (FEM) implemented in ABAQUS/Standard [64], see for example [26].
+Our virtual tissue slabs are subjected to displacement-controlled symmetric biaxial loading, with loading ratios $\lambda_{1}:\lambda_{2}=(1:2),(1:1),(2:1)$ defined in terms of the prescribed edge stretches along the two in-plane loading directions.
+The maximum stretch applied to the samples is $15\%$ of the initial edge length, consistent with maximal values used in experiments [11].
 We use linear tetrahedral elements to discretize the domain and record nodal displacements along with reaction forces for a sequence of $n_{t}$ load steps (see Figure 1)
-Leveraging the local volume change of deformation $J=\det\left(\bm{F}\right)$, we multiplicatively decompose the deformation gradient $\bm{F}$ (5) into its volumetric $\bm{F}^{\rm{vol}}$ and isochoric $\bar{\bm{F}}$ parts (Flory, 1961),
+To define the constitutive response, we use the local volume change $J=\det\left(\bm{F}\right)$ and multiplicatively decompose the deformation gradient $\bm{F}$ (5) into its volumetric $\bm{F}^{\rm{vol}}$ and isochoric $\bar{\bm{F}}$ parts [65],
 $\bm{F}=\bm{F}^{\rm{vol}}\bar{\bm{F}}\quad\mbox{with}\quad\bm{F}^{\rm{vol}}=J^{\frac{1}{3}}\bm{I}\quad\mbox{and}\quad\bar{\bm{F}}=J^{-\frac{1}{3}}\bm{F}.$ (26)
-and introduce the following deformation (pseudo-)invariants (Spencer, 1984; Menzel, 2004):
+We then introduce the following deformation (pseudo-)invariants [66, 67]:
 $\bar{I}_{1}$ $=$ $[\bar{\bm{F}}^{{\rm{T}}{}}\bar{\bm{F}}]:\bm{I}$ (27)
@@
 where $\bm{f}^{0}$, $\bm{s}^{0}$, $\bm{n}^{0}$ denote the local microstructure fiber, sheet, and normal unit vector orientations in the undeformed reference configuration.
-Sometimes these invariants are centered around zero, in which case the following notation is used: $K_{1}=I_{1}-3,K_{4ff}=I_{4ff}-1,K_{4nn}=I_{4nn}-1$.
-We adopt the orthotropic, compressible Holzapfel-Ogden (HO) material model (Holzapfel and Ogden, 2009; Guan et al., 2018) using the following strain energy density function:
+We adopt the orthotropic, compressible Holzapfel-Ogden (HO) material model [14, 15] using the following strain energy density function:
 $\psi(\bm{F},\bm{\theta})$ $=\frac{a}{2b}\left[\exp\left(b(\bar{I}_{1}-3)\right)-1\right]+\sum_{i\,=\,f,n}\frac{a_{i}}{2b_{i}}\left[\exp\left(b_{i}\left<\bar{I}_{4ii}-1\right>^{2}\right)-1\right]$ (28)
 $+\frac{a_{fs}}{2b_{fs}}\left[\exp\left(b_{fs}\bar{I}_{4fs}^{2}\right)-1\right]+\frac{K}{2}\left(\frac{J^{2}-1}{2}-\ln J\right)$
-to compute the First Piola Kirchoff stress tensor $\bm{P}$ following (7).
-This variation of the HO-model was implemented using the universal material subroutine (Peirlinck et al., 2024).
-An analytical form of the stress tensor $\bm{P}$ is provided in B.
+to compute the First Piola Kirchhoff stress tensor $\bm{P}$ following (6). This variation of the HO-model was implemented using the recently proposed universal user material subroutine in Abaqus [52].
 In (28), $\bm{\theta}=\{a,b,a_{f},b_{f},a_{n},b_{n},a_{fs},b_{fs},K\}$ (see Table 1) denotes the to-be-inferred ground-truth material parameters.
-The Macaulay brackets $\langle x\rangle$ ensure that fiber and sheet term strain energy contributions activate only under tension (Wittrick, 1965; Wriggers et al., 2016; Peirlinck et al., 2024).
-Lastly, a bulk modulus is chosen based on the distribution of $J$ in the circ-long-hom samples and common compressibility considerations of the myocardium (McEvoy et al., 2018; Yin et al., 1996).
+The Macaulay brackets $\langle x\rangle$ ensure that fiber and sheet term strain energy contributions activate only under tension [68, 69, 52].
+Lastly, a bulk modulus is chosen based on the distribution of $J$ in the circ-long-hom samples and common compressibility considerations of the myocardium [70, 71] to limit volumetric strains while retaining numerical stability.
+To quantify how microstructural and geometric heterogeneity enrich the deformation space relevant for inference, Figures 45 report the resulting distributions of the key invariants activated under biaxial loading, which motivates our single-test identification setup.
 Figure 4: Deformation profiles of geometrically homogeneous tissue slabs with varying microstructural heterogeneity.
 Maximum principal stretches are shown for circ-long-hom (top row), circ-rad-hom (middle row), and rad-heli-hom (bottom row) specimens subjected to biaxial loading protocols ($\lambda_{1}:\lambda_{2}=(1:2),(1:1),(2:1)$) up to $\lambda^{\text{max}}=1.15$.
-Rightmost columns display distributions of invariants $I_{1}-3$, $I_{4ff}-1$, $I_{4nn}-1$, and $I_{4fs}$ for each specimen type respectively using approximated kernel density estimations denoted as $\hat{\pi}(I_{*})$.
-The kernel density estimation is truncated at an arbitrary level of 20 to show the spread of the invariants more clearly.
+Rightmost columns display distributions of invariants $I_{1}-3$, $I_{4ff}-1$, $I_{4nn}-1$, and $I_{4fs}$ for each specimen type respectively using approximated kernel density estimations denoted as $\mathrm{KDE}(I_{*})$ The kernel density estimation is truncated at an arbitrary level of 20 to show the spread of the invariants more clearly.
 Figure 4 presents the deformation response of geometrically homogeneous slabs with varying microstructural architectures.
 The qualitative results of the maximum principal stretch are shown for all loading conditions and homogeneous tissue samples.
-On the right side of the figure we show kernel density estimations of the biaxially activated invariants - truncated at $20$ to visualize the spread of these invariants.
-The circ-long-hom configuration represents a classic homogenized myocardial sample, as excised by (Sommer et al., 2015), with assumed uniform fiber-sheet-normal orientations.
+On the right side of the figure we show kernel density estimations of the biaxially activated invariants, truncated at $20$ to visualize the spread of these invariants.
+The circ-long-hom configuration represents a classic homogenized myocardial sample, as excised by [11], with assumed uniform fiber-sheet-normal orientations.
 Under biaxial loading, this results in a uniform deformation pattern with isotropic invariants $I_{1}$ ranging from $3.158$ to $3.257$ across the three loading protocols.
 Anisotropic contributions are moderate, with $I_{4ff}$ and $I_{4nn}$ spanning ranges from $1.025$ to $1.139$ and $1.024$ to $1.144$ respectively.
-Fiber-sheet shear contributions remain negligible, with $I_{4fs}$ limited to 0 up to machine precision.
+Fiber-sheet shear contributions remain negligible, with $I_{4fs}$ limited to $0.000$ up to machine precision.
 Introducing transmural fiber variation in the circ-rad-hom slab increases local deformation heterogeneity.
-The isotropic invariant $I_{1}$ expands to a broader range of $3.160$ to $3.367$, while $I_{4ff}$ and $I_{4nn}$ reach $1.194$ and $1.252$, respectively.
-Notably, fiber-sheet shear contributions increase, with $I_{4fs}$ reaching $0.099$.
+The isotropic deformation invariant $I_{1}$ expands to a broader range of $3.160$ to $3.367$, while $I_{4ff}$ and $I_{4nn}$ reach $1.194$ and $1.252$, respectively.
+Notably, fiber-sheet shear contributions increase, with $I_{4fs}$ reaching up to $0.099$.
 In the rad-heli-hom specimen, which introduces a more asymmetric helical fiber distribution, deformation heterogeneity is further amplified.
@@
 In particular, the rad-heli-hom configuration shows noticeable thickness bulging coinciding with bands of reduced strain, where fibers align in-plane.
-These bands are visually distinct due to their contrasting strain magnitudes relative to surrounding regions.
-These results underline the capacity of microstructural heterogeneity alone, even in geometrically simple specimens, to enrich deformation diversity.
-Such diversity is critical for reliably inferring nonlinear orthotropic constitutive behavior, motivating their role as training configurations in this study.
+Overall, microstructure-induced heterogeneity alone enriches the activated invariant space under a single biaxial protocol, which motivates using these slabs as training configurations in this study.
 Figure 5: Deformation profiles of geometrically heterogeneous tissue slabs with varying microstructural heterogeneity.
 Maximum principal stretches circ-long-het (top row), circ-rad-het (middle row), and rad-heli-het (bottom row) specimens under biaxial loading ($\lambda_{1}:\lambda_{2}=(1:2),(1:1),(2:1)$).
-Invariant distributions of $I_{1}-3$, $\bar{I}_{4ff}-1$, $\bar{I}_{4nn}-1$, and $I_{4fs}$ are shown in the rightmost panels using approximated kernel density estimations denoted as $\hat{\pi}(I_{*})$.
+Invariant distributions of $I_{1}-3$, $\bar{I}_{4ff}-1$, $\bar{I}_{4nn}-1$, and $I_{4fs}$ are shown in the rightmost panels using approximated kernel density estimations denoted as $\mathrm{KDE}(I_{*})$ .̇.
+denoted as $\mathrm{KDE}(I_{*})$.
 The kernel density estimation is truncated at an arbitrary level of 20 to show the spread of the invariants more clearly.
@@
 These observations reinforce the rationale for using such heterogeneous configurations as training specimens, enhancing parameter identifiability through enriched mechanical diversity.
-4 Orthotropic constitutive parameter inference
+4 Inference results and robustness to measurement noise
+4.1 Noise-free inference: identifiability gains from microstructural and geometrical heterogeneity
 […]
-$e_{\theta}$ $=\frac{|\hat{\theta}-\theta_{\text{gt}}|}{\theta_{\text{gt}}}\qquad$ $\tilde{\theta}$ $=\frac{\hat{\theta}-\mu_{\hat{\theta}}}{\mu_{\hat{\theta}}}\qquad$ $s_{\tilde{\theta}}$ $=\frac{\sqrt{N(N-1)}}{N-2}\cdot\frac{\frac{1}{N}\sum_{n=1}^{N}(\tilde{\theta}_{n}-\mu_{\tilde{\theta}})^{3}}{\left(\frac{1}{N}\sum_{n=1}^{N}(\tilde{\theta}_{n}-\mu_{\tilde{\theta}})^{2}\right)^{3/2}}$ (28)
-with $N$ equaling the amount of samples drawn from the posterior distribution.
-To gain insight into the shape, mean, and spread of the approximated posterior distribution, we draw 1,0001{,}000 samples from the inferred guide distribution shown in the bottom rows of Figures 6 and 7 respectively.
-Figure 6: Orthotropic constitutive parameter inference from geometrically homogeneous tissue slabs with varying microstructure.
-From left to right: a circ-long-hom specimen, a circ-rad-hom specimen, and a rad-heli-hom specimen.
-From top to bottom: (i) evolution of inferred material parameters $\hat{\theta}$ over the epochs, (ii) evolution of relative parameter errors $e_{\theta}$, (iii) final relative errors at final epoch, and (iv) rescaled posterior distributions.
-Distributions are shown in grayscale when unbiased skewness $|s_{\theta}|<0.5$.
-Despite microstructural heterogeneity, geometric homogeneity limits identifiability of shear-related parameters ($a_{fs}$, $b_{fs}$), which remain poorly inferred.
-In contrast, isotropic and primary anisotropic parameters are consistently identified with high accuracy, exhibiting narrow and near-Gaussian posterior distributions.
-Increasing microstructural heterogeneity improves convergence rates and reduces inference error for shear parameters, though challenges persist for $b_{fs}$, particularly in the rad-heli-hom specimen.
-Focusing on geometrically homogeneous samples (Figure 6), we evaluated three configurations with increasing microstructural heterogeneity: fully homogeneous (circ-long-hom), symmetric transmural fiber rotation (circ-rad-hom), and asymmetric helical fiber variation (rad-heli-hom).
-For the circ-long-hom configuration, all parameters converged rapidly except for those associated with the fiber-sheet shear coupling parameters ($a_{fs}$, $b_{fs}$).
-Final relative errors were minimal for isotropic parameters ($e_{a}=4.77\times 10^{-5}$, $e_{b}=-2.30\times 10^{-6}$) and remained low for anisotropic terms in $f$ and $n$ directions (i.e., $e_{a_{f}}=-5.50\times 10^{-5}$, $e_{b_{f}}=-1.47\times 10^{-6}$, $e_{a_{n}}=9.66\times 10^{-5}$ , $e_{b_{n}}=1.29\times 10^{-5}$) and the bulk modulus ($e_{k}=2.423\times 10^{-6}$).
-In contrast, shear-related parameters exhibited substantial errors ($e_{a_{fs}}=-141.7$, $e_{b_{fs}}=0.92$).
-These trends stabilized within 8,0008{,}000 epochs, except for $b_{fs}$, which remained poorly identified.
-Introducing symmetric transmural fiber heterogeneity, i.e.
-in the circ-rad-hom slab, markedly improved fiber-sheet shear parameter identification, reducing relative errors to $e_{a_{fs}}=8.05\times 10^{-3}$ and $e_{b_{fs}}=0.17$, while maintaining accuracy for isotropic and anisotropic terms (see Table 4).
-Convergence for $b_{fs}$ was slower, requiring up to 30,00030{,}000 epochs.
-Further increasing fiber and sheet deformation heterogeneity (Figure 4), that is in the rad-heli-hom slab, yielded similar improvements , though inferring $b_{fs}$ remained challenging, with a relative error of $0.374$ and no clear stabilization after 100,000100{,}000 epochs.
-Other parameters exhibited relative errors below $10^{-4}$, comparable to the circ-rad-hom case (Table 4).
-Posterior distributions (Figure 6 - bottom row) confirmed these trends.
-Well-identified parameters displayed narrow, symmetric distributions, with standard deviations several orders of magnitude smaller than their means.
-In contrast, $a_{fs}$ and $b_{fs}$ exhibited skewed or truncated distributions, particularly in the circ-long-hom configuration, indicating limited identifiability.
-Notably, skewness magnitudes exceeded $|s_{\theta}|>0.5$ only for $a_{fs}$ and $b_{fs}$, while other inferred parameter distributions remained near-Gaussian.
-Overall, these results highlight that geometric homogeneity severely limits identifiability of shear-related parameters, while even moderate microstructural heterogeneity substantially improves inference quality.
-Figure 7: Orthotropic constitutive parameter inference from geometrically heterogeneous tissue slabs with varying microstructure.
-From left to right: a circ-long-het specimen, a circ-rad-het specimen, and a rad-heli-het specimen.
-From top to bottom: (i) evolution of inferred material parameters $\hat{\theta}$ over the epochs, (ii) evolution of relative parameter errors $e_{\theta}$, (iii) final relative errors at final epoch, and (iv) rescaled posterior distributions.
-Distributions are shown in grayscale when unbiased skewness $|s_{\theta}|<0.5$.
-Geometric heterogeneity significantly enhances identifiability of shear-related parameters, especially $b_{fs}$, reducing relative errors to sub-percent levels in circ-rad-het and rad-heli-het specimens.
-Posterior distributions for these parameters become narrower and less skewed compared to homogeneous counterparts.
-All other parameters maintain high accuracy and exhibit symmetric, low-variance distributions.
-These results underline the critical role of geometric heterogeneity in resolving parameters sensitive to shear-deformation modes.
-Focusing on geometrically heterogeneous samples (Figure 7), we evaluated three configurations with increasing microstructural complexity: fully homogeneous fibers (circ-long-het), symmetric transmural fiber rotation (circ-rad-het), and asymmetric helical fiber variation (rad-heli-het).
-For the circ-long-het configuration, all parameters were accurately identified except for $b_{fs}$ (relative error of roughly $1\%$).
-Relative errors remained low for isotropic parameters ($e_{a}=1.83\times 10^{-5}$, $e_{b}=1.22\times 10^{-5}$), anisotropic parameters in $f$ and $n$ directions ($e_{a_{f}}=-4.18\times 10^{-5}$, $e_{b_{f}}=9.77\times 10^{-6}$, $e_{a_{n}}=7.87\times 10^{-5}$, $e_{b_{n}}=-1.40\times 10^{-5}$), and the bulk modulus ($e_{k}=4.84\times 10^{-6}$).
-However, fiber-sheet shear coupling remained challenging, with errors of $e_{a_{fs}}=1.02\times 10^{-2}$ and $e_{b_{fs}}=0.879$.
-All parameters stabilized within 8,0008{,}000 epochs.
-Introducing symmetric fiber heterogeneity in the circ-rad-het slab markedly improved shear parameter inference, reducing $e_{a_{fs}}$ and $e_{b_{fs}}$ to $2.52\times 10^{-3}$ and $1.54\times 10^{-3}$, respectively.
-Isotropic and anisotropic parameters remained highly accurate (e.g., $e_{a}=-1.64\times 10^{-5}$, $e_{b}=-2.80\times 10^{-5}$), with convergence achieved within 5,0005{,}000 epochs.
-Further increasing microstructural complexity in the rad-heli-het configuration yielded comparable results.
-Relative errors for $a_{fs}$ and $b_{fs}$ were $1.14\times 10^{-3}$ and $6.90\times 10^{-3}$, respectively, while other parameters maintained relative errors below $10^{-4}$, as shown in the third row of Figure 7 and Table 4.
-Convergence required up to 10,00010{,}000 epochs, slightly slower than the circ-rad-het case.
-Posterior distributions (Figure 7 - bottom row) confirmed these trends.
-Parameters associated with isotropic and anisotropic invariants displayed narrow, near-Gaussian distributions with rescaled standard deviations between $10^{-3}$ and $10^{-5}$.
-In contrast, shear-related parameters ($a_{fs}$, $b_{fs}$) exhibited wider, skewed distributions, particularly in circ-long-het.
-However, skewness and standard deviations were notably reduced compared to homogeneous slab configurations, indicating improved identifiability.
-In general, these results demonstrate that geometric heterogeneity substantially enhances inference of shear-related parameters, effectively complementing microstructural variations to improve parameter identifiability.
-In general, we notice in the top rows of figures 6 and 7 that the parameters $b$ , $b_{f}$, $a_{n}$ and $a_{fs}$ drop in value after roughly 1,0001{,}000 epochs, only to reach a minimal value around 5,0005{,}000 epochs and then shoot up to converge to their optimal values.
-Similar behavior can be seen for the remaining parameters, though they stabilize at lower values than their initial values.
-In the second row of these figures we can also see interesting jumps between $10$ and 5,0005{,}000 epochs.
-Comparing these jumps to the parameter values in the top row reveals that these are the epochs where the estimated parameter value crosses the ground-truth value without achieving a stabilizing solution.
-This underlines the changing posterior geometry and thus also how the loss landscape changes due to parameter optimization.
-The summary of all parameter fitting results under noise free conditions is provided in C.
-5 Impact of measurement noise
-We examined inference robustness by introducing displacement measurement noise at two levels: low ($1\times 10^{-4}\,\mathrm{mm}$) and high ($1\times 10^{-3}\,\mathrm{mm}$), applied to geometrically heterogeneous training samples, because they showed structurally better performance in the previous sections.
-The choice of displacement noise levels is based on similar works (Flaschel et al., 2021; Joshi et al., 2022; Thakolkaran et al., 2022; Rahmani et al., 2013).
-We note here that Gaussian white noise can be considered harsher than what is present in realistic data because the spatial correlation between measurements is not taken into account.
-Some relevant works have implemented spatially correlated noise, or adapted fictitious noise to DIC measurements directly (Jafari et al., 2025; Peyraut and Genet, 2025).
-By considering Gaussian white noise, we aim to validate how robust our method is.
-[…]
-In Figure 8, we assess parameter inference performance with low levels of displacement noise ($\sigma=10^{-4}$) for geometrically heterogeneous tissue slabs with varying levels of microstructural heterogeneity.
-From left to right, results are shown for (i) a fully homogeneous circ-long-het specimen, (ii) a circ-rad-het specimen with moderate transmural fiber rotation, and (iii) a rad-heli-het specimen exhibiting pronounced helical fiber variation.
-In the top row of the figure, we observe that the circ-long-het configuration fails to converge for the parameter value $a_{fs}$.
-Consequently, more stochastic behavior as well as a slight downward drift can be observed in $b_{fs}$.
-Although most parameter values reach good relative error values ranging from $1\times 10^{-4}$ to $5\times 10^{-3}$, the parameters associated with $I_{4fs}$ show relative errors of the same order of magnitude as the true value.
-This error is further highlighted by the skewness and clear truncation of the posterior distribution shown in the bottom row.
-The spread in values also highlights the uncertainty around these parameter values.
-The helically distributed fiber configurations show no such skewness in any of their parameter distributions.
-They do obtain higher relative errors in their parameter values.
-Ranging between $5\times 10^{-3}$ to $1\times 10^{-1}$ for the circ-rad configuration and $3\times 10^{-4}$ to $2\times 10^{-2}$ (excluding $a_{fs}$ and $b_{fs}$).
-Strikingly, the shear parameters stabilize in the circ-rad-het and rad-heli-het configurations after roughly 5,0005{,}000 and 40,00040{,}000 epochs respectively, even though their relative errors are remarkably high.
-Combined with the lack of skewness and truncation in their distributions, this indicates that the parameter values have moved sufficiently far away from the prior and have settled in a (sub-optimal) local minimum.
-[…]
-We assess the performance of the parameter fitting approach for the high-noise case in a similar way.
-In Figure 9, we use the same lay-out as before and observe striking differences with the low displacement noise setting.
-All configurations now fail to characterize the shearing parameters $a_{fs}$ and $b_{fs}$.
-Moreover, we now also observe that the circ-rad-het configuration fails to show stabilizing behavior for the $a_{n}$ and $b_{n}$ parameters.
-We also observe an increase in relative error for all parameter values in all configurations.
-Remarkably, the circ-long-het specimen obtains the lowest relative errors for all parameters.
-The circ-rad-het configuration clearly performs the worst of the three showing relative errors ranging from $2\times 10^{-1}$ to $1$.
-Compared to the low-noise case this indicates a sudden drop in fitting performance.
-All configurations show overly skew distributions in the $a_{fs}$ and $b_{fs}$ parameters, although the circ-rad configuration also shows skewness in the distributions for $a_{n}$ and $b_{n}$.
-The remaining distributions show standard deviations orders of magnitude lower than the mean, indicating the confidence the model has in these values.
-Similarly to the low-noise case we see that parameter that do stabilize, do so around 8,0008{,}000.
-The rad-heli-het configuration seems to show stabilizing behavior for all its parameters roughly 2,0002{,}000 epochs later.
-Summarizing Figures 7, 8, and 9, we observe that the introduction of noise steadily degrades the relative error in the inferred parameter values.
-Furthermore, some inferred constitutive parameters start converging towards numerical zero as the amount of displacement measurement noise increases.
-D provides a summary of the obtained constitutive parameter values, the relative error achieved, and their standard deviation.
-6 Validation
-We reproduce multimodal, homogeneous loading protocols, i.e.
-classical incompressible biaxial tension and triaxial shear tests, following the experimental protocols of (Dokos et al., 2002; Sommer et al., 2015).
-Constitutive parameters inferred from our unsupervised single-shot heterogeneous biaxial protocol are then used to assess, in a controlled fashion, the predictive capability of the proposed framework relative to established multimodal calibration pipelines for myocardial tissue (Guan et al., 2018; Peirlinck et al., 2018; Kakaletsis et al., 2021; Martonová et al., 2024).
-Specifically, we prescribe homogeneous deformations with deformation gradient:
-$\bm{F}=\left[\begin{array}[]{lll}\lambda_{\rm{f}}&\gamma_{\rm{fs}}&\gamma_{\rm{fn}}\\ \gamma_{\rm{sf}}&\lambda_{\rm{s}}&\gamma_{\rm{sn}}\\ \gamma_{\rm{nf}}&\gamma_{\rm{ns}}&\lambda_{\rm{n}}\\ \end{array}\right].$ (29)
-For the biaxial tension cases, the fiber and normal stretches $\lambda_{\rm f}$ and $\lambda_{\rm n}$ are controlled, the sheet stretch follows from incompressibility, $\lambda_{\rm s}=1/(\lambda_{\rm f}\lambda_{\rm n})$, and all shear strains are set to zero, $\gamma_{\rm{fs}}=\gamma_{\rm{sf}}=\gamma_{\rm{nf}}=\gamma_{\rm{fn}}=\gamma_{\rm{sn}}=\gamma_{\rm{ns}}\equiv 0$.
-The hydrostatic pressure $p$ is determined by enforcing zero normal traction in the sheet direction, $\sigma_{\rm ss}=0$, which yields the following non-zero normal stresses:
-[…]
-Here $\psi_{\star}$ indicates the partial derivative of the strain energy density function with respect to the corresponding invariant $I_{\star}$, i.e.
-$\partial\psi/\partial I_{\star}$.
-In the triaxial shear tests, the principal stretches are fixed to unity, $\lambda_{\rm f}=\lambda_{\rm s}=\lambda_{\rm n}\equiv 1$, and only one shear component is activated per test; all others are set to zero, $\gamma_{\rm{fs}}=\gamma_{\rm{sf}}=\gamma_{\rm{nf}}=\gamma_{\rm{fn}}=\gamma_{\rm{sn}}=\gamma_{\rm{ns}}\equiv 0$, with a single component allowed to vary.
-Each test therefore excites a single shear mode.
-Given that $\gamma_{\star}\geq 0$, the resulting shear stresses are:
-$\sigma_{\rm sf}=2\,\gamma_{\rm fs}\,\psi_{1}+\psi_{\rm 4fs},$ $\sigma_{\rm fs}=2\,\gamma_{\rm sf}\!\left[\psi_{1}+\psi_{\rm 4ff}\right]+\psi_{\rm 4fs},$ (31)
-$\sigma_{\rm nf}=2\,\gamma_{\rm fn}\!\left[\psi_{1}+\psi_{\rm 4nn}\right],$ $\sigma_{\rm fn}=2\,\gamma_{\rm nf}\!\left[\psi_{1}+\psi_{\rm 4ff}\right],$
-$\sigma_{\rm ns}=2\,\gamma_{\rm sn}\!\left[\psi_{1}+\psi_{\rm 4nn}\right],$ $\sigma_{\rm sn}=2\,\gamma_{\rm ns}\psi_{1}.$
-Figure 10: Noise-free validation single-shot versus multi-modal biaxial and triaxial shear tests.
-Comparing Cauchy stresses based on the ground-truth parameter values, shown in tri markers, and the inferred noise-free full-field single-shot biaxial test parameter values sets, shown in shaded colors representing the $95\%$ confidence intervals resulting from the inferred posterior parameter set distributions.
-$R^{2}$ scores for inferred versus ground-truth stress values, for each individual biaxial tensile and triaxial shear test, are shown in the top left corners.
-Figure 10 reports noise-free validation across three heterogeneous configurations: circ–long–het (left), circ–rad–het (middle), and rad–heli–het (right).
-Each column compares Cauchy stresses from ground-truth parameters (tri markers) with predictions from the single-shot, noise-free full-field inference constitutive parameter sets (shaded 95% credible intervals) under classical biaxial tension and triaxial shear protocols.
-In the circ–long–het configuration, the inferred biaxial stress–stretch curves match the ground-truth with $R^{2}=1.00$ for both biaxial tension cases.
-The nf-, sn-, and ns-shear responses are also indistinguishable from the ground-truth, whereas the cross-fiber shear components show lower agreement, with $R^{2}=0.89$ for $\sigma_{\mathrm{fs}}$ and $R^{2}=0.77$ for $\sigma_{\mathrm{sf}}$ respectively.
-Aggregated over all tests in this configuration, the overall coefficient of determination is $R^{2}=0.96$.
-These deviations are consistent with limited excitation of the $I_{4fs}$ shear-coupling invariant in the circ–long-het configuration as shown in Figure 5.
-In the circ–rad–het configuration, the inferred constitutive parameter sets reproduce the ground-truth stresses across biaxial tension and all shear tests with $R^{2}=1.00$ throughout.
-Posterior confidence intervals remain tight, indicating precise recovery of the active parameters for this orientation.
-In the rad–heli–het configuration, inferred parameter set curves overlay the ground-truth for all biaxial tensile and shear tests with $R^{2}=1.00$.
-Overall, the noise-free validation shows that single-shot full-field inference generalizes across all three heterogeneous configurations, with slight deviations confined to cross-fiber shear in the circ–long–het case.
-Figure 11: High-noise validation single-shot versus multi-modal biaxial and triaxial shear tests.
-Comparing Cauchy stresses based on the ground-truth parameter values, shown in tri markers, and the inferred high-noise $\sigma=10^{-3}$ full-field single-shot biaxial test parameter values sets, shown in shaded colors representing the $95\%$ confidence intervals resulting from the inferred posterior parameter set distributions.
-$R^{2}$ scores for inferred versus ground-truth stress values, for each individual biaxial tensile and triaxial shear test, are shown in the top left corners.
-Figure 11 reports high-noise ($\sigma=10^{-3}$) validation across three heterogeneous configurations: circ–long–het (left), circ–rad–het (middle), and rad–heli–het (right).
-Each column compares Cauchy stresses from ground-truth parameters with predictions from the single-shot full-field inference under displacement noise $\sigma=10^{-3}$ (shaded 95% confidence intervals).
-In the circ–long–het configuration, inferred biaxial tension behavior remains perfectly aligned with the ground-truth stress behavior.
+$\hat{e}_{\theta}$ $=\frac{\mu_{\theta}-\theta_{\text{gt}}}{\theta_{\text{gt}}}\qquad$ $e_{\theta}$ $=\frac{\theta-\theta_{\text{gt}}}{\theta_{\text{gt}}}$ (29)
+Here, we define $\mu_{\theta}=\mu^{q}$ as the mean of the underlying untruncated normal distribution used in the guide distribution. Furthermore, we define the unbiased skewness $s_{\theta}$:
+$s_{\tilde{\theta}}$ $=\frac{\sqrt{N(N-1)}}{N-2}\cdot\frac{\frac{1}{N}\sum_{n=1}^{N}(\tilde{\theta}_{n}-\mu_{\tilde{\theta}})^{3}}{\left(\frac{1}{N}\sum_{n=1}^{N}(\tilde{\theta}_{n}-\mu_{\tilde{\theta}})^{2}\right)^{3/2}}\qquad\text{with}$ $\tilde{\theta}$ $=\frac{\theta-\mu_{\theta}}{\mu_{\theta}}$ (30)
+with $\tilde{\theta}$ denoting the relative error between samples taken from the guide distribution and the scaled posterior, $\mu_{\tilde{\theta}}$ denoting the mean of the distribution $\tilde{\theta}$, and $N$ equaling the number of samples drawn from the posterior distribution.
+Unless stated otherwise, we draw N=1, [^0] N=1{,}000 samples from the inferred guide distribution shown in the bottom rows of Figures 6 and 7, respectively, to visualize the shape, mean, and spread of the approximated posterior distribution.
+For completeness, the full set of converged parameter estimates, posterior scales (the standard deviation of the underlying and untruncated normal distribution), and relative errors for all six training specimens is reported in B (Tables 34).
+In the main text, we therefore focus on the parameter groups that drive the qualitative differences between Figures 6 and 7, and summarize the remaining terms compactly.
+Figure 6: Noise-free parameter inference on geometrically homogeneous tissue slabs with varying microstructure.
+From left to right: circ-long-hom, circ-rad-hom, and rad-heli-hom specimens.
+From top to bottom: (i) evolution of inferred material parameters $\theta$ over optimization epochs, (ii) evolution of absolute relative location parameter errors $|\hat{e}_{\theta}|$, (iii) absolute relative errors $|e_{\theta}|$ at the final epoch with corresponding 95% credibility interval, and (iv) rescaled posterior distributions $\tilde{\theta}$ (drawn from the variational guide).
+The plots show rapid convergence and small final errors for most parameters across all three specimens, while the shear-coupling parameters ($a_{fs}$, $b_{fs}$) display slower convergence and broader posteriors in several cases.
+Distributions are shown in grayscale when the absolute unbiased skewness satisfies $|s_{\tilde{\theta}}|<0.5$.
+Focusing on the geometrically homogeneous specimens (Figure 6), the plots show rapid convergence and narrow posteriors for the isotropic $(a,b,K)$ and primary anisotropic $(a_{f},b_{f},a_{n},b_{n})$ parameter groups across all three microstructural settings.
+Quantitatively, these parameters reach small final deviations, with the largest magnitudes remaining below $2.3\times 10^{-4}$ in this homogeneous-geometry group (Table 4).
+In contrast, the fiber–sheet shear-coupling terms $(a_{fs},\,b_{fs})$ are the dominant outliers.
+In the fully homogeneous circ-long-hom specimen, the shear-coupling terms remain poorly identified, with large final errors ($\hat{e}_{a_{fs}}=-1.42\times 10^{2}$ and $\hat{e}_{b_{fs}}=9.20\times 10^{-1}$; Table 4).
+Introducing transmural and helical fiber variation improves identifiability of $a_{fs}$ substantially (from $\mathcal{O}(10^{2})$ to $\mathcal{O}(10^{-3}\!-\!10^{-2})$; Table 4), while $b_{fs}$ remains the least accurate inferred parameter in the homogeneous-geometry group (e.g., $\hat{e}_{b_{fs}}=1.72\times 10^{-1}$ for circ-rad-hom and $3.74\times 10^{-1}$ for rad-heli-hom).
+Notably, $b_{fs}$ also converges more slowly than the remaining parameters: it stabilizes only after $\sim 3\times 10^{4}$ epochs in circ-rad-hom, and it shows no clear plateau by $10^{5}$ epochs in rad-heli-hom, whereas most other parameters stabilize within $\sim 8\times 10^{3}$ epochs (rows (i)–(ii)).
+These trends are consistent with the broader and more skewed posteriors for $(a_{fs},\,b_{fs})$ in Figure 6 (bottom row).
+Figure 7: Noise-free parameter inference on geometrically heterogeneous tissue slabs with varying microstructure.
+From left to right: circ-long-het, circ-rad-het, and rad-heli-het specimens.
+From top to bottom: (i) evolution of inferred material parameters $\theta$ over optimization epochs, (ii) evolution of absolute relative location parameter errors $|\hat{e}_{\theta}|$, (iii) absolute relative errors $|e_{\theta}|$ at the final epoch with corresponding 95% credibility interval, and (iv) rescaled posterior distributions $\tilde{\theta}$ (drawn from the variational guide).
+The plots show stable convergence for all parameter groups across the three specimens, and the posterior samples indicate narrower spreads for the shear-coupling parameters ($a_{fs}$, $b_{fs}$) compared with the homogeneous counterparts.
+Distributions are shown in grayscale when the absolute unbiased skewness satisfies $|s_{\tilde{\theta}}|<0.5$.
+For the geometrically heterogeneous specimens (Figure 7), the plots show stable convergence across parameter groups and generally narrower posterior spreads than in Figure 6.
+For the isotropic and primary anisotropic parameters, final deviations remain small overall, with the largest magnitudes staying below $5\times 10^{-4}$ across the heterogeneous-geometry group (Table 4).
+The most pronounced quantitative change occurs in the inferred shear-coupling parameters.
+With combined geometric heterogeneity and fiber rotation, both $a_{fs}$ and $b_{fs}$ converge to small final deviations: in circ-rad-het, $\hat{e}_{a_{fs}}=-2.52\times 10^{-3}$ and $\hat{e}_{b_{fs}}=-1.54\times 10^{-3}$, and in rad-heli-het, $\hat{e}_{a_{fs}}=-1.14\times 10^{-3}$ and $\hat{e}_{b_{fs}}=6.90\times 10^{-3}$ (Table 4).
+In contrast, without fiber rotation (circ-long-het), $b_{fs}$ remains comparatively inaccurate ($\hat{e}_{b_{fs}}=0.879$), consistent with the broader posterior in the corresponding column of Figure 7.
+Compared with the homogeneous-geometry cases in Figure 6, the rotated-fiber heterogeneous specimens (circ-rad-het and rad-heli-het) reach a stable $b_{fs}$ trajectory with small final deviations within $\mathcal{O}(10^{4})$ epochs ($\sim 5\times 10^{3}$ and $\sim 10^{4}$ epochs, respectively), rather than requiring $\sim 3\times 10^{4}$ to $10^{5}$ epochs, consistent with the tighter posterior spreads in row (iv).
+Across both Figures 6 and 7, several parameters exhibit non-monotonic transients before converging.
+In particular, multiple traces (including $b$, $b_{f}$, $a_{n}$, and $a_{fs}$) show a decrease after roughly $10^{3}$ epochs, followed by recovery towards their converged values by $\mathcal{O}(10^{4})$ epochs.
+Consistent excursions appear in the error plots over the same epoch range, reflecting epochs where the inferred trajectories pass through the ground-truth value (i.e., where the signed relative deviation changes sign) without immediately stabilizing.
+This behavior is consistent with strong parameter coupling and a nonconvex loss landscape in nonlinear orthotropic inference, where the optimizer can traverse multiple basins before stabilizing.
+4.2 Robustness to displacement noise
+We next assess robustness to displacement measurement noise by perturbing the measured full-field kinematics with additive Gaussian white noise at two levels: low ($\sigma_{u}=1\times 10^{-4}\,\mathrm{mm}$) and high ($\sigma_{u}=1\times 10^{-3}\,\mathrm{mm}$).
+We intentionally avoid modeling spatially correlated noise here.
+Instead, we use spatially uncorrelated Gaussian white noise as a conservative stress test of the inference framework, with noise magnitudes chosen in line with prior inverse and Bayesian identification studies [30, 31, 32, 72].
+Realistic displacement fields, for example those obtained with DIC, typically exhibit spatial correlations.
+Related work has incorporated correlated noise models or DIC-aware fictitious noise to mimic these effects more closely [29, 73].
+In this study, we deliberately adopt the simpler white-noise approximation to probe robustness under coarse, worst-case perturbations, and we apply it to the geometrically heterogeneous training specimens, since these provided the most reliable baseline inference performance in the noise-free setting (Section 4.1).
+Figure 8: Parameter inference on geometrically and microstructurally heterogeneous tissue slabs under low displacement measurement noise.
+From left to right: circ-long-het, circ-rad-het, and rad-heli-het specimens.
+From top to bottom: (i) evolution of inferred material parameters $\theta$ over optimization epochs, (ii) evolution of absolute relative location parameter errors $|\hat{e}_{\theta}|$, (iii) absolute relative errors $|e_{\theta}|$ at the final epoch with corresponding 95% credibility interval, and (iv) rescaled posterior distributions $\tilde{\theta}$ (drawn from the variational guide).
+The plots show that most parameters maintain stable convergence under low noise, while $a_{fs}$ and $b_{fs}$ exhibit the largest variability across specimens, reflected by slower convergence and broader posteriors.
+Here, uncorrelated Gaussian displacement noise with standard deviation $\sigma_{u}=10^{-4}\,\mathrm{mm}$ is added to the full-field kinematics.
+Distributions are shown in grayscale when the absolute unbiased skewness satisfies $|s_{\tilde{\theta}}|<0.5$.
+Figure 8 summarizes inference under low displacement noise for the three geometrically heterogeneous training specimens.
+Across all three microstructural settings, the trajectories in rows (i) and (ii) indicate that the low-noise setting largely preserves the convergence patterns of the noise-free heterogeneous baseline in Figure 7, and differences emerge most clearly in the shear-coupling terms.
+For circ-long-het, final relative errors for isotropic and primary anisotropic parameters stay within a narrow band, with representative values between $10^{-4}$ and $5\times 10^{-3}$ (Table 7).
+In contrast, the parameter trajectory for $a_{fs}$ does not settle to a clear plateau in row (i), and $b_{fs}$ exhibits increased stochasticity together with a mild downward drift.
+The corresponding relative errors for the shear-coupling terms are of order one, with $|\hat{e}_{a_{fs}}|\approx 1$ and $|\hat{e}_{b_{fs}}|\approx 0.9$.
+This behavior is mirrored in the posterior samples in row (iv), which show pronounced skewness and visible truncation for $(a_{fs},b_{fs})$, consistent with elevated uncertainty and weak practical identifiability under this noise level.
+Whereas in the noise-free heterogeneous setting (Figure 7) the shear parameter $a_{fs}$ could still be identified up to a relative error on the order of $10^{-2}$, the low-noise case now produces errors for both $a_{fs}$ and $b_{fs}$ of the same order of magnitude as their true values.
+In the helically distributed fiber specimens circ-rad-het and rad-heli-het, the posterior samples in row (iv) do not exhibit comparable skewness or truncation for the inferred parameters.
+Compared with the noise-free case, the final relative errors for isotropic and primary anisotropic parameters are slightly higher, remaining mostly between $5\times 10^{-3}$ and $10^{-1}$ for circ-rad-het and between $3\times 10^{-4}$ and $2\times 10^{-2}$ for rad-heli-het (excluding $a_{fs}$ and $b_{fs}$), while the majority of parameters still converge to clearly defined plateaus in rows (i) and (ii).
+For the shear-coupling terms, the posterior shapes of $(a_{fs},b_{fs})$ become more regular in row (iv), yet their error traces in row (ii) plateau at comparatively elevated levels.
+In the noise-free case (Figure 7), we achieved rotated-fiber shear-coupling errors $|\hat{e}_{a_{fs}}|$ of order $10^{-3}$ and $|\hat{e}_{b_{fs}}|$ of at most $10^{-2}$ in the circ-rad-het and rad-heli-het specimens (Table 4).
+Under low noise here, these errors increase to magnitudes between roughly $5\times 10^{-1}$ and order one, with the largest outlier observed for $b_{fs}$ in rad-heli-het (Table 7).
+Strikingly, the shear-coupling parameters stabilize in the circ-rad-het and rad-heli-het configurations after roughly 5,0005{,}000 and 40,00040{,}000 epochs respectively, even though their relative errors remain high.
+Combined with the lack of skewness and truncation in their posteriors, this pattern is consistent with the parameters having moved away from the prior and having settled in a suboptimal local minimum under low-noise conditions.
+Figure 9: Parameter inference on geometrically and microstructurally heterogeneous tissue slabs under high displacement measurement noise.
+From left to right: circ-long-het, circ-rad-het, and rad-heli-het specimens.
+From top to bottom: (i) evolution of inferred material parameters $\theta$ over optimization epochs, (ii) evolution of absolute relative location parameter errors $|\hat{e}_{\theta}|$, (iii) absolute relative errors $|e_{\theta}|$ at the final epoch with corresponding 95% credibility interval, and (iv) rescaled posterior distributions $\tilde{\theta}$ (drawn from the variational guide).
+The plots show increased variability and slower stabilization for several parameters compared with the low-noise case, with the shear-coupling parameters ($a_{fs}$, $b_{fs}$) exhibiting the largest spread in the posterior samples across specimens.
+Here, displacement noise with standard deviation $\sigma=10^{-3}\,\mathrm{mm}$ is added to the full-field kinematics.
+Distributions are shown in grayscale when the absolute unbiased skewness satisfies $|s_{\tilde{\theta}}|<0.5$.
+Figure 9 assesses parameter inference under high displacement measurement noise ($\sigma_{u}=10^{-3}\,\mathrm{mm}$) for the three geometrically heterogeneous specimens, using the same layout as in Figure 8.
+Relative to the low-noise case, the trajectories in rows (i) and (ii) show a marked loss of stability and accuracy, most prominently in the shear-coupling parameters $(a_{fs},b_{fs})$.
+Across all three specimens, these shear terms no longer attain reliable plateaus, their estimates for $a_{fs}$ collapse towards numerical near-zero values, and the corresponding relative errors remain of order one (Table 7).
+In addition, the primary anisotropic sheet-direction parameters $(a_{n},b_{n})$ in the circ-rad-het specimen fail to settle to stable values and reach relative errors close to unity, indicating a breakdown of identifiability in this parameter group under high noise.
+By contrast, the isotropic and primary anisotropic parameters remain comparatively more robust in circ-long-het, where final relative errors for these parameters mostly lie between $10^{-3}$ and $5\times 10^{-2}$.
+In the rad-heli-het specimen, these relative parameter errors fall between $5\times 10^{-2}$ and $10^{-1}$ (Table 7).
+The circ-rad-het configuration clearly performs worst overall, with many isotropic and primary anisotropic parameters exhibiting relative errors in the range $2\times 10^{-1}$ to $1$, representing a sharp degradation compared with both the noise-free and low-noise cases.
+Similar to the low-noise setting, parameters that do stabilize under high noise typically do so around 8,0008{,}000 epochs for circ-long-het and circ-rad-het, whereas the rad-heli-het specimen requires up to about 10,00010{,}000 epochs for its stable parameter groups.
+C provides the corresponding noisy hyperparameter and constitutive parameter estimates, together with their relative errors and posterior scale parameters (Tables 57), which quantitatively underpin the trends observed in Figures 79.
+Comparing Figures 7, 8, and 9 show a clear, parameter-selective degradation of identifiability as displacement noise increases from zero to low to high levels.
+In the noise-free heterogeneous baseline (Figure 7), isotropic and primary anisotropic parameters achieve small final deviations, with $|\hat{e}_{\theta}|$ typically below $5\times 10^{-4}$, while the shear-coupling terms $(a_{fs},b_{fs})$ are already more sensitive in circ-long-het but remain well identified in circ-rad-het and rad-heli-het (Table 4).
+Under low noise (Figure 8), these isotropic and primary anisotropic errors increase by roughly one to two orders of magnitude yet mostly remain below $10^{-1}$, whereas the shear-coupling errors in circ-long-het, circ-rad-het, and rad-heli-het rise to order one or larger, consistent with the broadened and, in some cases, truncated posteriors for $(a_{fs},b_{fs})$.
+At high noise (Figure 9), this trend intensifies: the shear-coupling parameters lose practical identifiability in all three specimens, with $|\hat{e}_{a_{fs}}|$ and $|\hat{e}_{b_{fs}}|$ remaining of order one and the corresponding posterior samples strongly skewed, while selected primary anisotropic parameters in circ-rad-het (notably $a_{f}$, $a_{n}$, and $b_{n}$) also exhibit near-unity relative errors and non-stabilizing trajectories.
+By contrast, most isotropic parameters and the bulk modulus $K$ retain comparatively moderate deviations even at the highest noise level, particularly in circ-long-het and rad-heli-het, indicating a degree of robustness in the volumetric and baseline stiffness sector despite the pronounced degradation in shear-related and directional anisotropic terms.
+5 Validation beyond training
+We now assess how well parameter sets inferred from single-shot heterogeneous biaxial tests generalize beyond the specific training protocol.
+To this end, we first compare Cauchy stress responses of inferred parameter sets under classical homogeneous uni-modal deformation modes against the ground-truth parameterized Holzapfel–Ogden law.
+Subsequently, we probe more complex mixed-modal deformation states, sampled from a physiologically constrained invariant space, and compare the resulting strain energy between inferred and ground-truth parameters to assess generalization under heterogeneous load paths.
+5.1 Uni-modal deformation mode validation
+We reproduce homogeneous, uni-modal loading protocols, i.e.
+classical incompressible biaxial tension and triaxial shear tests, following the experimental protocols of [10, 11].
+We consider incompressible biaxial tension, where $\lambda_{\rm f}$ and $\lambda_{\rm n}$ are prescribed and $\lambda_{\rm s}$ follows from incompressibility, and triaxial shear tests, where the principal stretches are fixed to unity and a single shear component is activated per test.
+The resulting stress expressions are summarized in D.
+Constitutive parameters inferred from our unsupervised single-shot heterogeneous biaxial protocol are then used to assess, in a controlled fashion, the predictive capability of the proposed framework relative to established multi-modal calibration pipelines for myocardial tissue [12, 15, 16, 26].
+Figure 10: Noise-free single-shot inference validation under homogeneous uni-modal loading.
+Cauchy stresses predicted by the ground-truth HO parameter set (tri markers) and by the inferred noise-free single-shot full-field posterior (shaded $95\%$ credible intervals) for three heterogeneous configurations: circ–long–het (left), circ–rad–het (middle), and rad–heli–het (right).
+Rows correspond to classical incompressible biaxial tension and triaxial shear modes.
+Panel-wise $R^{2}$ values quantify agreement between posterior mean predictions and ground-truth stresses.
+Noise-free inference generalizes perfectly across all modes for circ-rad-het and rad-heli-het ($R^{2}\approx 1$ throughout), while circ-long-het shows the largest deviations in cross-fiber shear ($\sigma_{\mathrm{fs}}$, $\sigma_{\mathrm{sf}}$).
+Figure 10 reports noise-free validation across the three geometrically heterogeneous configurations: circ–long–het (left), circ–rad–het (middle), and rad–heli–het (right).
+Each column compares Cauchy stresses from the ground-truth parameter set (tri markers) with predictions based on the noise-free single-shot full-field inference parameter sets (shaded $95\%$ credible intervals) under classical biaxial tension and triaxial shear tests.
+Leveraging parameters inferred from the noise-free circ-long–het slab, the inferred biaxial stress-stretch curves match the ground-truth mechanical response of the tissue with $R^{2}=1.00$ for both biaxial tension cases.
+The inferred nf-, sn-, and ns-shear responses are also indistinguishable from the ground-truth response, whereas the inferred cross-fiber shear behavior shows lower agreement with respect to the ground-truth model, with $R^{2}=0.89$ and $R^{2}=0.77$ for $\sigma_{\mathrm{fs}}$ and $\sigma_{\mathrm{sf}}$ respectively.
+Aggregated over all tests in this configuration, the overall coefficient of determination amounts to $R^{2}=0.96$.
+These deviations are consistent with the limited deformational excitation of the $I_{4fs}$ shear-coupling invariant in the circ-long-het specimen (Figure 5) and with the weaker identifiability of $(a_{fs},b_{fs})$ observed in the previous section.
+Leveraging the circ-rad-het slab, the inferred constitutive parameter sets perfectly reproduce the ground-truth stresses across all biaxial tension and shear tests with $R^{2}=1.00$ in every panel.
+Posterior confidence intervals remain tight, indicating precise recovery of the underlying constitutive parameters for this orientation, including the shear-coupling response.
+Focusing on the rad-heli-het specimen, the inferred stress curves again overlay the ground-truth for all biaxial tensile and shear tests with $R^{2}=1.00$.
+Overall, the noise-free validation shows that single-shot full-field inference generalizes reliably across all three heterogeneous configurations, with slight deviations confined to cross-fiber shear in the circ-long-het case, which is in agreement with the parameter-level identifiability analysis in Figure 7.
+Figure 11: High-noise single-shot inference validation under homogeneous uni-modal loading.
+Cauchy stresses predicted by the ground-truth HO parameter set (tri markers) and by the inferred single-shot full-field posterior under displacement noise $\sigma_{u}=10^{-3}\,\mathrm{mm}$ (shaded $95\%$ credible intervals) for circ–long–het (left), circ–rad–het (middle), and rad–heli–het (right).
+Rows correspond to the same biaxial tension and triaxial shear protocols as in Figure 10.
+Panel-wise $R^{2}$ values quantify agreement between posterior mean predictions and ground-truth stresses.
+High noise primarily degrades cross-fiber shear predictivity and widens credible intervals, most notably for circ-rad-het, while biaxial tension and non-cross-fiber shear modes remain well captured, particularly for circ-long-het and rad-heli-het.
+Figure 11 reports high-noise ($\sigma_{u}=10^{-3}$) validation for the same three heterogeneous single-shot slab specimens.
+Inferring from displacement and reaction force data in the circ–long–het specimen, the inferred biaxial tension behavior remains well aligned with the ground-truth, with panel-wise $R^{2}$ values close to one.
 Similarly, the nf-, sn-, and ns-shear responses retain correct magnitude and slope.
-Noise primarily affects the inferred cross-fiber shear stress $\sigma_{\mathrm{fs}}$ and $\sigma_{\mathrm{sf}}$ response, where credible intervals widen and small deviations between ground-truth and inferred behaviors appear.
-Aggregated over all tests in this configuration, the overall coefficient of determination is $R^{2}=0.96$.
-In the circ–rad–het configuration, the stresses computed from the inferred parameter sets under noisy full-field biaxial tensile tests, no longer fit the ground-truth stress curves perfectly.
-Mean trends are preserved across biaxial tensile and shear protocols, but panel-wise $R^{2}$ values are lower than those in the high-noise circ-long-het and rad-heli-het configurations.
-Nevertheless, the aggregated coefficient of determination for this configuration still remains $R^{2}=0.77$.
-In the rad–heli–het configuration, we see the best overall inferred stress responses across all biaxial tensile and shear stress responses, with $R^{2}$-values greater than 0.88 across the board.
-Here, the total aggregated coefficient of determination equals $R^{2}=0.93$.
-Overall, despite elevated displacement and thus deformation noise, the single-shot SVI inference approach maintains very good quantitative agreement with ground-truth stresses across all three heterogeneous configurations.
+Noise primarily affects the inferred cross-fiber shear stresses $\sigma_{\mathrm{fs}}$ and $\sigma_{\mathrm{sf}}$, where credible intervals widen and small systematic deviations between ground-truth and inferred curves appear.
+Aggregated over all tests in this configuration, the overall coefficient of determination remains high at $R^{2}=0.96$.
+Leveraging the circ–rad–het slab, stresses computed from the high-noise inferred parameter sets no longer match the uni-modal ground-truth responses perfectly.
+Mean trends are preserved across biaxial and shear protocols, but panel-wise $R^{2}$ values are lower than in the circ-long-het and rad-heli-het configurations, particularly in cross-fiber and sheet-direction shear.
+Nevertheless, the aggregated coefficient of determination still reaches $R^{2}=0.77$.
+Lastly, we note the best uni-modal predictive nature of the inferred parameter set from the high-noise rad-heli-het specimen, with with all panel-wise $R^{2}$-values greater than $0.88$ and an aggregated coefficient of determination of $R^{2}=0.93$.
+Overall, despite elevated displacement noise and the associated degradation in certain parameter estimates reported in Section 4.2, the single-shot stochastic variational inference framework maintains very good quantitative agreement with uni-modal ground-truth stresses across all three heterogeneous configurations, especially in the biaxial tension and non-cross-fiber triaxial shear modes.
+5.2 Mixed-modal deformation validation
+To probe generalization under mixed deformation modes, we sample deformation states in invariant spaces between the homogeneous loading paths.
+We draw $(I_{1},I_{4ff},I_{4nn},I_{4fs})$ uniformly within the minimum and maximum values induced by the homogeneous protocols in the uni-modal deformation, see also (31), and retain only physiologically admissible states.
+Specifically, we accept samples that satisfy near-incompressibility $0.9999\leq J\leq 1.0001$ and whose maximum principal Cauchy stress does not exceed $20.3~\mathrm{kPa}$, consistent with reported peak stresses for left-ventricular biaxial tests on human myocardial tissue [11].
+The invariant-to-kinematics reconstruction and stress evaluation used for filtering are detailed in E.
+We repeat this procedure until 5,0005{,}000 admissible mixed-modal states are obtained.
+Figure 12: Mixed-modal validation set construction in invariant space.
+Invariant trajectories induced by uni-modal biaxial and single-shear protocols (colored curves) compared with randomly sampled invariant combinations that satisfy (quasi-)incompressibility and a physiological bound on maximum principal stress detailed in E.
+The uni-modal protocols excite low-dimensional paths, whereas the accepted multi-modal deformation modes populate the space between these uni-modal paths, yielding mixed-modal deformation states for further and richer out-of-training validation.
+Figure 12 illustrates the resulting invariant coverage.
+The thin sweeping colored arcs represent the uni-modal loading conditions from the previous section, highlighting that these tests excite the invariants along low-dimensional paths.
+In contrast, our randomly sampled deformation gradients populate the space between these arcs and thus correspond to more complex mixed-modal testing protocols.
+The accepted samples lie between $3$ and $3.25$ in $I_{1}$, between $0.57$ and $1.32$ for $I_{4ff}$ and $I_{4nn}$, and between 0.5-0.5 and $0.5$ for $I_{4fs}$, and they are spread out approximately uniformly within this constrained region.
+This construction supports the view that the resulting invariant set spans a physiologically relevant range of deformation states that is well suited for out-of-training validation of the inferred versus ground-truth constitutive parameter set.
+Figure 13: Strain-energy generalization under mixed-modal deformations.
+Correlation between strain energy predicted by inferred parameter posteriors, $\Psi$, and the ground-truth strain energy, $\Psi^{\mathrm{GT}}$, over 5,0005{,}000 accepted mixed-modal deformation states sampled in invariant space (Figure 12).
+Results are shown for geometrically heterogeneous training specimens under both low and high displacement measurement noise, i.e.
+$\sigma_{u}=10^{-4}$ and $10^{-3}\,\mathrm{mm}$ respectively.
+Hexbin density emphasizes the dominant response region, and the black diagonal indicates perfect agreement.
+Despite noise-driven degradation of select parameter groups, predicted energies remain strongly correlated with the ground-truth across mixed-modal states (all $R^{2}\geq 0.9$).
+More specifically, we compute strain energies using both the ground-truth parameter set and the inferred parameter sets obtained from both low and high displacement measurement noise-affected heterogeneous training experiments.
+The correlation between the strain energy predicted by the inferred parameter sets $\Psi$ and the ground truth $\Psi^{\mathrm{GT}}$ is shown as a hexbin plot in Figure 13.
+The black diagonal centerline indicates optimal behavior, i.e.
+the inferred mechanical response matches the expected ground-truth response perfectly.
+In line with our prior results for the weakest identifiability of the noise-affected circ-long–het specimen in the parameter inference and uni-modal stress inference analysis, this specimen also yields the lowest strain-energy agreement across the multi-modal deformation modes with $R^{2}=0.91$ in the low-noise setting and $R^{2}=0.931$ in the high-noise setting.
+Across the three heterogeneous slabs tested, we observe larger deviations away from the diagonal, indicating that a subset of randomly sampled invariant combinations, together with corresponding posterior samples, lead to strain energies that differ more strongly from the ground-truth.
+The circ-rad-het specimen shows a very high coefficient of determination in the low-noise case ($R^{2}=0.989$), and its performance slightly degrades to $R^{2}=0.915$ in the high-noise case.
+Despite this drop, circ-rad-het exhibits the tightest clustering of hexbin intensities around the diagonal, suggesting that most mixed-modal states are still well captured, with a relatively small number of outliers contributing to the $R^{2}$ reduction.
+Lastly, the most consistent performing sample is the rad-heli-het with an $R^{2}$ score of $0.962$ in the low noise setting and $0.968$ in the high noise setting.
+Overall, the mixed-modal strain-energy validation indicates that, even though high displacement noise can substantially perturb individual parameter estimates—particularly the shear-coupling and some directional anisotropic terms—the inferred models retain strong predictive capability at the level of strain energy across a broad set of physiologically admissible deformation states (all $R^{2}\geq 0.9$).
+This finding complements the stress-based validation and suggests that the proposed single-shot full-field identification framework can generalize reliably beyond the specific training protocol, provided that geometric and microstructural heterogeneity are sufficiently rich and measurement noise remains within a moderate range.
 6 Discussion
 The ability to infer complex orthotropic hyperelastic models from a single full-field biaxial tensile testing experiment represents an interesting avenue in biomechanical tissue characterization.
-Traditional workflows typically rely on multiple specialized loading configurations and assume simplified model forms to ensure parameter identifiability (Dokos et al., 2002; Sommer et al., 2015; Avazmohammadi et al., 2018).
-In contrast, our framework leverages specimen-intrinsic microstructural (and geometrical) heterogeneity, combined with stochastic variational inference, to robustly recover full sets of nonlinear, cross-correlated material parameters under realistic noise conditions.
-This approach not only reduces experimental tissue manipulation but also expands opportunities to characterize spatially heterogeneous biological tissues, where sample availability or repeat testing is limited.
-Moreover, it lays the groundwork for practical integration of physics-informed Bayesian inference into future experimental biomechanics and computational modeling pipelines.
-Additionally, we offer the possibility of taking inter-sample and intra-sample variability into account for materials where this was previously not an option.
-Power of tissue testing heterogeneity in constitutive parameter identifiability.
-Our results show that introducing both geometric and microstructural heterogeneity in the training specimen substantially improves the identifiability of constitutive model parameters.
-The addition of a simple circular occlusion generates spatial strain gradients and local stress concentrations that increase the sensitivity of internal forces to variations in model parameters (Attar et al., 2022; Conde et al., 2023; Pottier et al., 2011; Pierron and Grédiac, 2021).
-Similarly, strategically aligning fiber orientations relative to the applied boundary conditions enhances the excitation of multiple deformation invariants (Zhang et al., 2022; Kakaletsis et al., 2021; Linka et al., 2021).
-These experimental design choices enabled stable recovery of all parameters associated with dominant energy modes ($I_{1}$, $I_{4ff}$), and even parameters associated with more subtle contributions ($I_{4fs}$) when measurement noise was low.
-Importantly, both forms of heterogeneity complemented each other: geometric features enriched the macroscopic stress states, while fiber orientation tuning ensured multiaxial activation of the microstructure.
-These findings suggest that future experimental protocols could deliberately incorporate both microstructural and geometrical tailoring to maximize parameter observability.
-Homogeneous loading conditions heavily impact orthotropic inference power.
-Samples with homogeneously distributed fiber orientations and without geometric features exhibited markedly worse parameter convergence (Pottier et al., 2011).
-In these homogeneous specimens, the dominant loading paths aligned primarily with the $I_{1}$ and $I_{4ff}$ invariants, while parameters associated with fiber-sheet coupling ($I_{4fs}$) received little excitation, as previously postulated by (Holzapfel and Ogden, 2009; Guan et al., 2018).
-As a result, posterior distributions for associated parameters such as $a_{fs}$ and $b_{fs}$ remained close to their prior, showing truncation effects and large uncertainty.
-Interestingly, under high measurement noise, homogeneous configurations experienced smaller decrease in performance in validation compared to heterogeneous specimens.
-This reflects a robustness trade-off: simpler configurations concentrate strain energy into a few modes, reducing sensitivity to noise at the cost of losing full identifiability.
-These findings highlight that while homogeneity can be beneficial under extreme noise, achieving accurate full-model inference requires careful experimental enrichment.
-Measurement noise, as expected, degrades inference performance.
-At a displacement noise level of $1\times 10^{-4}$ mm, the method still achieved low relative errors and tightly concentrated posterior distributions for most constitutive parameters.
-The shear related parameter $a_{fs}$ and $b_{fs}$, however, show larger deviation from the ground-truth.
-At $1\times 10^{-3}$ mm noise, significant broadening and bias in the posterior appeared, particularly for parameters linked to minor strain energy contributions.
-Namely, the fourth coupling invariant $I_{4fs}$ showed strong vulnerability to noise, with associated parameters $a_{fs}$ and $b_{fs}$ tending toward zero or staying close to the prior resulting in a truncated posterior distribution, which is in agreement with the loss in accuracy for anisotropic features in (Rahmani et al., 2013).
-This effectively removed their contribution to the strain energy function in the inferred models.
-Despite these deviations, the validation strain energy and stresses remained close to the ground-truth.
-This suggests that some parameter errors can be tolerated without substantial loss of predictive accuracy for macroscopic mechanical behavior.
-Improving noise robustness through enhanced experimental design, probabilistic displacement modeling, or variational reweighting strategies offer important future directions.
-(Marek et al., 2019; Jafari et al., 2025; Peyraut and Genet, 2025)
-Single-shot inferred material models validate well to multimodal testing protocols - even under high noise.
-We show that we can reproduce the behavior from classical homogeneous multimodal biaxial tensile and triaxial shear testing.
-Even at high noise levels, we see better overall agreement with the ground-truth stresses from a single heterogeneous full-field biaxial tensile test than we generally see for separate biaxial tensile-only, or triaxial shear-only training.
-Interestingly, the circ-rad-het configuration seems to be most sensitive to displacement noise when considering pure biaxial tensile and triaxial shear test loading, despite this configuration’s richer heterogeneous deformation profile shown in Figure 5.
-Both the circ-long-het and rad-heli-het samples show good extrapolative behavior, even under high noise levels when the shearing component in the strain energy vanishes, forming interesting slicing configurations for future experimental studies of myocardial tissue behavior.
+Conventional constitutive identification workflows for complex orthotropic tissues, suchs as the heart muscle, typically rely on assumed microstructural homogeneity across the tissue specimen and leverage multiple specialized uni-modal loading protocols to identify constitutive parameter sets [10, 11, 74].
+This reflects a broader challenge for orthotropic HO myocardium laws: parameter uniqueness is generally difficult to guarantee under limited excitation, motivating multi-modal protocols, sensitivity-informed reparameterizations, or model reduction [15, 75, 76, 77] This work demonstrates that a single heterogeneous biaxial experiment, when combined with full-field kinematics and sparse reaction-force measurements, can support uncertainty-aware inference of a nonlinear orthotropic Holzapfel–Ogden (HO) parameter set using stochastic variational inference.
+Beyond the identifiability perspective, this single-shot strategy reduces experimental tissue manipulation and is particularly attractive for heterogeneous biological tissues, where sample availability or repeat testing is limited.
+Across the considered slab configurations, two consistent findings emerge: (i) identifiability is primarily governed by the degree to which the experiment excites the relevant invariant subspace, and (ii) even when individual HO parameters are not recovered accurately under varying levels of displacement measurement noise, the inferred models can remain highly predictive for stresses and strain energy over both uni-modal and mixed-modal deformation states.
+These observations clarify when single-shot full-field inference can be expected to recover the full parameter set, and when it should instead be interpreted as delivering a reliable posterior predictive constitutive response with quantified uncertainty.
+Heterogeneity as an experimental design principle for orthotropic constitutive parameter inference.
+Dominant gains in constitutive parameter identifiability arise from deliberately enriching the deformation space within a single test.
+Microstructural heterogeneity, introduced through slab orientation inducing specific within-specimen heterogeneity through transmural fiber rotation, and geometric heterogeneity, introduced through a central occlusion, both broaden the activated distributions of $I_{1}$, $I_{4ff}$, $I_{4nn}$, and $I_{4fs}$ (Figures 4-5).
+These microstructural or geometric experimental protocol choices increase the sensitivity of the equilibrium residuals to the full parameter vector $\bm{\theta}$.
+In the noise-free setting, this explains the contrast between homogeneous and heterogeneous specimens in Figures 6-7.
+Most isotropic and primary anisotropic parameters are inferred robustly across configurations, while the shear-coupling parameters $(a_{fs},b_{fs})$ become practically identifiable only when the single slab biaxial tension experiment produces appreciable $I_{4fs}$ excitation.
+The circ-rad-het and rad-heli-het slab configurations provide the most consistent recovery of these terms, reflecting the combined effect of geometry-induced strain gradients and orientation-driven multiaxial activation of the microstructure.
+In this sense, our results translate the multi-modal calibration message from prior myocardium studies into a single-shot setting: rather than combining separate stretch and shear protocols, we aim to excite the same (otherwise weakly constrained) coupling modes within one heterogeneous experiment [10, 11, 15].
+These results further suggest that future protocols could deliberately combine geometrical and microstructural tailoring to maximize parameter observability within a single specimen.
+Experimental displacement measurement noise selectively impacts constitutive parameter identifiability.
+Displacement noise degrades constitutive parameter identifiability selectively rather than uniformly.
+For low noise ($\sigma_{u}=10^{-4}$ mm), Figures 8-9 show stable convergence and comparatively narrow posteriors for the isotropic and primary anisotropic parameters.
+In contrast, $(a_{fs},b_{fs})$ shows the largest posterior broadening and, in some cases, pronounced skewness or truncation of the inferred posterior parameter distribution.
+For high noise ($\sigma_{u}=10^{-3}$ mm), the shear-coupling terms lose practical identifiability across specimens, and additional degradation can appear in directional anisotropic parameters, most clearly in circ-rad-het.
+These results are consistent with global sensitivity and inverse uncertainty quantification studies of the underlying Holzapfel-Ogden constitutive model [43, 77, 78], which show that parameter sensitivity and practical identifiability depend strongly on the deformation modes and invariant support represented in the available data.
+Two implications follow.
+Full-field single-shot inference strongly benefits from experimental designs that amplify invariants associated with weaker energy modes.
+In addition, posterior-shape diagnostics such as truncation and skewness provide useful indicators of noise-driven practical non-identifiability and should be interpreted alongside point estimates.
+Identifiability versus predictive capability: how important is it to identify all the underlying constitutive parameters.
+An important result is that imperfect recovery of ground-truth HO parameters does not necessarily imply poor constitutive prediction, as also noted in prior HO calibration and in vivo estimation studies [75].
+Figures 10-11 show that posterior predictive stresses under classical homogeneous uni-modal protocols remain accurate in aggregate, even at high noise.
+Figure 13 shows strong strain-energy correlation across 5,000 physiologically admissible mixed-modal deformation states, with all $R^{2}\geq 0.9$.
+When viewed together with the parameter-level degradation in Figures 8-9, these results indicate that distinct parameter combinations can yield near-indistinguishable macroscopic responses over the rich explored invariant deformation space [15, 77] In such cases, the posterior predictive response, together with its credible interval becomes more relevant for downstream simulation than any single best-fit parameter vector and is consistent with inference perspectives that emphasize parameter ranges rather than unique point estimates.
+This perspective is particularly pertinent for soft tissues, where measurement noise, limited loading diversity, and intrinsic variability constrain practical identifiability.
+Some heterogeneous configurations remain more predictive under noise.
+Our validation results also reveal slab-specific trends.
+The circ-rad-het configuration can be more sensitive to displacement noise in the uni-modal and mixed-modal validations, despite exhibiting a rich heterogeneous deformation profile in Figure 5.
+This observation indicates that heterogeneity alone is not sufficient.
+Identifiability also depends on how displacement measurement noise perturbs the invariants that control weaker energy modes and how these perturbations propagate through the equilibrium residuals, which is consistent with the broader observation that the subset of HO terms supported by data depends on the deformation modes represented in the calibration dataset [15].
+In contrast, rad-heli-het exhibits the most consistent predictive performance across stress and energy validations in Figures 10-13, making it a promising slicing orientation when a single-shot protocol is prioritized and robust in-plane excitation of multiple microstructural constituent contributions is feasible.
+Outlook.
+Some limitations motivate future work.
+The results are obtained on synthetic data with known boundary conditions, known microstructural fields, and an assumed HO model form.
+Future experimental work will need to address boundary-condition uncertainty, microstructure dispersion error or misregistration, and model discrepancy [79, 80].
+More broadly, automated model discovery frameworks provide an interesting avenue to deal with both practical non-identifiability and model-form discrepancy in settings with complex loading patterns and varying amounts of measurement noise [16, 30, 32, 33, 53, 81] Complementary directions include explicit probabilistic models for displacement measurement error, experimental designs that target weakly activated invariants, and alternative variational weighting strategies to improve robustness under noisy full-field data [29, 73, 82].
+Variational families that capture posterior dependencies more accurately, combined with posterior predictive checks tailored to equilibrium residuals, would further strengthen uncertainty quantification.
+Finally, the present study focused on spatially homogeneous material parameters with spatially varying structure fields.
+Extending the framework to spatially varying parameter fields with appropriate regularization could be valuable for truly heterogeneous tissues, while raising additional identifiability and experimental design challenges [83, 84].
+Overall, our results support single-shot, uncertainty-aware identification of nonlinear orthotropic constitutive behavior from a single heterogeneous biaxial test.
+Our main experimental design message is that the relevant invariant subspace must be sufficiently activated, with particular emphasis on the weak shear-coupling mode.
+At the same time, the validation results show that posterior predictive accuracy can remain high even when individual parameters are noise-limited.
 7 Conclusion
 We presented a physics-informed stochastic variational framework that recovers the full constitutive parameter set of a nonlinear orthotropic hyperelastic model from a single biaxial stretch experiment.
-By exploiting intrinsic microstructural heterogeneity, or introducing simple geometric heterogeneity features when needed, the method identifies not only the dominant tensile parameters but also the shear-coupling terms that are typically weakly excited in a biaxial tensile test.
-Our results demonstrate that an effective balance between underlying microstructure and biaxial loading improves identifiability, and this balance can be set during slicing and tissue harvesting.
-Circumferential–radial slabs and rotated cuts that place both fiber and sheet directions in the loading plane enable full recovery under noise-free conditions.
-Under realistic displacement measurement noise, the approach remains robust and yields stresses and strain energies in close agreement with ground-truth simulations.
-Adding geometric heterogeneity, for example a central occlusion, systematically improves convergence and tightens posterior distributions, particularly for shear parameters.
+By exploiting intrinsic microstructural heterogeneity, or introducing simple geometric heterogeneity when needed, the method identifies not only the dominant tensile parameters but also shear-coupling terms that are typically weakly excited in biaxial tension.
+Our results demonstrate that an effective balance between underlying microstructure and biaxial loading improves identifiability, and that this balance can be set through slicing and tissue harvesting.
+Circumferential–radial slabs and rotated cuts that place both fiber and sheet directions in the loading plane enable recovery of the full parameter set under noise-free conditions.
+Under realistic displacement measurement noise, the approach remains robust and yields stresses and strain energies in close agreement with the ground-truth mechanical response, while providing posterior uncertainty estimates for parameters and predicted stresses and strain energies.
+Adding geometric heterogeneity, a central occlusion for example, systematically improves convergence and tightens posterior distributions, particularly for shear-related parameters.
 These results indicate that slicing protocols and experimental design can be used deliberately to enhance observability in full-field inverse characterization.
-More broadly, the framework enables localized characterization of orthotropic nonlinear biological tissues using a single-mode setup.
-This reduces experimental tissue manipulation burden, improves reproducibility, and supports studies where available tissue is limited.
+More broadly, the framework enables localized characterization of orthotropic nonlinear biological tissues from a single heterogeneous full-field experiment.
+The proposed framework therefore provides a practical route to single-shot, uncertainty-aware constitutive characterization that is compatible with emerging full-field experimental modalities.
+This reduces experimental tissue manipulation, improves reproducibility, and supports studies where available tissue is limited.
+Acknowledgements
+M.P. acknowledges support through the NWO Veni Talent Award 20058.
 Appendix A Tissue slab geometry and loading conditions
@@
 Figure 1: Synthetic data generation: training sample dimensions and loading conditions Tissue slab geometry and applied boundary conditions for the training samples.
-We subject all training samples undergo to a biaxial loading protocol $\lambda_{1}:\lambda_{2}-(1:2),(1:1),(2:1)$ up to $15\%$ of the initial length of the sample.
+We subject all training samples to a biaxial loading protocol $\lambda_{1}:\lambda_{2}-(1:2),(1:1),(2:1)$ up to $15\%$ of the initial length of the sample.
 All samples are of size 10 $\times$ 10 $\times$ 1 mm3.
 The geometrically heterogeneous tissue samples have a circular occlusion in the center with a radius of $R=1$ mm.
-Appendix B Analytical derivations
-We summarize the analytical derivation of the Holzapfel-Ogden model stress used in our code below:
-$\bm{P}\left(\bm{F},\bm{\theta}\right)=$ $2\bm{F}\frac{\partial\psi(\bm{F})}{\partial\bm{C}}$ (32)
-$=$ $2\bm{F}\left(\frac{\partial\bar{\psi}(\bar{I}_{1})}{\partial\bar{I}_{1}}\frac{\partial\bar{I}_{1}}{\partial\bm{C}}+\frac{\partial\bar{\psi}(\bar{I}_{4f})}{\partial\bar{I}_{4f}}\frac{\partial\bar{I}_{4f}}{\partial\bm{C}}\right$.
-$+\left.\frac{\partial\bar{\psi}(\bar{I}_{4s})}{\partial\bar{I}_{4s}}\frac{\partial\bar{I}_{4s}}{\partial\bm{C}}+\frac{\partial\bar{\psi}(\bar{I}_{4fs})}{\partial\bar{I}_{4fs}}\frac{\partial\bar{I}_{4fs}}{\partial\bm{C}}+K\frac{\bm{J}^{2}-1}{2}\bm{C}^{-1}\right)$
-$\frac{\partial\bar{\psi}(\bar{I}_{1})}{\partial\bar{I}_{1}}$ $=\frac{a}{2}exp(b(\bar{I}_{1}-3))$ (33)
-$\frac{\partial\bar{\psi}(\bar{I}_{4i})}{\partial\bar{I}_{4i}}$ $=a_{i}\langle\bar{I}_{4i}-1\rangle\exp(b_{i}\langle\bar{I}_{4i}-1\rangle^{2})$
-$\frac{\partial\bar{\psi}(\bar{I}_{4fs})}{\partial\bar{I}_{4fs}}$ $=a_{fs}\bar{I}_{4fs}exp(b_{fs}\bar{I}_{4fs}^{2})$
-$\frac{\partial\bar{I}_{1}}{\partial\bm{C}}$ $=J^{-2/3}\frac{\partial I_{1}}{\partial\bm{C}}+I_{1}\frac{\partial J^{-2/3}}{\partial\bm{C}}=\bm{J}^{-2/3}\bm{I}-\frac{1}{3}\bar{I}_{1}\bm{C}^{-1}$ (34)
-$\frac{\partial\bar{I}_{4i}}{\partial\bm{C}}$ $=J^{-2/3}\frac{\partial I_{4i}}{\partial\bm{C}}+I_{4i}\frac{\partial J^{-2/3}}{\partial\bm{C}}=\bm{J}^{-2/3}(\bm{i}\otimes\bm{i})-\frac{1}{3}\bar{I}_{4i}\bm{C}^{-1}$
-$\frac{\partial\bar{I}_{4fs}}{\partial\bm{C}}$ $=J^{-2/3}\frac{\partial I_{4fs}}{\partial\bm{C}}+I_{4fs}\frac{\partial J^{-2/3}}{\partial\bm{C}}=\frac{1}{2}\bm{J}^{-2/3}(\bm{f}\otimes\bm{s}+\bm{s}\otimes\bm{f})-\frac{1}{3}\bar{I}_{4fs}\bm{C}^{-1}$
-And the Kullback-Leibler derivation:
-$KL\left[\pi(\bm{\theta},V_{s},\sigma^{2}|\bm{b})||q(\bm{\theta},V_{s},\sigma^{2})\right]$ $=KL\left[\pi(\bm{\Phi}|\bm{b})||q(\bm{\Phi})\right]=\int q(\bm{\Phi})log\left[\frac{q(\bm{\Phi})}{\pi(\bm{\Phi}|\bm{b})}\right]d\bm{\Phi}$ (35)
-$=\int q(\bm{\Phi})log\left[\frac{q(\bm{\Phi})\pi(\bm{b})}{\pi(\bm{\Phi},\bm{b})}\right]d\bm{\Phi}$
-$=\int q(\bm{\Phi})log\left[\frac{q(\bm{\Phi})}{\pi(\bm{\Phi},\bm{b})}\right]d\bm{\Phi}+\int q(\bm{\Phi})log[\pi(\bm{b})]d\bm{\Phi}$
-$=\int q(\bm{\Phi})log\left[\frac{q(\bm{\Phi})}{\pi(\bm{\Phi},\bm{b})}\right]d\bm{\Phi}+log[\pi(\bm{b})]$
-$=-\int q(\bm{\Phi})log\left[\frac{\pi(\bm{b}|\bm{\Phi})\pi(\bm{\Phi})}{q(\bm{\Phi})}\right]d\bm{\Phi}+log[\pi(\bm{b})]$
-$=-\int q(\bm{\Phi})\left(log\left[\frac{\pi(\bm{\Phi})}{q(\bm{\Phi})}\right]+log[\pi(\bm{b}|\bm{\Phi})]\right)d\bm{\Phi}+log[\pi(\bm{b})]$
-$=-\int q(\bm{\Phi})log\left[\frac{\pi(\bm{\Phi})}{q(\bm{\Phi})}\right]d\bm{\Phi}-\int q(\bm{\Phi})log[\pi(\bm{b}|\bm{\Phi})]d\bm{\Phi}+log[\pi(\bm{b})]$
-Appendix C Noise-free orthotropic constitutive parameter inference results
+Appendix B Noise-free orthotropic constitutive parameter inference results
 This appendix contains the numerical results for the parameter fitting procedure when no displacement noise is applied to the experiments.
@@
 The hyperparameters are both for inverse-gamma distributions where $\alpha_{\star}$, $\beta_{\star}$ are the scale and rate parameters.
-Hyperparameters circ-long-hom circ-rad-hom rad-heli-hom circ-long-het circ-rad-het rad-heli-het
-$\alpha_{vs}$ $1.914\times 10^{1}$ $1.928\times 10^{1}$ $1.928\times 10^{1}$ $1.916\times 10^{1}$ $1.930\times 10^{1}$ $1.930\times 10^{1}$
+circ-long-hom circ-rad-hom rad-heli-hom circ-long-het circ-rad-het rad-heli-het
+$\alpha_{v_{s}}$ $1.914\times 10^{1}$ $1.928\times 10^{1}$ $1.928\times 10^{1}$ $1.916\times 10^{1}$ $1.930\times 10^{1}$ $1.930\times 10^{1}$
 $\alpha_{\sigma}$ $5.215\times 10^{6}$ $5.050\times 10^{6}$ $5.101\times 10^{6}$ $5.400\times 10^{6}$ $5.127\times 10^{6}$ $5.292\times 10^{6}$
-$\beta_{vs}$ $7.401\times 10^{7}$ $7.201\times 10^{7}$ $7.169\times 10^{7}$ $924\times 10^{7} 7.7$$20\times 10^{7} 7.7$$714\times 10^{7}$
+$\beta_{v_{s}}$ $7.401\times 10^{7}$ $7.201\times 10^{7}$ $7.169\times 10^{7}$ $924\times 10^{7} 7.7$$20\times 10^{7} 7.7$$714\times 10^{7}$
 $\beta_{\sigma}$ $1.901\times 10^{1}$ $1.836\times 10^{1}$ $1.854\times 10^{1}$ $1.834\times 10^{1}$ $1.733\times 10^{1}$ $1.789\times 10^{1}$
@@
 Parameters circ-long-hom circ-rad-hom rad-heli-hom circ-long-het circ-rad-het rad-heli-het
-mean:
+Inferred $\mu_{\bm{\theta}}$:
 $a$ $8.090\times 10^{-4}$ $8.090\times 10^{-4}$ $8.090\times 10^{-4}$ $8.090\times 10^{-4}$ $8.090\times 10^{-4}$ $8.090\times 10^{-4}$
@@
 $K$ $0\times 10^{-1} 1.000×10$$0\times 10^{-1}$ $0\times 10^{-1} 1.000×10$$0\times 10^{-1}$ $0\times 10^{-1} 1.000×10$$0\times 10^{-1}$
-variance:
+Inferred $\sigma_{\bm{\theta}}$:
 $\sigma_{a}$ $1.792\times 10^{-8}$ $1.513\times 10^{-8}$ $1.518\times 10^{-8}$ $1.756\times 10^{-8}$ $1.684\times 10^{-8}$ $1.600\times 10^{-8}$
 $\sigma_{b}$ $1.223\times 10^{-4}$ $9.733\times 10^{-5}$ $9.650\times 10^{-5}$ $1.268\times 10^{-4}$ $064\times 10^{-4} 1.1$$125\times 10^{-4}$
-$\sigma_{af}$ $4.608\times 10^{-9}$ $2.851\times 10^{-8}$ $2.972\times 10^{-8}$ $5.507\times 10^{-9}$ $4.784\times 10^{-8}$ $2.798\times 10^{-8}$
-$\sigma_{bf}$ $2.990\times 10^{-5}$ $2.072\times 10^{-4}$ $2.109\times 10^{-4}$ $3.661\times 10^{-5}$ $6.346\times 10^{-4}$ $2.022\times 10^{-4}$
-$\sigma_{an}$ $514\times 10^{-9} 1.1$$100\times 10^{-7}$ $8.551\times 10^{-9}$ $1.380\times 10^{-9}$ $1.084\times 10^{-7}$ $8.144\times 10^{-9}$
-$\sigma_{bn}$ $6.611\times 10^{-5}$ $3.699\times 10^{-2}$ $4.123\times 10^{-4}$ $7.526\times 10^{-5}$ $3.117\times 10^{-2}$ $3.880\times 10^{-4}$
-$\sigma_{afs}$ $1.906\times 10^{0}$ $4.088\times 10^{-5}$ $4.207\times 10^{-5}$ $6.564\times 10^{-5}$ $4.102\times 10^{-6}$ $1.034\times 10^{-5}$
-$\sigma_{bfs}$ $1.370\times 10^{1}$ $4.410\times 10^{0}$ $3.996\times 10^{0}$ $1.283\times 10^{1}$ $4.534\times 10^{-2}$ $2.728\times 10^{-1}$
+$\sigma_{a_{f}}$ $4.608\times 10^{-9}$ $2.851\times 10^{-8}$ $2.972\times 10^{-8}$ $5.507\times 10^{-9}$ $4.784\times 10^{-8}$ $2.798\times 10^{-8}$
+$\sigma_{b_{f}}$ $2.990\times 10^{-5}$ $2.072\times 10^{-4}$ $2.109\times 10^{-4}$ $3.661\times 10^{-5}$ $6.346\times 10^{-4}$ $2.022\times 10^{-4}$
+$\sigma_{a_{n}}$ $514\times 10^{-9} 1.1$$100\times 10^{-7}$ $8.551\times 10^{-9}$ $1.380\times 10^{-9}$ $1.084\times 10^{-7}$ $8.144\times 10^{-9}$
+$\sigma_{b_{n}}$ $6.611\times 10^{-5}$ $3.699\times 10^{-2}$ $4.123\times 10^{-4}$ $7.526\times 10^{-5}$ $3.117\times 10^{-2}$ $3.880\times 10^{-4}$
+$\sigma_{a_{fs}}$ $1.906\times 10^{0}$ $4.088\times 10^{-5}$ $4.207\times 10^{-5}$ $6.564\times 10^{-5}$ $4.102\times 10^{-6}$ $1.034\times 10^{-5}$
+$\sigma_{b_{fs}}$ $1.370\times 10^{1}$ $4.410\times 10^{0}$ $3.996\times 10^{0}$ $1.283\times 10^{1}$ $4.534\times 10^{-2}$ $2.728\times 10^{-1}$
 $\sigma_{K}$ $2.363\times 10^{-7}$ $6.853\times 10^{-7}$ $5.688\times 10^{-7}$ $2.296\times 10^{-7}$ $8.327\times 10^{-7}$ $5.612\times 10^{-7}$
@@
 Relative errors circ-long-hom circ-rad-hom rad-heli-hom circ-long-het circ-rad-het rad-heli-het
-$e_{rel,a}$ $4.765\times 10^{-5}$ $1.864\times 10^{-5}$ $2.925\times 10^{-5}$ $1.832\times 10^{-5}$ $1.644\times 10^{-5}$ $2.593\times 10^{-5}$
-$e_{rel,b}$ $2.298\times 10^{-6}$ $3.659\times 10^{-5}$ $3.854\times 10^{-5}$ $1.223\times 10^{-5}$ $2.795\times 10^{-5}$ $3.428\times 10^{-5}$
-$e_{rel,af}$ $5.495\times 10^{-5}$ $8.243\times 10^{-5}$ $7.541\times 10^{-5}$ $4.184\times 10^{-5}$ $1.356\times 10^{-5}$ $8.322\times 10^{-5}$
-$e_{rel,bf}$ $1.471\times 10^{-6}$ $4.349\times 10^{-5}$ $4.047\times 10^{-5}$ $9.769\times 10^{-6}$ $8.735\times 10^{-6}$ $3.724\times 10^{-5}$
-$e_{rel,an}$ $9.655\times 10^{-5}$ $1.844\times 10^{-4}$ $3.895\times 10^{-5}$ $7.867\times 10^{-5}$ $2.017\times 10^{-4}$ $4.702\times 10^{-5}$
-$e_{rel,bn}$ $1.290\times 10^{-5}$ $2.250\times 10^{-4}$ $5.078\times 10^{-5}$ $1.396\times 10^{-5}$ $4.682\times 10^{-4}$ $4.864\times 10^{-5}$
-$e_{rel,afs}$ $1.417\times 10^{2}$ $8.050\times 10^{-3}$ $1.901\times 10^{-3}$ $1.023\times 10^{-2}$ $2.520\times 10^{-3}$ $1.137\times 10^{-3}$
-$e_{rel,bfs}$ $9.201\times 10^{-1}$ $1.716\times 10^{-1}$ $3.743\times 10^{-1}$ $8.794\times 10^{-1}$ $1.535\times 10^{-3}$ $6.900\times 10^{-3}$
-$e_{rel,K}$ $2.423\times 10^{-6}$ $2.932\times 10^{-5}$ $4.502\times 10^{-5}$ $4.837\times 10^{-6}$ $1.661\times 10^{-5}$ $4.184\times 10^{-5}$
-Appendix D Noise-affected orthotropic constitutive parameter inference results
-This appendix contains the numerical results for the parameter fitting procedure when no displacement noise is applied to the experiments.
+$\hat{e}_{a}$ $4.765\times 10^{-5}$ $1.864\times 10^{-5}$ $2.925\times 10^{-5}$ $1.832\times 10^{-5}$ $1.644\times 10^{-5}$ $2.593\times 10^{-5}$
+$\hat{e}_{b}$ $2.298\times 10^{-6}$ $3.659\times 10^{-5}$ $3.854\times 10^{-5}$ $1.223\times 10^{-5}$ $2.795\times 10^{-5}$ $3.428\times 10^{-5}$
+$\hat{e}_{a_{f}}$ $5.495\times 10^{-5}$ $8.243\times 10^{-5}$ $7.541\times 10^{-5}$ $4.184\times 10^{-5}$ $1.356\times 10^{-5}$ $8.322\times 10^{-5}$
+$\hat{e}_{b_{f}}$ $1.471\times 10^{-6}$ $4.349\times 10^{-5}$ $4.047\times 10^{-5}$ $9.769\times 10^{-6}$ $8.735\times 10^{-6}$ $3.724\times 10^{-5}$
+$\hat{e}_{a_{n}}$ $9.655\times 10^{-5}$ $1.844\times 10^{-4}$ $3.895\times 10^{-5}$ $7.867\times 10^{-5}$ $2.017\times 10^{-4}$ $4.702\times 10^{-5}$
+$\hat{e}_{b_{n}}$ $1.290\times 10^{-5}$ $2.250\times 10^{-4}$ $5.078\times 10^{-5}$ $1.396\times 10^{-5}$ $4.682\times 10^{-4}$ $4.864\times 10^{-5}$
+$\hat{e}_{a_{fs}}$ $1.417\times 10^{2}$ $8.050\times 10^{-3}$ $1.901\times 10^{-3}$ $1.023\times 10^{-2}$ $2.520\times 10^{-3}$ $1.137\times 10^{-3}$
+$\hat{e}_{b_{fs}}$ $9.201\times 10^{-1}$ $1.716\times 10^{-1}$ $3.743\times 10^{-1}$ $8.794\times 10^{-1}$ $1.535\times 10^{-3}$ $6.900\times 10^{-3}$
+$\hat{e}_{K}$ $2.423\times 10^{-6}$ $2.932\times 10^{-5}$ $4.502\times 10^{-5}$ $4.837\times 10^{-6}$ $1.661\times 10^{-5}$ $4.184\times 10^{-5}$
+Appendix C Noise-affected orthotropic constitutive parameter inference results
+This appendix contains the numerical results for the parameter fitting procedure when displacement noise is applied to the experiments.
 In Table 5, we show the optimized hyperparameters at the final epoch.
@@
 The hyperparameters are both for inverse-gamma distributions where $\alpha_{\star}$, $\beta_{\star}$ are the scale and rate parameters.
-Hyperparameters circ-long-hom circ-rad-hom rad-heli-hom circ-long-het circ-rad-het rad-heli-het
-$\sigma=1e^{-4}$ $\sigma=1e^{-4}$ $\sigma=1e^{-4}$ $\sigma=1e^{-3}$ $\sigma=1e^{-3}$ $\sigma=1e^{-3}$
-$\alpha_{vs}$ 1.922e+01 1.930e+01 1.928e+01 1.922e+01 1.918e+01 1.923e+01
+circ-long-het circ-rad-het rad-heli-het circ-long-het circ-rad-het rad-heli-het
+$\sigma_{u}=1e^{-4}$ $\sigma_{u}=1e^{-4}$ $\sigma_{u}=1e^{-4}$ $\sigma_{u}=1e^{-3}$ $\sigma_{u}=1e^{-3}$ $\sigma_{u}=1e^{-3}$
+$\alpha_{v_{s}}$ 1.922e+01 1.930e+01 1.928e+01 1.922e+01 1.918e+01 1.923e+01
 $\alpha_{\sigma}$ 2.615e+06 2.678e+06 2.683e+06 8.849e+05 1.047e+06 9.640e+05
-$\beta_{vs}$ 1.977e+07 2.076e+07 2.479e+07 2.168e+06 1.052e+06 2.585e+06
+$\beta_{v_{s}}$ 1.977e+07 2.076e+07 2.479e+07 2.168e+06 1.052e+06 2.585e+06
 $\beta_{\sigma}$ 3.598e+01 3.166e+01 3.250e+01 1.079e+02 8.942e+01 9.801e+01
@@
 The corresponding standard deviation has also been reported.
-Parameters circ-long-hom circ-rad-hom rad-heli-hom circ-long-het circ-rad-het rad-heli-het
-$\sigma=1e^{-4}$ $\sigma=1e^{-4}$ $\sigma=1e^{-4}$ $\sigma=1e^{-3}$ $\sigma=1e^{-3}$ $\sigma=1e^{-3}$
-mean:
+Parameters circ-long-het circ-rad-het rad-heli-het circ-long-het circ-rad-het rad-heli-het
+$\sigma_{u}=1e^{-4}$ $\sigma_{u}=1e^{-4}$ $\sigma_{u}=1e^{-4}$ $\sigma_{u}=1e^{-3}$ $\sigma_{u}=1e^{-3}$ $\sigma_{u}=1e^{-3}$
+Inferred $\mu_{\bm{\theta}}$:
 $a$ 8.086e-04 8.051e-04 8.054e-04 8.101e-04 9.909e-04 9.237e-04
@@
 $K$ 1.000e-01 9.952e-02 1.001e-01 9.922e-02 7.003e-02 9.155e-02
-variance:
+Inferred $\sigma_{\bm{\theta}}$:
 $\sigma_{a}$ 6.561e-08 5.847e-08 5.717e-08 5.420e-07 3.137e-07 4.306e-07
@@
 The relative error at the final epoch between the parameter values and the ground-truth for samples with circular occlusion, additive displacement noise at high ($\sigma=1e-3$) and low ($\sigma=1e-4$) levels and various microstructures.
-Parameters circ-long-hom circ-rad-hom rad-heli-hom circ-long-het circ-rad-het rad-heli-het
-$\sigma=1e^{-4}$ $\sigma=1e^{-4}$ $\sigma=1e^{-4}$ $\sigma=1e^{-3}$ $\sigma=1e^{-3}$ $\sigma=1e^{-3}$
-$e_{rel,a}$ 4.632e-04 4.835e-03 4.466e-03 -1.355e-03 -2.248e-01 -1.418e-01
-$e_{rel,b}$ -1.793e-04 -8.132e-03 -2.026e-03 -4.671e-02 -1.738e-01 -5.960e-02
-$e_{rel,af}$ -4.334e-03 -4.031e-02 -1.806e-02 -4.855e-02 -9.203e-01 -6.801e-02
-$e_{rel,bf}$ 1.847e-03 3.258e-02 1.073e-02 3.515e-02 5.563e-01 5.470e-02
-$e_{rel,an}$ 5.049e-03 -4.218e-02 -7.769e-03 -2.432e-02 1.000e+$0$ -9.956e-02
-$e_{rel,bn}$ -1.119e-03 9.979e-02 2.589e-04 2.969e-02 9.946e-01 2.204e-02
-$e_{rel,afs}$ 1.000e+$0$ 5.264e-01 9.225e-01 1.000e+$0$ 1.000e+$0$ 1.000e+$0$
-$e_{rel,bfs}$ 9.069e-01 -5.964e-01 -3.369e+$0$ 9.118e-01 9.657e-01 9.974e-01
-$e_{rel,K}$ -1.137e-04 4.845e-03 -1.068e-03 7.770e-03 2.997e-01 8.453e-02
+Parameters circ-long-het circ-rad-het rad-heli-het circ-long-het circ-rad-het rad-heli-het
+$\sigma_{u}=1e^{-4}$ $\sigma_{u}=1e^{-4}$ $\sigma_{u}=1e^{-4}$ $\sigma_{u}=1e^{-3}$ $\sigma_{u}=1e^{-3}$ $\sigma_{u}=1e^{-3}$
+$\hat{e}_{a}$ 4.632e-04 4.835e-03 4.466e-03 -1.355e-03 -2.248e-01 -1.418e-01
+$\hat{e}_{b}$ -1.793e-04 -8.132e-03 -2.026e-03 -4.671e-02 -1.738e-01 -5.960e-02
+$\hat{e}_{a_{f}}$ -4.334e-03 -4.031e-02 -1.806e-02 -4.855e-02 -9.203e-01 -6.801e-02
+$\hat{e}_{b_{f}}$ 1.847e-03 3.258e-02 1.073e-02 3.515e-02 5.563e-01 5.470e-02
+$\hat{e}_{a_{n}}$ 5.049e-03 -4.218e-02 -7.769e-03 -2.432e-02 1.000e+$0$ -9.956e-02
+$\hat{e}_{b_{n}}$ -1.119e-03 9.979e-02 2.589e-04 2.969e-02 9.946e-01 2.204e-02
+$\hat{e}_{a_{fs}}$ 1.000e+$0$ 5.264e-01 9.225e-01 1.000e+$0$ 1.000e+$0$ 1.000e+$0$
+$\hat{e}_{b_{fs}}$ 9.069e-01 -5.964e-01 -3.369e+$0$ 9.118e-01 9.657e-01 9.974e-01
+$\hat{e}_{K}$ -1.137e-04 4.845e-03 -1.068e-03 7.770e-03 2.997e-01 8.453e-02
+Appendix D Uni-modal deformation mode validation protocols and stress expressions
+This appendix summarizes the homogeneous deformation modes used for validation in Section 5.1.
+We reproduce incompressible biaxial tension and triaxial shear protocols following [10, 11] and evaluate the resulting Cauchy stress responses.
+We prescribe a spatially homogeneous deformation gradient
+$\bm{F}=\left[\begin{array}[]{lll}\lambda_{\rm{f}}&\gamma_{\rm{fs}}&\gamma_{\rm{fn}}\\ \gamma_{\rm{sf}}&\lambda_{\rm{s}}&\gamma_{\rm{sn}}\\ \gamma_{\rm{nf}}&\gamma_{\rm{ns}}&\lambda_{\rm{n}}\\ \end{array}\right].$ (31)
+For biaxial tension, $\lambda_{\rm f}$ and $\lambda_{\rm n}$ are prescribed, incompressibility implies $\lambda_{\rm s}=1/(\lambda_{\rm f}\lambda_{\rm n})$, and all shear components are set to zero.
+The hydrostatic pressure is obtained by enforcing $\sigma_{\rm ss}=0$, which yields the following non-zero stresses
+[…]
+Here $\psi_{\star}$ denotes $\psi_{\star}=\partial\psi/\partial I_{\star}$.
+For the triaxial simple shear cases we consider, $\lambda_{\rm f}=\lambda_{\rm s}=\lambda_{\rm n}\equiv 1$ and a single shear component is activated per test while all others are set to zero.
+For $\gamma_{\star}\geq 0$, the non-zero shear stresses read
+$\sigma_{\rm sf}=2\,\gamma_{\rm fs}\,\psi_{1}+\psi_{\rm 4fs},$ $\sigma_{\rm fs}=2\,\gamma_{\rm sf}\!\left[\psi_{1}+\psi_{\rm 4ff}\right]+\psi_{\rm 4fs},$ (33)
+$\sigma_{\rm nf}=2\,\gamma_{\rm fn}\!\left[\psi_{1}+\psi_{\rm 4nn}\right],$ $\sigma_{\rm fn}=2\,\gamma_{\rm nf}\!\left[\psi_{1}+\psi_{\rm 4ff}\right],$
+$\sigma_{\rm ns}=2\,\gamma_{\rm sn}\!\left[\psi_{1}+\psi_{\rm 4nn}\right],$ $\sigma_{\rm sn}=2\,\gamma_{\rm ns}\psi_{1}.$
+Appendix E Mixed-modal invariant sampling and admissibility filtering
+We sample mixed-modal deformation states by drawing invariants uniformly between the extrema induced by the homogeneous deformation modes in Eq.
+(31).
+We then reconstruct $\bm{C}$ in the local fiber-sheet-normal basis, compute $\sqrt{\bm{C}}$ via Schur decomposition, and evaluate the maximum principal Cauchy stress used for admissibility filtering.
+We take the minimum and maximum values of the invariants induced by Eq. (31) and uniformly sample in between:
+$I_{\star}\sim\mathcal{U}(I_{\star}^{\min},I_{\star}^{\max}),\quad I_{\star}\in\{I_{1},I_{4ff},I_{4nn},I_{4fs}\}.$ (34)
+where $I_{\star}^{\min}$ and $I_{\star}^{\max}$ denote the minimum and maximum deformation invariants seen in the uni-modal biaxial tension and triaxial simple shear experiments studied in Section 5.1.
+We reconstruct the right Cauchy-Green tensor using the identity $I_{1}=I_{4ff}+I_{4ss}+I_{4nn}$ and assuming the same orientation basis as in the uni-modal spatially homogeneous loading case:
+$C_{11}=I_{4ff},\qquad$ $C_{13}=C_{31}=C_{23}=C_{32}=0,$ (35)
+$C_{22}=I_{1}-I_{4ff}-I_{4nn},\qquad$ $C_{12}=C_{21}=I_{4fs},$
+$C_{33}=I_{4nn}.$
+Using the Schur decomposition [85], we compute the right stretch tensor $\bm{U}=\sqrt{\bm{C}}$ and use the polar decomposition $\bm{F}=\bm{R}\bm{U}$.
+We then evaluate the Cauchy stress and extract the maximum principal stress:
+$\sigma_{1}$ $=\mathcal{F}_{1}(J^{-1}\bm{F}\bm{S}\bm{F}^{T})=\mathcal{F}_{1}(J^{-1}\bm{R}\bm{U}\bm{S}\bm{U}\bm{R}^{T})=\mathcal{F}_{1}(J^{-1}\bm{U}\bm{S}\bm{U}),$ (36)
+where $\mathcal{F}_{1}(\bm{\sigma})=\max\operatorname{eig}(\bm{\sigma})$ and $\bm{S}=\partial\Psi/\partial\bm{C}$.
+We accept only physiologically relevant samples satisfying $0.9999\leq J\leq 1.0001$, aligned with the tissue’s assumed quasi-incompressibility, and $\sigma_{1}\leq 20.3~\mathrm{kPa}$, aligned with the maximum principal stresses measured in experimental protocols on human myocardial tissue [11] (see Table 3 showing peak Cauchy biaxial tensile stresses in the left ventricle).
+We repeat sampling until 5,0005{,}000 states are retained.